A Complete Bibliography of Publications in Journal of Computational and Applied Mathematics (2025–2029)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

             E-mail: beebe@math.utah.edu, beebe@acm.org,
                       beebe@computer.org (Internet)
           WWW URL: https://www.math.utah.edu/~beebe/

10 June 2025 Version 1.16

Title word cross-reference

(2 + 1) [MS25b, MO25]. 0 [Kar25b]. 1 [AKKM25, CJ25, Kar25b]. 2 [ABC^{+}25, LMZ25, MKE26]. 2 × 2 [KR25]. α [SAP^{+}25]. AXB = C [NZ25]. B [GP25a]. c [XQW25]. C^{0} [GS25]. C^{1} [GP25a, LM25a]. D [ABC^{+}25]. E [NGU25a]. H(curl^{2}) [WLZZ25]. H_{∞} [SC25]. K [XGH25, ZHX25, DJ25, GJ25, JFP25]. k/n(G) [YSP25]. L1 [CG25]. L^{2} [ZC25a, Xie25]. l_{1/2} [CLJX25]. L_{2} [LSPV25]. L_{∞} [ZZJJ25]. M [YK25]. R [CdlHG25]. M [NDSZ25]. n [HY25b, Ran25]. n: G [DJ25]. p [CK25, ZLLL25]. q [KR25]. R [XNL25]. R^{n} [BS25a, CA25b]. R^{s} [DP25]. R_{0} [SWD25]. ρ [TST25]. s [LBM25]. t [WCYL26]. W̃_{2}(m, m − 1) [SHK25]. φ [TST25]. w [AAN25].

-bar [ABC^{+}25]. -conforming [WLZZ25]. -convex [LBM25]. -dimensional [MS25b, MO25, Ran25]. -Laplacian [CK25]. -lateral [XGH25]. -means [XQW25]. -metric [YK25]. -norm [Xie25]. -numerical [KR25]. -order [ZHX25]. -out-of- [DJ25, SY25]. -out-of-n [GJ25]. -person [HY25b]. -record [JFP25]. -splines [GP25a]. -stable [SAP^{+}25]. -tensors [NDSZ25]. -type [SWD25].

1 [Ano25a, Ano25b, Ano25c, Ano25d, Ano25e, Ano25f, Ano25g, Ano26a]. 15 [Ano25h, Ano25i, Ano25j, Ano25k, Ano25l, Ano25m].

2025 [Ano25n, Ano25a, Ano25b, Ano25h, Ano25c, Ano25i, Ano25-34, Ano25d, Ano25j, Ano25-36, Ano25-35, Ano25e, Ano25k, Ano25l, Ano25f, Ano25-37, Ano25g, Ano25m, Ano25-38]. 2026 [Ano26a].

421 [MR25].

Abbas [NGU25a]. abrupt [WBH^{+}26]. abscissae [NT25a]. absolute [LC25b, LL25]. accelerated [BRS25, FLC^{+}25, HHF25, LTL^{+}25, SMDW25, TTZL25, ZLZ25]. acceleration [FLC^{+}25, TYJH25]. accuracy [BAPPF25b, MCL25, QSW25]. accurate [PFYH25, Wan25b, YWW25]. acoustic [LLG26]. acquired [Wan25c]. Activation [LU25]. active [HU25]. actuarial [HY25a]. ADAM [SWWY25]. ADAM-type [SWWY25]. Adaptive [MWTW25, PP25, TLSW25, CLYY25, CGuRZ25, HLF^{+}25, JCCZ25, JWS^{+}25, KS25b, LLG26, OAO25, TGDC25, XL25, ZMG25, ZLC^{+}25, CCGRR26]. adaptivity [BS25b, KS25c, YTT25]. addition [LU25]. additive [Ran25]. adjustments [AGGPDRC25]. ADMM [ZLLL25, ZPGY25]. ADMM-type [ZLLL25]. adopted [HJC26]. ADR [EES25, YWS25]. Advantages [VN25]. advection [CFCP26, FBM25, XGYH25, EES25]. adversarial [AGGM25]. affinely [Sah25]. against [SBAM25]. agent [LSCZ25]. aggregation [PLW^{+}25]. aggregation-based [PLW^{+}25]. aging [NRS25]. Agros [KPK25]. AHP [Wan25c]. airflow [LB25]. algebra [ZZM25]. Algebraic [RV25, DGMN25, PLW^{+}25]. algorithm [AK25a, AIF26, dAB25, BE26, CPSW25, CLJX25, CXTL25, CLY25b, CLQY25, Che25b, CXLP25, Don25, FA25, GCT25, GLTW25, GJWV25, HBG25, HLZ25, HGZL25, IRS25, JN25a, JWZZ25, JJM25, LZFP25, LC25b, LLZ25a, LTL^{+}25, LWF25b, LJY25, LY25, LR25, LZZ^{+}25b, OAO25, SK25b, SQLZ25, SWWY25, SSW^{+}25, SWD25, TTZL25, TSL25, TYJH25, TZ25b, WKQ25, WAXM25, XNL25, XLLJ25, YLC25, YLW^{+}25, YYJZ25, ZLZ25, ZLC^{+}25, ZLLL25, ZPGY25, ZHLD25]. algorithms [CGuRZ25, FBZ25, HHF25, JWS^{+}25, LSCZ25, MYL25, MG25, NZ25, SSLP25, ZS25]. Allen [HKH^{+}25, HLQF25, WWC25, XC25, ZSYZ25]. ALM [YLW^{+}25]. ALM-based [YLW^{+}25]. Almost [CCLH25, BS25c]. Alpha [PW25, WZZ25]. Alpha-maxmin [WZZ25]. Alpha-robust [PW25]. altering [LU25]. alternating [XNL25, NDSZ25]. alternative [LWLW25, WLL25]. American [GHMN25, YL25b, ZZ25b]. Analysis [CCT25, DLR25, JFP25, LCZ25, MY25, WZTL25, ADY25, AXF26, BFQ25, CLWY25, CG25, DCT^{+}25, DL25, DH25, FLLZ25, FFDL25, FADA25, JL25, Kar25a, KR26, LW25, LTZ25, LP25b, LC25a, LGDN25, LMZ25, LR25, Moh25, MS25b, PXZZ25, QZG^{+}25, SX25, SDS25, SGRS25, SBAM25, TC25, VS25, WAXM25, WCW25, XK25, YSP25, ZLS25, ZSYZ25, ZC25b, ZFCM25]. analysis-suitable [XK25]. analytic [SH25]. Analytical [MS25b]. analyzers [WCYL26]. analyzing [HLL25, YAO^{+}25b]. Anderson [NDSZ25]. aneurysm [DMS^{+}25]. anisotropic [FSC25, KS25b, PXZZ25, WX25a]. annular [YWW25]. Appl [MR25]. applicability [LP25b]. Application [CA25b, FKVZ25, AIF26, AP25, BRPP25, CXTL25, CLY25b, CLQY25, CXLP25, FBM25, GP25a, IRS25, JN25b, LBM25, LC25b, LWLW25, LY25, MZL25, Moh25, QZG^{+}25, TSM25, TA25, WLCM25, XQW25, YLC25, YWS25, Yao25a, YYJZ25, ZYTH25, ZLLL25, ZPGY25, ZDX25]. applications [HY25a, uDJKA26, SK25a, SW25, SWD25, TSL25, WuR25, WZ25b, ZHLD25]. applied [EYJKL26, MYL25, MR23, MR25, XGYH25]. approach [ADJL25, CDPT25, GBB25, HBG25, SDS25, SGRS25]. Approaches [ZFCM25, CK25]. Approximate [KSC25, ZNL25]. approximated [CAN25]. approximates [Xie25]. approximating [SHK25]. Approximation [KCM25, AKKM25, CA25a, CdlHG25, GBFNR26, GP25b, HJ25, JN25b, NGU25a, Öza25, SZD25, TC25, XNL25, XLLJ25, YT25, YC25]. approximations [GHZ25, KR26, YK25]. April [Ano25n]. Aquila [VKA25]. arbitrary [AM25, LC25c, PY25]. Archimedean [LL25]. argument [FA25, Pho25]. argument-based [FA25]. arising [BCdR25, COPPS25, NSW25]. arrangement [MT25]. arrival [BC25]. arteries [BPT26]. artificial [Gho25]. Asian [AAN25]. assessing [GHMN25]. asset [LK25]. asset-liability [LK25]. assets [LG25, YL25b, ZFCM25]. assimilation [DLR25, WJ25]. associated [BCD25]. asymmetric [SK25b]. asymptotic [ADJL25, EES25, FLS25, ZC25b]. Augmented [WX25b, HHF25, ZNL25]. August [Ano25a, Ano25b, Ano25h]. authentic [LGDN25]. Auto [Wan25a]. Auto-stabilized [Wan25a]. automatic [Don25]. autoregressive [HY25a, SW25]. average [XLLJ25]. AVF [MSR26].

B [LWL^{+}25, RV25, RK25]. B-spline [LWL^{+}25, RV25, RK25]. Backward [KKV25, LTZ25, CP25a, IRS25, YY25]. balance [GBFNR26]. balanced [GBFNR26]. ball [JL25]. Banach [OAO25, ZZM25]. bank [ZFCM25]. bar [ABC^{+}25]. barrier [HKY25]. based [ADJL25, AIF26, AS25, ANZ25, dAB25, CKC25, CLY25a, CLY25b, DP25, FA25, FLC^{+}25, FWL25, GHMN25, GCT25, GBFNR26, GLTW25, GHRK25, GD25, GLL^{+}25, HLF^{+}25, HNHR25, HLZ25, JFP25, JN25a, JJM25, JWS^{+}25, Kar25a, KS25d, LP25a, LZFP25, LWL^{+}25, LRG25, LG25, LTL^{+}25, LLZ25b, LWF25b, LJY25, LWZZ25, MS25a, MYL25, MWTW25, MHO25, MG25, OAO25, PLW^{+}25, RK25, SK25b, SSW^{+}25, TTZL25, TSL25, TYJH25, Wan25c, WCS25, WBH^{+}26, WRBT25, XJQ25, XNL25, XCGZ25, YLW^{+}25, YT25, You25, YPLH25, YYJZ25, ZuI25, ZLC^{+}25, ZZ25a, ZPGY25, ZDX25]. bases [HKM25]. Basis [GCN25, KCM25, MHO25, VKGM25, WYZ25, YLW^{+}25]. batch [SSLP25]. Bayesian [NAF25, SDS25, YXM25]. BDF2 [YTT25, ZWZJ25]. beam [CCT25]. behavior [EES25, KKV25]. behavioral [COPPS25]. Bénard [MHP25]. Benjamin [KM25]. Bessel [LC25c, ZC25b]. between [WZZ25, Xie25]. Bézier [AM25]. Bézier-Smart [AM25]. BFGS [SMDW25]. bi [LG25]. bi-objective [LG25]. biased [WSL25]. bicriteria [SSLP25]. bidiagonalization [LTZ25]. bifurcation [IHA25, Kar25b, MO25]. biharmonic [BCD25, LWH^{+}25]. Bilevel [CDPT25, CGuRZ25, EYJKL26]. bilinear [SZD25, XJQ25]. BINO [YXM25]. bioconvective [UX25]. bioheat [QSW25]. biological [FBM25]. Biot [BLC25, KL25]. biperiodic [DdF25]. biquaternion [CKC25]. bivariate [SAP^{+}25, TA25]. blades [BHV25]. Blind [XL25]. Block [GH25a, AK25b, BKST25, BS25c, CLJX25, KSC25, NZ25, SHL^{+}25, TGDC25, ZPGY25]. block-by-block [KSC25]. blood [LLZ25a]. blurred [JJM25]. Board [Ano25o, Ano25p, Ano25q, Ano25r, Ano25s, Ano25t, Ano25u, Ano25v, Ano25w, Ano25x, Ano25y, Ano25z, Ano25-27, Ano25-28, Ano25-29, Ano25-30, Ano25-31, Ano25-32, Ano25-33, Ano26b]. body [KR26]. Boltzmann [XCGZ25]. Bona [KM25]. book [DG25]. bordered [BS25c]. Bouligand [FLC^{+}25]. bound [ABP25, BPT25a, GWHY25, HLQF25, HGZL25, SWD25]. bound-preserving [ABP25, BPT25a, GWHY25]. boundaries [HU25]. Boundary [MMK25, ADLP25, BACF25, BCdR25, BCD25, CK25, GXC25, LFH25, LP25b, WRBT25, XC25, ZMWD25]. boundary-layer [BACF25]. boundary-precise [ZMWD25]. bounded [FKVZ25, HY25b]. Boundedness [EG25]. Bounds [CP25a, AXF26, BRPP25, KR25, SH25]. branch [HGZL25]. branch-reduction-bound [HGZL25]. Bregman [LY25]. Brinkman [BLC25, JLC25, LZZ25a, LWF25b]. Brownian [AAN25, ZZX25]. Broyden [TZ25b]. Broyden-like [TZ25b]. bubble [GS25]. bubble-enriched [GS25]. bus [SY25].

Cahn [GXC25, HKH^{+}25, HLQF25, PNL25, WWC25, XC25, ZSYZ25]. calculus [AXF26, HuR25, Öza25, Sak25]. Calibration [AARK25]. Can [NT25b]. candidate [LZFP25]. canonical [El 26]. capabilities [BDvdHV25]. capacity [FSC25]. capturing [WBH^{+}26]. Caputo [BY25, GCT25, HNHR25, JU25, VMRBG25]. cardiac [RNC25]. cardinality [WKQ25]. Carlo [DP25, SDS25]. cascade [BHV25]. case [Hag25]. category [HNHR25]. Cauchy [ANZ25, LF25]. causality [BDvdHV25]. causality-preservation [BDvdHV25]. cell [GH25b]. censored [WSL25]. censoring [NAF25]. centers [YLW^{+}25]. CG [AIF26, Che25b, HLZ25, MYL25, TTZL25]. CG-based [AIF26]. chain [EYJKL26, HBG25]. chains [BDFMM25]. chance [CLY25a]. channelized [HLC25]. channelized-media [HLC25]. chaos [IHA25, Kar25b, MO25, YAO^{+}25b]. chaotic [ADY25]. characterization [RV25]. Charged [Yin25]. Charged-Particle [Yin25]. Chebyshev [Moi25, NT25a, PM25b]. chemistry [GJWV25]. chemotaxis [QZG^{+}25]. chevron [KKV25]. class [Buo25, DS25, EYJKL26, NSW25, SK25a, TSM25, WX25a, WCS25, XHLH25, uRGYZ25]. classes [NRS25]. classical [JN25b]. climate [BPT^{+}25c, DCT^{+}25]. Closed [Wan25c]. Closed-form [Wan25c]. cluster [DCT^{+}25]. clustering [WCYL26, XQW25]. CoCoA [HGGPD^{+}25]. coding [LC25a, ZGWH25]. coefficient [QSW25]. coefficients [HHF25, WLL25, YY25]. coherence [LC25a, LLZ25b]. cold [MHWY25]. collected [DJ25]. Collocation [COPPS25, PT26, AD25, GHD25, HRST25, NKN25]. Color [CKC25, ZLLL25]. combination [YLW^{+}25]. combined [Che25a, HY25a]. Common [YK25, CGuRZ25, CXLP25, HU25, SY25, WCYL26]. communication [LSCZ25]. commutative [Hag25, YY25]. compact [BACF25, ZZJJ25]. comparative [TTZL25]. Comparing [VS25]. Comparison [BY25, Sak25, HRO^{+}25, Wan25c]. competing [WZTL25, YSP25]. complementarity [CLY25b, FWL25, GLL^{+}25, HV25, MG25, WCS25, XLLJ25]. complementary [WZTL25]. completion [AS25, ZLLL25]. complex [FLLZ25, FKVZ25, FA25, HCCC25, LCZ25, MG25, TSL25, YWS25]. complex-valued [MG25]. Complexity [ZS25, IHA25]. components [DJ25, GJ25]. Composite [CGuRZ25, DWZ25, HU25, LGDN25]. compressed [CLJX25]. compressibility [Gho25, YL25a]. compressible [ZZ25a]. compression [SHL^{+}25, ZPGY25]. compressive [LWLW25, ZHLD25]. Comput [MR25]. computation [CLZ25, GR25, Pho25, Sak25, WSKH26]. Computational [dAB25, UX25, AD25]. Computationally [YT25]. computations [ZWSC25]. computed [LSPV25]. Computing [JVW25, SHL^{+}25, Yao25a]. condition [GXC25, NGU25a, XC25]. Conditional [PJPW25, CPSW25, LLLL25]. conditioned [WLCM25]. conditions [ADLP25, EYJKL26, JU25, LP25b, LLZ25b, Ran25, SS25, Ton25, WRBT25, Xie25, ZYTH25, ZHLD25]. conduction [OTD^{+}25]. conductivity [Sai25]. cone [HGZL25, SWD25]. conformable [dAB25, GLTW25, MO25, BY25]. conforming [HLWZ25, HMZ25, WLZZ25]. conjugate [AMRS25, Don25, JWS^{+}25, ZHLD25]. connected [CS25]. connectivity [ZLY25]. conservation [ADJL25, FC25]. conservative [GD25, MMD25, Yin25]. conserved [ZC25a]. conserving [ZMG25]. consistency [MCL25]. consistent [HLF^{+}25, TGDC25, YL25a]. constant [BE26, TA25]. constants [LZZ25a]. constrained [AIF26, CLJX25, CLY25a, GS25, HHF25, JN25a, JJM25, LS25b, LWLW25, LTL^{+}25, PP25, WKQ25, YLW^{+}25, YYJZ25]. constraint [BAPPF25a, CGuRZ25, GHZ25, LYC25, TSL25, ZS25]. constraints [LG25, RNC25, TLSW25, Wan25a, WZ25a, ZLY25, ZFCM25]. Construction [HKM25, YC25, NRS25]. consumption [CNQY25]. contact [KR26, LYC25, XHLH25, ZBN^{+}25]. contaminated [SK25a]. continua [HLC25]. continuation [LR25]. continuity [Hun25, JJM25]. continuous [DLR25, MRRÁY25, YY25, ZCH^{+}25]. continuously [SQLZ25]. contractive [KIF^{+}25]. contractivity [YY25]. contrast [XGYH25]. Control [PSF25, AP25, GHZ25, GCT25, GLTW25, GS25, HU25, JWZZ25, Kar25b, LS25b, MZL25, PP25, RNC25, SC25, SKM25, SKM26, SWWY25, TSM25, VMRBG25, XDZ25, YS25, ZFCM25]. controllability [Sak25, SX25]. controlled [KIF^{+}25]. controller [JCCZ25]. convection [ABP25, CJ25, DZZ25, KS25b, QZG^{+}25, XCGZ25, YCLS25]. convection-diffusion [ABP25, CJ25, QZG^{+}25, YCLS25]. convection-dominated [DZZ25]. Convergence [DL25, IRS25, JL25, KR26, LP25b, LMZ25, LR25, Yin25, CLWY25, DS25, Hag25, HWZ25, LGDN25, Ngu25b, ZHF25]. Convergent [FLS25, DWZ25, SRMGMD25, ZZJJ25, uRGYZ25]. converging [KS25a]. Convex [AGGM25, AIF26, HHF25, LBM25, LTL^{+}25, TYJH25, XL25, YLW^{+}25, YPLH25, ZPGY25]. convex-constrained [AIF26]. convexifications [EYJKL26]. convexity [Wan25a, WZ25a]. Convolution [BHV25, XLZ25]. convolutional [ASRL25, LC25a]. cooperative [XGH25]. copositive [HLS25]. core [BPT25b, MF25]. correction [CGuRZ25, MRRÁY25]. corrector [AK25a]. correlation [PJPW25]. COS [ACO25]. cost [AAN25, FF25]. costs [PW25]. Coupled [LFH25, AHLS26, BLC25, CP25b, CG25, FSC25, JCCZ25, MAHL25, ZMG25]. coupling [WRBT25]. Cournot [Kar25b, Shi26]. covariance [BAPPF25a]. covariant [Ngu25b]. covering [Hag25]. CP [XNL25]. CP-decomposition [XNL25]. crack [KR26]. cracks [CLYY25]. Crank [HLQF25, LCZ25, WLL25]. credible [NAF25]. credit [HBG25]. criterion [YT25]. cross [XJQ25]. Crouzeix [LZZ25a]. crystal [LLZH25, WX25a]. Cubic [LWL^{+}25, GP25a, Ngu25b, RV25]. cumulant [CGR25]. cumulative [KZG25]. curl [WLZZ25]. curves [FKVZ25, LM25a, PSF25, RV25]. CVaR [XLLJ25]. cyclic [BS25c, BRPP25].

D

[ABC^{+}25, AKKM25, CJ25, LMZ25, MKE26]. damped [SMDW25]. damping [Don25, ZHF25]. dangling [SHL^{+}25]. Darcy [GH25b, HJ25, HYZ^{+}25, JLC25, You25]. Darcy-transport [JLC25]. Data [PDT^{+}25, CPSW25, CGR25, DJ25, DLR25, LW25, MXW^{+}25, SK25a, SQLZ25, WZTL25, WKQ25, WJ25, WCYL26, WSL25, XDZ25]. Data-driven [PDT^{+}25, CPSW25]. database [HRO^{+}25]. DEA [HC25, LG25]. deblurring [FF25]. December [Ano25c, Ano25i]. decentralized [ZLZ25]. decision [CNQY25, Kar25a, Wan25c]. decision-making [CNQY25, Wan25c]. decomposed [JWZZ25, LLLL25]. decomposition [CKC25, CWX25, CS25, JS25, KL25, LP25b, LC25c, MHP25, SGRS25, XNL25, XCGZ25]. decomposition-based [XCGZ25]. deconvolution [XL25]. Decoupled [LWZZ25, JWZZ25, PFYH25]. Deep [MR25, SSM25, ZLS25, MR23]. defined [ZMWD25]. degradation [Son25, YSP25]. degree [AM25]. DEIM [CWX25]. delay [GHD25, HRST25, KPTS25, LZZ^{+}25b, PDT^{+}25, PK25, XDZ25, Yu25]. delayed [FG25, GCT25]. delays [JCCZ25, PDT^{+}25, PT26, SC25]. delegation [Kar25b, Shi26]. delivery [BPT26]. delta [Moh25]. dendritic [WX25a]. dengue [LP25a]. denoising [FF25, VKGM25, ZLC^{+}25, ZGWH25]. density [DL25, HYZ^{+}25]. dependence [BFQ25, PJPW25, WZZ25]. dependent [ABP25, CNQY25, FLLZ25, OTD^{+}25, QSW25, QZG^{+}25, WZTL25, YSP25]. Derivative [AIF26, GCT25, Hag25, HNHR25, LP25a, LTL^{+}25, LJY25, Ngu25b]. Derivative-free [AIF26, Hag25, LTL^{+}25, LJY25]. derivatives [dAB25, EGM25, GLTW25, JU25, PS25]. descent [dAB25]. design [VKA25]. detecting [DCT^{+}25, YLC25]. detection [BPT^{+}25c, CGR25]. Detective [PM25a]. Determination [QWLF25]. determined [TST25]. determining [KKV25, QSW25]. Deterministic [GR25, TGDC25]. Developing [HLL25]. DFIG [SC25]. diagnostics [ZLC^{+}25]. diagonal [BS25c, DdF25]. Difference [AGGM25, CTZ25, FG25, FADA25, FC25, GWHY25, HKH^{+}25, JS25, LGZ25, MWTW25, Wan25b, ZYTH25, ZZJJ25]. differences [GNUV25]. differencing [AAKW25, HFSK25]. differentiable [SQLZ25]. differential [AFA25, AD25, FADA25, GNUV25, GBB25, Gho25, HWZ25, HNHR25, HB25, HY25b, KSC25, LCR25, LLG25, MKE26, PDT^{+}25, PT26, Ran25, RM25, SX25, SKM25, VMRBG25, XLZ25, Yu25, ZZX25]. diffuse [DGYH25]. Diffusion [EES25, ABP25, ACO25, AAKW25, BPT25a, CFCP26, CJ25, DH25, Fan25, FSC25, FF25, FGGdO25, FBM25, GP25b, KS25b, LWA25, MAHL25, PXZZ25, PY25, QZG^{+}25, VKGM25, WLL25, XGYH25, YCLS25]. diffusive [LM25b, Wan25b]. diffusive-viscous [LM25b]. diffusivity [LWA25]. digital [DP25]. Dilated [XLZ25]. Dimension [XJQ25, BACF25]. dimensional [AAKW25, Che25a, HLQF25, IHA25, KKHK25, MS25b, MO25, NKN25, Ran25, WLL25, WZ25b, WCYL26, ZMG25]. dimensions [ADJL25]. Dirac [CLYY25]. direct [LDZY25]. direction [WLL25]. discontinuous [BPT25a, BLC25, Che25a, DZZ25, HMZ25, JLC25, LW25, PM25a, QZG^{+}25, YK25, YCLS25, ZZ25a, ADPP25]. discovery [PDT^{+}25]. Discrete [GJ25, ADJL25, ACO25, AGGPDRC25, BRS25, CCLH25, DH25, FG25, GCT25, GH25a, Kar25b, LZFP25, LGZ25, LZZ^{+}25b, MHWY25, MWTW25, PDT^{+}25, Sak25, SC25, Vas25, WWC25]. discrete-time [AGGPDRC25, CCLH25, FG25, LGZ25, MHWY25]. discretization [Buo25, GH25b, HYZ^{+}25, KL25, LB25, SWWY25, SSM25, YWS25, ZLS25]. discretization-invariant [ZLS25]. Dispersion [SBAM25]. displacement [BFQ25, FLS25]. dissipation [LM25b]. dissipation-preserving [LM25b]. distance [Xie25]. Distributed [LSCZ25, HRST25, YS25]. distributed-order [HRST25]. distribution [CP25a, JFP25, KZG25, NBG25, NN25, WCYL26, YWY25]. distributions [ACO25, NRS25, NAF25, SK25a, SK25b, SAP^{+}25, TA25, WZTL25]. div [HLWZ25]. Dolichobrachistochrone [GBB25]. Domain [KS25a, KL25, LP25b, LLLL25, MHP25, XCGZ25]. domains [LM25b, YWW25, ZMG25]. dominated [DZZ25]. Double [CXLP25, AARK25, LTL^{+}25, YLC25]. double-step [LTL^{+}25]. Doubly [LSLL25]. Drazin [ZZM25]. drift [BPT25a]. drift-diffusion [BPT25a]. driven [CPSW25, CTZ25, PDT^{+}25, Ran25, ZZX25]. drug [BPT26]. DtN [LLG26]. dual [AK25a, CQ25, HKM25, JU25, LR25, YLW^{+}25, ZHF25, uRGYZ25, LCR25]. dual-inertial [uRGYZ25]. duality [JU25]. Dubrovsky [MS25b]. DuNN [LCR25]. duopoly [Kar25b]. Dynamic [CY25, IHA25, GBB25, GXC25, KR26, LP25a, VKA25, XC25]. dynamical [Buo25, ZHF25]. Dynamics [DH25, FG25, Shi26, Yin25, ADY25, CPSW25, LGZ25, MO25, YAO^{+}25b].

each [GJ25]. Earth [WBH^{+}26]. Editorial [Ano25o, Ano25p, Ano25q, Ano25r, Ano25s, Ano25t, Ano25u, Ano25v, Ano25w, Ano25x, Ano25y, Ano25z, Ano25-27, Ano25-28, Ano25-29, Ano25-30, Ano25-31, Ano25-32, Ano25-33, Ano26b]. Effectiveness [NGU25a]. Effects [LB25, Son25]. efficiency [BPT26, LG25, LZZ25a]. Efficient [CM25, FBZ25, HCCC25, LM25b, WZ25a, ZNL25, AK25a, CPSW25, CLJX25, CLY25a, CJ25, GHZ25, GHD25, GJWV25, HLQF25, JN25a, LC25b, LLZ25a, MMD25, SGRS25, SKM26, SWWY25, SSW^{+}25, WX25a, WKQ25, YL25a, YWW25, YT25, YYJZ25, ZLC^{+}25, ZHLD25]. efficiently [SHL^{+}25]. EIEQ [ZSYZ25]. eigenpairs [Yao25a]. eigensolutions [ABB25]. eigenvalue [GJWV25, HLS25, Sah25, TC25, WAXM25, ZDD25]. eigenvalue-eigenvector [ZDD25]. eigenvalues [DdF25, EG25, LS25a, QWLF25, Sah25, ZDD25]. eigenvector [ZDD25]. elastic [LWL^{+}25, LGDN25, ZBN^{+}25]. elasticity [HMZ25]. elastodynamic [FLS25]. electrical [ABC^{+}25, LW25]. electromagnetic [LFH25]. electrophysiology [RNC25]. Element [Vas25, XGYH25, ABP25, BS25a, CLYY25, CA25b, Che25a, CLWY25, CP25b, CH25, CG25, CKL^{+}25, DGMN25, DL25, FSC25, FFDL25, GHZ25, HLWZ25, HWZ25, HMZ25, LS25a, LCZ25, LYC25, LZZ25a, LLG26, LMZ25, LSPV25, OTD^{+}25, QZG^{+}25, Sak25, TZ25a, VS25, Wan25a, WLZZ25, WZ25a, WWC25, WCW25, XGX25, YWS25, YS25, ZBN^{+}25]. element-free [LCZ25]. elements [NT25b, SWWY25, VS25, ZZM25]. ellipse [SH25]. elliptic [ABB25, BS25a, CLYY25, GHZ25, GS25, LCR25, SWWY25, YWW25, ZH25]. Embedding [CKŚ25, WJ25]. Emden [SGRS25]. empirical [GH25a]. enabling [LU25]. Energy [ZMG25, DWZ25, LYC25, LLZH25, LWZZ25, MCL25, PFYH25, XC25, ZS25, ZSYZ25, ZC25a]. energy- [ZC25a]. Energy-conserving [ZMG25]. energy-optimized [MCL25]. energy-stable [LLZH25]. engineering [VKA25]. enhanced [AK25b, FGGdO25, HLF^{+}25]. Enhancing [CA25a]. enriched [DGMN25, GS25]. Entropy [WBH^{+}26, IHA25, LG25]. Entropy-based [WBH^{+}26]. environment [AAN25]. environments [Hun25]. epidemic [KS25a]. equality [YLW^{+}25, ZDD25]. equality-constrained [YLW^{+}25]. equally [ZYJ25]. equation [AFA25, AKKM25, BKST25, BPT25a, BACF25, BCD25, COPPS25, CP25b, CKL^{+}25, CTZ25, DZZ25, Fan25, FSC25, GXC25, GP25b, HKH^{+}25, HLQF25, JS25, KKHK25, KIF^{+}25, KM25, LCZ25, LWH^{+}25, LWA25, LLZH25, LLBX25, LLXB25, LF25, MS25b, NZ25, PNL25, QSW25, RK25, SSM25, WWC25, Wan25b, XC25, XGX25, YWS25, Yao25a, YT25, ZCH^{+}25, ZSYZ25, Zha25, ZZJJ25, ZZ25b, ZDX25]. equations [AIF26, ABP25, AHLS26, AD25, ADLP25, BPT25b, CLYY25, CAN25, CLWY25, CH25, CG25, DKM^{+}25, EMID26, FLLZ25, FFDL25, FADA25, FBM25, GNUV25, GHD25, GH25b, GD25, HWZ25, HNHR25, HB25, HLC25, HRST25, JJM25, KKV25, KSC25, LFH25, LC25b, LWLW25, LZZ25a, LM25b, LTL^{+}25, LCR25, LLG25, LJY25, MS25a, MYL25, MKE26, MO25, NSW25, PDT^{+}25, PXZZ25, PY25, PT26, PM25b, PK25, QZG^{+}25, QWZ25, Ran25, RM25, SX25, SQLZ25, SKM25, SGRS25, SBAM25, TZ25a, TZ25b, VMRBG25, WLL25, WZ25a, WZ25b, XCZ25, XLZ25, YTT25, YWW25, YCLS25, Yu25, YYJZ25, ZZX25, ZHX25]. equilibria [HY25b, XGH25]. equilibrium [CXLP25, PY25, YLC25, uRGYZ25]. equilibrium-diffusion [PY25]. equivalent [HGZL25, SS25]. Ericksen [ZWZJ25]. Erratum [MR25]. Error [DZZ25, DGYH25, GS25, SH25, WCW25, XCZ25, ZSYZ25, AMRS25, CA25a, CA25b, CP25b, CG25, JWZZ25, LTZ25, MWTW25, MR23, MR25, NBG25, PXZZ25, SWD25, TC25, WWC25, Xie25, YTT25, ZC25a]. estimate [DZZ25, JWZZ25]. estimates [AMRS25, CA25a, CA25b, CP25b, DS25, DGYH25, GS25, MWTW25, MR23, MR25, SWD25, WWC25, XCZ25, YTT25, ZC25a]. Estimating [BC25]. Estimation [BAPPF25a, BE26, FBM25, LSLL25, LLG25, PJPW25, SAP^{+}25]. estimations [PXZZ25, WSL25]. estimators [NN25]. Euler [Ran25, SKM25]. Eulerian [ADJL25, PY25]. evaluation [LC25c, SY25, Son25, TTZL25, XCZ25]. even [DLR25]. event [JCCZ25]. event-triggered [JCCZ25]. evolution [DCT^{+}25]. evolutionary [XHLH25]. Exact [PK25]. exactly [HHF25]. example [OTD^{+}25, ZFCM25]. exchange [Zho25]. exciting [ACO25]. Existence [MKE26, MAHL25, XGH25, TSM25]. exotic [WX25b]. expansion [CA25b, DKM^{+}25]. expansions [FLS25]. experiments [JN25a]. explicit [Hag25, KS25a, Ran25, ZSYZ25, ZWZJ25, ZZX25]. explicit-invariant [ZSYZ25]. Exploring [EES25]. exponent [IHA25, MS25a]. exponential [AARK25, AAKW25, CCLH25, CM25, DGMN25, HFSK25, JCCZ25, KS25d, NAF25, NPA^{+}25]. exponential-type [KS25d]. extended [SK25a]. extending [CLQY25]. extension [ACO25, CQ25, KS25c, MF25, SRMGMD25, ZLS25]. exterior [QWLF25, Sak25]. extragradient [CGuRZ25, CXLP25, OAO25, uRGYZ25]. extrapolation [LC25c].

Fabrizio [GCT25, JU25]. facility [LZFP25, MHWY25]. factor [Don25, LLG25, WCYL26]. factorization [ZNL25]. factors [LP25a]. Fading [BCD25]. failure [DJ25, GJ25]. failures [SDS25, YSP25]. family [uDJKA26, WZTL25]. Fast [EGM25, ZHF25, Fah25, GP25b, JWZZ25, LCZ25, ZGL^{+}25]. FastAGMGar [PLW^{+}25]. FDTD [HLL25]. FE [LDZY25, ZC25a]. feasibility [CXTL25]. feature [BPT^{+}25c, HLF^{+}25, LWH^{+}25, WBH^{+}26]. February [Ano25-34]. feedforward [YC25]. Feller [CdlHG25]. FEM [GHRK25, KR26]. FEM-based [GHRK25]. FEMs [PXZZ25]. Fickian [FGGdO25]. field [Gho25, LLZH25, LWZZ25, MT25, WX25a, Yin25, BAPPF25a]. fields [KPK25, BAPPF25b]. fifth [IHA25]. fifth-dimensional [IHA25]. filled [SQLZ25, SSW^{+}25]. filter [ANZ25]. finance [HBG25, SS25]. find [SSLP25]. finding [FA25]. Finite [BDFMM25, CTZ25, FFDL25, FC25, KPTS25, SZD25, VS25, Vas25, XGYH25, ABP25, BS25a, CLYY25, CA25b, CLZ25, Che25a, CH25, CKL^{+}25, DGMN25, DL25, FG25, FSC25, FADA25, GNUV25, GHZ25, GBFNR26, GWHY25, HKH^{+}25, HWZ25, HMZ25, LS25a, LYC25, LGZ25, LLG26, LMZ25, LWF25b, LSPV25, MWTW25, NT25b, PY25, QZG^{+}25, Sak25, SWWY25, TZ25a, Wan25a, WZ25a, WWC25, Wan25b, XGX25, XCGZ25, YWS25, YS25, ZBN^{+}25, Zha25]. Finite-frequency [SZD25]. finite-part [CLZ25]. finitely [PS25]. first [Buo25, CPSW25, CA25b, EMID26, XC25]. first-order [CPSW25, CA25b, XC25]. Fisher [KKHK25]. fitting [SQLZ25]. fixed [ADY25, CGuRZ25, CXLP25, LSCZ25, MZL25, NGU25a, YK25, uRGYZ25]. fixed-point [CGuRZ25, uRGYZ25]. fixed-time [MZL25]. flow [AHLS26, BHV25, Che25a, CLY25b, GCN25, Gho25, GWHY25, HYZ^{+}25, LLLL25, LWF25b, PFYH25, UX25, YL25a, You25, ZDX25]. flows [DGYH25, GHRK25, LLZ25a, XCGZ25, ZZ25a]. fluid [ADY25, BHV25, GHRK25]. flutter [VS25]. Fokker [RK25, ZCH^{+}25]. footrule [TST25]. Forchheimer [HYZ^{+}25]. forcing [WBH^{+}26]. forest [BPT^{+}25c]. form [Wan25c, XGH25]. formation [FBM25]. formula [CA25a, CA25b]. formulas [AXF26, SHK25]. formulation [FFDL25, ZuI25, ZZ25b]. Forward [YXM25, CP25a, IRS25]. four [CM25]. Fourier [LWH^{+}25]. Fourier-feature [LWH^{+}25]. fourth [AAKW25, HKH^{+}25, HLWZ25, uDJKA26, LMZ25, SGRS25]. fourth-order [AAKW25, HKH^{+}25, HLWZ25, uDJKA26, LMZ25, SGRS25]. Fowler [SGRS25]. FPMPE [ZWSC25]. fractal [KCM25]. Fractional [LBM25, ADY25, ASRL25, AXF26, ABH^{+}25, AAN25, BY25, dAB25, CFCP26, CH25, CG25, Fan25, GNUV25, GCT25, GLTW25, GP25b, HNHR25, HFSK25, HRST25, HuR25, JU25, JWZZ25, JN25b, KPTS25, Kar25b, KSC25, LP25a, LWA25, MMD25, MKE26, Öza25, RK25, RM25, SRMGMD25, SS25, SKM25, TZ25a, VMRBG25, WLL25, XDZ25, YAO^{+}25b, ZCH^{+}25, ZLC^{+}25, ZZX25]. fractional-order [dAB25, GLTW25, XDZ25, YAO^{+}25b, ZLC^{+}25]. fracture [Che25a]. fractured [ABH^{+}25]. fractures [LWF25b]. fragile [XDZ25]. frame [HLF^{+}25]. framework [GR25, Kar25a, KS25c]. Franklin [ZHX25]. Fréchet [NN25]. Fredholm [WZ25b]. Fredholm-Hammerstein [WZ25b]. free [AIF26, CK25, Hag25, HMZ25, LCZ25, LTL^{+}25, LJY25, SQLZ25, WJ25]. frequency [HB25, SZD25]. frequency-independent [HB25]. fully [MWTW25, PFYH25, WWC25]. fully-decoupled [PFYH25]. fully-discrete [WWC25]. function [CGR25, KZG25, MHO25, MPB25, Öza25, Pho25, SQLZ25, SSW^{+}25, WSL25, ZuI25, ZC25b]. function-based [ZuI25]. functional [COPPS25]. functionals [KKV25]. functions [CLZ25, DGMN25, EGM25, FLS25, HNHR25, KS25d, KCM25, LBM25, LU25, LSPV25, LC25c, MRRÁY25, Moh25, NT25a, OTD^{+}25, SHK25, SH25, SSW^{+}25, VKGM25, WZ25b, YC25]. fundamental [KC25]. fusion [HLF^{+}25]. Fuzzy [HC25, KS25d, AAN25, HJC26, Hun25, JN25a, SC25, Wan25c].

Galerkin

[BPT25a, BLC25, Che25a, DZZ25, HMZ25, JLC25, LCZ25, LW25, LM25b, LMZ25, PM25a, PM25b, QZG^{+}25, QWZ25, Wan25a, WZ25a, YCLS25, ZCH^{+}25, ZMG25, ZZ25a]. game [GBB25, Kar25b, LGZ25]. games [HY25b, Hun25, XGH25, ZYTH25]. gamma [Son25]. gap [QWLF25]. Gauss [dSMSdO^{+}25, SH25]. Gaussian [BAPPF25a, BAPPF25b, Ran25]. General [ZHX25, ASRL25, AGGPDRC25, CLWY25, HuR25, MRRÁY25, SSW^{+}25, WCW25, Yu25]. generalised [EG25]. generalization [MMD25, TYJH25]. Generalized [HFSK25, Vas25, dAB25, HY25b, HGZL25, Hun25, KM25, LC25b, LYC25, LLZ25b, MY25, MG25, MKE26, MS25b, MO25, NPA^{+}25, PM25a, Ran25, SKM26, SWD25, TSL25, WuR25, WZTL25, WAXM25, XGH25, XGX25, YWY25, ZGL^{+}25, ZZM25, ZPGY25, ZZJJ25, ZZ25b, XGYH25]. generated [PS25]. generating [CGR25]. generative [BDvdHV25]. Generic [MHO25]. genuine [CQ25]. Geometric [AAN25, CLY25a, DJ25]. Gibson [BFQ25]. Ginzburg [LCZ25]. Global [BRS25, SWD25, FA25, LSCZ25, SSW^{+}25]. globally [YY25]. GMsFEM [SSM25]. GMsHDG [PM25a]. Goal [BS25b]. Goal-oriented [BS25b]. gold [BPT26]. golden [LY25, OAO25]. Gompertz [JFP25]. Gordon [ZMG25]. governed [FBM25]. Gower [JMX25]. Graddiv [HLWZ25]. Graddiv-conforming [HLWZ25]. grade [UX25]. gradient [AMRS25, AP25, dAB25, CPSW25, CLQY25, Don25, GCT25, GLTW25, JWS^{+}25, ZHLD25]. gradient-based [GCT25, GLTW25]. Gramians [XJQ25]. graph [SNZ25, Vas25]. graphs [BAPPF25b, LSCZ25]. gravitational [Sai25]. Greedy [LLBX25, ABB25, LLXB25]. Greenwood [SAP^{+}25]. Gregory [ZBK25]. grey [BY25, ZDX25]. grid [CH25, GCN25, LRG25, TZ25a, WCW25, YL25b]. grids [EGM25, XCGZ25]. group [ZGWH25]. growth [LLZ25a, WX25a].

Hadamard [HNHR25]. Hamiltonian [BS25b]. Hammerstein [CAN25, WZ25b]. hand [Fah25]. Hankel [FBZ25]. harmonic [CS25, Sak25]. Hawkes [ACO25]. health [CNQY25, YLC25]. heat [ADY25, GD25, HYZ^{+}25, NKN25, OTD^{+}25]. heavy [JL25, SK25b]. heavy-ball [JL25]. heavy-tailed [SK25b]. Helmholtz [BCdR25, CTZ25]. Hemispheroidal [CS25]. hemivariational [XHLH25]. Henon [IHA25]. Hermite [ZMG25, LM25a, LM25b, WZ25b]. Hermitian [BRS25, GJWV25, Pho25, XNL25]. Hessenberg [BKST25]. Hestenes [Che25b]. Heston [ACO25, HKY25]. heterogeneity [LSLL25]. heterogeneous [ABH^{+}25, DJ25, GCN25, UX25]. heuristic [ZS25]. HH [AXF26, BRPP25]. hierarchical [Son25]. High [GWHY25, RK25, SBAM25, YCLS25, BAPPF25b, CWX25, DZZ25, HLQF25, MSR26, QSW25, WX25a, Wan25b, WCYL26, XGYH25]. high-accuracy [QSW25]. high-contrast [XGYH25]. high-dimensional [HLQF25, WCYL26]. High-order [GWHY25, RK25, SBAM25, YCLS25, WX25a, Wan25b]. Higher [LGDN25, LB25, XNL25]. Higher-order [LGDN25]. Highly [SGRS25, AKKM25, CLZ25, WX25b, ZuI25, ZC25b]. Hilbert [HKM25, SHK25, uRGYZ25]. Hilliard [GXC25, PNL25]. history [BFQ25]. HLLC [ZZ25a]. HLLC-based [ZZ25a]. HMC [SK25b]. hodograph [FKVZ25, PSF25, RV25]. homogeneous [UX25]. homogenization [AHLS26]. homotopy [EMID26, ZGL^{+}25, ZYJ25]. HOMS [LGDN25]. HOMS-PINN [LGDN25]. Hood [VS25]. horizon [LGZ25]. horizontal [MG25, WCS25]. HTL [YWS25]. HTL-WASI [YWS25]. human [LB25]. Hybrid [QZG^{+}25, Che25a, JWS^{+}25, dSMSdO^{+}25, XCZ25, YPLH25, Zho25]. hybrid-dimensional [Che25a]. hybridizable [PM25a]. hybridized [MYL25]. Hybridizing [VKA25]. hydrodynamics [Gho25, PY25]. hydrogels [DWZ25]. hyperbolic [El 26, GBFNR26, YWY25]. hypergeometric [Pho25]. hyperplane [WYZ25].

identification [CPSW25, LZZ^{+}25b, PJPW25]. identity [BRPP25]. IDFPM [JJM25]. IDFPM-based [JJM25]. II [NAF25]. III [CCT25]. ill [DKM^{+}25, FLC^{+}25, JL25, TLSW25, WLCM25, ZGL^{+}25]. ill-conditioned [WLCM25]. ill-posed [DKM^{+}25, FLC^{+}25, JL25, TLSW25, ZGL^{+}25]. image [CKC25, Che25b, CXLP25, HRO^{+}25, IRS25, JJM25, MYL25, TYJH25, VKGM25, ZLC^{+}25, ZLLL25, ZPGY25]. image-based [ZLC^{+}25]. images [HRO^{+}25]. imaging [MT25]. IMEX [MSR26, YTT25]. IMEX-AVF [MSR26]. immersed [ZH25]. impact [BPT^{+}25c, DG25]. impacts [DCT^{+}25, WBH^{+}26]. Impedance [LW25, ABC^{+}25]. imperfect [NSW25, SDS25]. implementation [BPT25b]. implementing [HGGPD^{+}25]. Implicit [MHP25, DGYH25, LS25b, WLL25, WCS25, XCGZ25, YCLS25, ZWZJ25]. implicit-explicit [ZWZJ25]. implicitly [ZMWD25]. implied [GHMN25]. importance [BPT^{+}25c]. imprecise [SW25, XQW25]. improve [HRO^{+}25]. Improved [WCS25, CXTL25, Don25, GP25b, HBG25, LC25a, OAO25, YLW^{+}25]. Improving [MCL25, Wan25c]. impulses [CCLH25]. inaccurate [WSKH26]. inclusion [HLS25, IRS25]. incomplete [MXW^{+}25]. incompressible [DL25, GHRK25, GWHY25, XCGZ25]. indefinite [TYJH25]. independent [HB25, LZZ25a]. index [JFP25, SAP^{+}25]. indicator [Kar25a]. indicator-based [Kar25a]. induced [CWX25, LWH^{+}25]. induction [PM25b]. inequalities [AXF26, Buo25, BRPP25, LBM25, LY25, TSM25, XHLH25]. inequality [OAO25, SS25]. Inertial [HHF25, CGuRZ25, CXLP25, IRS25, JCCZ25, LJY25, MYL25, YLC25, YYJZ25, ZHF25, uRGYZ25]. inertial-relaxed [MYL25]. inexact [CLY25b, ZLZ25]. inexactly [HHF25]. infeasible [AK25a]. inference [DJ25, KZG25, NBG25, NSW25, SDS25]. infinite [DLR25, JU25, LC25c]. Influence [Sai25]. information [BRPP25]. informed [LWH^{+}25, LGDN25, DMS^{+}25]. inhomogeneous [TC25]. initial [Ran25, YT25, ZGL^{+}25]. inner [GC25, GLL^{+}25, MHO25]. inner-outer [GLL^{+}25]. inpainting [HLF^{+}25]. input [HC25]. input-output [HC25]. Insights [CNQY25]. instabilities [FA25]. instability [SBAM25]. insurance [LK25, TA25]. integer [HY25a]. integer-valued [HY25a]. integrable [MO25]. Integral [MPB25, CAN25, EMID26, FLS25, GHZ25, GHD25, KIF^{+}25, LFH25, MS25a, WZ25b, ZZ25b, ZHX25]. integrals [CLZ25, LC25c, SHK25, WX25b, ZuI25, ZC25b]. integrated [Kar25a]. integrating [TYJH25]. integration [DP25, KS25c, MS25a, MRRÁY25, NPA^{+}25, NT25a, ZMWD25]. integrators [LWF25a, XGYH25]. integro [AFA25, AD25, FADA25, HNHR25, MKE26, PT26, RM25, SKM25, Yu25]. integro-differential [AFA25, AD25, FADA25, HNHR25, MKE26, PT26, RM25, SKM25, Yu25]. intensity [MSR26]. Inter [HLF^{+}25]. Inter-frame [HLF^{+}25]. interest [Zho25]. interface [BS25a, DGYH25, LP25b, WRBT25, ZH25]. interior [CJ25, GS25]. interlocking [HGGPD^{+}25]. interpolation [CA25a, CA25b, GH25a, LM25a, MHO25, OTD^{+}25, YL25b, ZYJ25]. interpolatory [CWX25]. interval [HJC26]. interval-valued [HJC26]. Intrusive [GCN25]. Invariant [QWZ25, PS25, ZLS25, ZSYZ25]. invariants [LWF25a]. inventory [CY25]. Inverse [Sah25, BCD25, CK25, CQ25, FBM25, GC25, GHRK25, HC25, LL25, MHO25, MXW^{+}25, MF25, NKN25, YXM25, ZZM25]. Inversion [YXM25]. inverted [WZTL25]. Invertible [LL25]. investigate [FA25]. investment [PW25, WZZ25]. investments [LK25]. invexity [SS25]. involving [HNHR25, JU25, LC25c, ZYTH25]. ions [PFYH25]. IPM [AK25a]. irregular [YWW25]. Ishikawa [LSCZ25]. Isogeometric [FLLZ25, LSPV25]. isotropic [TC25]. iteration [AK25a, GLL^{+}25, uDJKA26, NGU25a, TSL25, WCS25, WAXM25, ZGL^{+}25, ZWSC25, Ngu25b]. iteration-based [TSL25]. Iterative [dSMSdO^{+}25, BCdR25, GC25, GHZ25, LYC25, LTL^{+}25, MHO25, PNL25, WSKH26, ZLY25, ZPGY25].

J [MR25]. Jacobi [dSMSdO^{+}25]. January [Ano25d, Ano25j, Ano26a]. Jensen [BRPP25]. jobs [SSLP25, ZS25]. Joint [FBM25, ZGWH25, YWY25]. judgment [KZG25]. July [Ano25-35]. jump [ACO25, MSR26, ZYJ25]. jump-diffusion [ACO25]. jumps [AARK25]. June [Ano25-36].

Kac [DdF25]. Kaczmarz [LLBX25, LLXB25, NZ25, TGDC25, TLSW25, WYZ25]. Kawahara [SRMGMD25]. KdV [ZZJJ25, BPT25b]. Keller [FBM25]. kernel [HKM25, JN25b, KS25d, Öza25, ZuI25]. kernels [CAN25, Gho25, HuR25, SBAM25, ADPP25]. Kerr [CKL^{+}25]. Kerr-type [CKL^{+}25]. Key [NRS25]. kind [AD25, EMID26, GHD25, LF25]. kinematics [ZDX25]. kinetic [FA25]. Klein [ZMG25]. known [AK25a, FLS25]. Kolmogorov [KKHK25]. Konopelchenko [MS25b]. Kudryashov [MO25]. Kutta [Moi25, DZZ25, KS25c].

L [CWX25]. L-DEIM [CWX25]. Lagrange [WKQ25]. Lagrangian [ADJL25, HHF25, PY25, ZNL25]. Lanczos [LTZ25]. Landau [LCZ25]. Landweber [FLC^{+}25]. Lane [SGRS25]. language [CKŚ25]. Laplace [MR25, MR23, SK25a]. Laplacian [CK25]. large [AK25a, Che25b, DLR25, TGDC25]. large-scale [Che25b, TGDC25]. lateral [XGH25]. lattice [XCGZ25]. law [LLZ25a]. laws [ADJL25, FC25, GBFNR26]. layer [BACF25, ZBN^{+}25]. layers [ADY25, LU25]. LCP [SWD25]. leapfrog [SK25b]. learn [NT25b]. learned [ABC^{+}25, ZGL^{+}25]. learning [ADPP25, SSM25, VKA25]. Least [Ton25, AP25, TL25, WLCM25]. Lebesgue [BE26]. Leffler [JN25b, Moh25, Öza25]. Legendre [WuR25]. length [BC25, WSL25]. length-biased [WSL25]. Leslie [JMX25, ZWZJ25]. level [KC25]. Levenberg [TZ25b, TLSW25]. Levin [WX25b]. LFC [SC25]. liabilities [ZFCM25]. liability [LK25]. library [HGGPD^{+}25]. Lie [MS25b]. life [TA25, WSL25]. lifetime [JFP25, WZTL25]. lifetimes [DJ25]. LightGBM [HBG25]. Lightning [SNZ25]. like [CA25a, SS25, TZ25b]. likelihood [DJ25, LLG25]. limit [PY25]. limited [XJQ25]. line [CLYY25]. linear [AAKW25, CLJX25, DS25, DKM^{+}25, El 26, GHZ25, GD25, HV25, HB25, HGZL25, HMZ25, JL25, LLZH25, LZZ^{+}25b, MG25, NBG25, SX25, SZD25, TGDC25, VMRBG25, WSKH26, YL25a, Yu25, ZFCM25]. linearized [CG25, ZPGY25]. linearly [HHF25]. Liouville [GC25]. Lipschitz [CCGRR26, GR25, JJM25, LY25, YY25, ZZX25]. liquidity [LK25, LG25]. Lobatto [SH25]. Local [VKGM25, XK25, DZZ25, FC25, HLF^{+}25, HLC25, KL25, Xie25, XGH25, You25, YCLS25]. localized [HKM25]. locally [HJC26]. location [LZFP25]. locations [LZFP25]. locking [HMZ25]. locking-free [HMZ25]. Logarithmic [NN25, Wan25c]. logistic [AIF26, LJY25, YYJZ25]. long [LLZ25a, WKQ25]. long-only [WKQ25]. long-term [LLZ25a]. Longitudinal [SK25b]. loop [HY25b, LGZ25]. lot [SSLP25]. Low [FF25, BKST25, SZD25, YT25]. Low-cost [FF25]. low-rank [BKST25, SZD25, YT25]. lowest [CLWY25]. lowest-order [CLWY25]. LQ [LGZ25]. Lucas [LC25c, SKM26]. Luchko [HuR25]. Lyapunov [IHA25].

machine [SSLP25]. macromolecular [DWZ25]. magnetic [Yin25]. magnetohydrodynamic [DGYH25]. magnetohydrodynamics [DL25]. Mahony [KM25]. maintaining [LS25a]. making [CNQY25, Wan25c]. management [LK25, ZFCM25]. manifolds [JWS^{+}25, TTZL25]. Mann [LTL^{+}25, uRGYZ25]. Mann-type [uRGYZ25]. map [IHA25]. Mapped [WZ25b]. mapping [YWW25]. mappings [KIF^{+}25, NGU25a, YK25]. maps [Hun25, SC25]. March [Ano25e, Ano25k]. marker [GH25b]. marker-and-cell [GH25b]. markets [WZZ25]. Markov [BAPPF25a, BAPPF25b, BDFMM25]. Marquardt [TLSW25, TZ25b]. mass [GD25, ZC25a]. mass- [ZC25a]. MASSOR [MY25]. matched [OTD^{+}25]. matching [LLZ25b, SNZ25]. material [UX25]. materials [LGDN25]. Math [MR25]. mathematical [ITM25, SGRS25]. matrices [BS25c, CQ25, KR25, LL25, Sah25, Wan25c]. matrix [BAPPF25a, CKC25, DdF25, FWL25, GJWV25, JWZZ25, LC25b, MG25, NSW25, NZ25, Pho25, SKM26, WCS25, ZYTH25, ZNL25, ZDD25]. matrix-decomposed [JWZZ25]. maximization [CGR25]. Maximum [DJ25, LLG25, HLQF25, WZTL25, YCLS25]. maximum-principle-preserving [YCLS25]. maxmin [WZZ25]. Maxwell [ADLP25, CKL^{+}25, FLLZ25, WZ25a]. May [Ano25l, Ano25f]. Mean [Hag25, PW25, WZZ25]. Mean-square [Hag25]. mean-variance [PW25, WZZ25]. means [XQW25]. measles [ITM25]. measure [SX25]. measures [LG25]. mechanics [XHLH25]. mechanism [MHWY25]. media [ABH^{+}25, CKL^{+}25, FLLZ25, GCN25, GWHY25, HLC25, LWF25b, PM25a, XGYH25, YL25a, You25]. median [ANZ25]. medical [CXTL25, ZLC^{+}25]. medium [Sai25, TC25]. memristive [JCCZ25]. Mercer [AXF26]. mesh [KS25b, LMZ25, VN25, You25, DMS^{+}25]. Mesh-informed [DMS^{+}25]. meshes [Wan25a, WZ25a, WCW25]. Meshfree [Gho25, NPA^{+}25]. meshless [AD25, GNUV25]. metasurfaces [HLL25]. Method [Vas25, ADJL25, ABB25, AMRS25, AFA25, ABP25, AD25, ACO25, BKST25, BS25a, BPT25a, BCdR25, BLC25, dAB25, BCD25, COPPS25, CLYY25, CAN25, Che25a, CLWY25, CP25b, CG25, CJ25, DL25, DZZ25, EMID26, FADA25, FLC^{+}25, FWL25, GC25, GCN25, GHMN25, GHD25, GXC25, GS25, GLL^{+}25, Hag25, HKH^{+}25, HLWZ25, HJ25, HWZ25, HNHR25, HLC25, HLZ25, HLL25, HMZ25, JS25, JLC25, JL25, KC25, KSC25, KR26, LCZ25, LW25, LP25b, LWL^{+}25, LWH^{+}25, LM25a, LYC25, LWLW25, LWA25, LRG25, LLG26, LGDN25, LTL^{+}25, LMZ25, LCR25, LLXB25, LWF25b, LSPV25, LDZY25, LF25, MY25, MS25a, MHP25, MSR26, dSMSdO^{+}25, MCL25, MXW^{+}25, MR23, MR25, NKN25, NN25, NPA^{+}25, NDSZ25, OTD^{+}25, PM25a, PT26, PLW^{+}25, QSW25, QWZ25, QWLF25, SRMGMD25, SKM25, SKM26, SMDW25, TGDC25, TZ25a, TYJH25, TLSW25, VS25, VKGM25, WuR25, WLL25, WLZZ25, WLCM25, WYZ25]. method [WCW25, XNL25, XGX25, XCGZ25, XQW25, YL25a, YTT25, YWW25, Yao25a, YAO^{+}25b, ZLY25, ZMWD25, ZCH^{+}25, ZMG25, ZWZJ25, ZBN^{+}25, ZZ25a, Zha25, ZWSC25, ZNL25, ZHX25]. method-based [TYJH25]. Methods [XGYH25, AS25, AP25, CM25, CH25, CKL^{+}25, CTZ25, GH25a, GWHY25, HV25, KS25c, LM25b, LLBX25, LC25c, MWTW25, Moi25, NSW25, PM25b, PNL25, WX25b, Wan25a, WX25a, WZ25a, WCS25, Wan25b, Yin25, YCLS25, Yu25, ZYJ25, ZZX25, uRGYZ25]. metric [FKVZ25, KIF^{+}25, YK25]. MFPIO [ZWSC25]. microsphere [DWZ25]. Milstein [YY25]. Milstein-type [YY25]. minimization [LSPV25]. minimizers [Xie25]. minimizing [LYC25]. Minkowski [ZYTH25]. missing [WCYL26]. Mittag [JN25b, Moh25, Öza25]. Mittag-Leffler [JN25b, Moh25, Öza25]. Mixed [ZBN^{+}25, Che25a, HC25, KR26, LY25, QZG^{+}25]. Mixtures [WCYL26]. mKDV [BPT25b]. mobile [MZL25]. mobility [LWZZ25]. Model [WJ25, ABH^{+}25, ACO25, CCT25, CY25, CLY25a, DH25, DGYH25, EES25, FG25, GD25, HKY25, HJC26, HC25, ITM25, JMX25, Kar25b, KS25a, LS25a, LP25a, LRG25, LWF25b, LWZZ25, MMD25, QZG^{+}25, RNC25, SW25, Shi26, SZD25, WX25a, XLLJ25, YL25a, ZLZ25, ZWZJ25, ZLLL25, ZZ25b, ZDX25]. modeling [Gho25, MHWY25, MMK25, SK25a, Son25, YXM25]. modelling [BDvdHV25]. models [BY25, Che25a, CNQY25, GBFNR26, HFSK25, LG25, LDZY25, NBG25, SK25b, Vas25, WBH^{+}26, WRBT25, Zho25, DMS^{+}25]. modification [FKVZ25]. modifications [PSF25]. Modified [EMID26, BKST25, CKL^{+}25, DKM^{+}25, JMX25, KKHK25, KM25, LWLW25, QWLF25, SAP^{+}25, ZHLD25]. Modulus [MG25, FWL25, GLL^{+}25, WCS25]. Modulus-based [MG25, FWL25, GLL^{+}25, WCS25]. moisture [MMK25]. moment [Ton25]. moments [CP25a, NN25, YWY25]. momentum [LLXB25, TGDC25]. Mond [JU25]. monodomain [RNC25]. Monolithic [GH25b]. monotone [AIF26, CLQY25, IRS25, LJY25, MYL25, XLLJ25, YYJZ25]. monotonic [Moh25]. Monte [DP25, SDS25]. Montgomery [BRPP25]. Moore [BFQ25, CQ25]. MOR [ABB25]. Moreau [RNC25]. motion [AAN25, AP25, ZZX25]. MP [Yao25a]. MP-Newton [Yao25a]. Multi [KC25, BPT25b, CDPT25, HLC25, HC25, KSC25, LGDN25, LSCZ25, LLG25, MMD25, SHL^{+}25, SY25, XGH25, ZMG25, ZBN^{+}25]. multi-agent [LSCZ25]. multi-conservative [MMD25]. multi-continua [HLC25]. multi-core [BPT25b]. multi-dimensional [ZMG25]. multi-factor [LLG25]. multi-layer [ZBN^{+}25]. Multi-level [KC25]. multi-objective [HC25]. multi-period [CDPT25]. multi-player [XGH25]. multi-scale [LGDN25]. multi-state [SY25]. multi-step [SHL^{+}25]. multi-term [KSC25]. Multicontinuum [AHLS26]. multigrid [GH25b, PLW^{+}25]. multilevel [XGX25, You25]. multilinear [NDSZ25, ZWSC25]. multiobjective [CLQY25, Hun25]. multiple [BDFMM25, CXTL25, Fah25, HJC26, uDJKA26, YL25b]. multiple-sets [CXTL25]. multiplicative [HGZL25, LBM25]. Multiscale [ABH^{+}25, Vas25, VKGM25, XGYH25, HLC25, LLZ25a, LYC25, NPA^{+}25, PM25a, XLZ25]. multiterm [KPTS25]. Multivariate [Son25, CGR25, DCT^{+}25, Öza25, SK25b, WCYL26, Xie25, YWY25, ZDX25]. Munk [BACF25]. MVHS [HLZ25].

n [GJ25]. nanoparticles [BPT26]. narrowing [BPT26]. Nash [HY25b]. nature [PSF25]. Navier [FFDL25]. near [MT25]. near-field [MT25]. negative [WZZ25]. Nesterov [FLC^{+}25, SMDW25]. nets [DP25]. Network [LCR25, BHV25, GHMN25, LWH^{+}25, LGDN25, MZL25, Vas25, YC25, ZH25]. network-based [GHMN25]. networks [GHRK25, JCCZ25, NT25b, ZLS25, ADPP25]. Neural [ADPP25, LCR25, YXM25, BHV25, GHMN25, GHRK25, JCCZ25, LWH^{+}25, LGDN25, MZL25, NT25b, XLZ25, YC25, ZH25]. neurons [LU25]. neutral [AFA25, FADA25, KPTS25, Yu25]. Newton [CLY25b, HV25, KR26, LBM25, LC25b, WKQ25, Yao25a]. Newton-type [HV25, LBM25, LC25b]. Nicolson [HLQF25, LCZ25, WLL25]. no [LS25a]. nodally [ABP25, BPT25a]. noise [ANZ25, CPSW25, CTZ25, Hag25, MAHL25, YPLH25]. noises [Ran25]. Non [ASRL25, ADLP25, CFCP26, DG25, FGGdO25, GCN25, XDZ25, AAKW25, CLQY25, EGM25, FC25, Hag25, HLC25, LL25, LSLL25, LY25, RM25, SX25, TLSW25, XL25, YY25, YPLH25]. non-Archimedean [LL25]. non-commutative [Hag25]. non-convex [XL25, YPLH25]. Non-convolutional [ASRL25]. Non-Fickian [FGGdO25]. Non-fragile [XDZ25]. non-globally [YY25]. Non-Intrusive [GCN25]. non-linear [AAKW25, SX25]. non-Lipschitz [LY25]. non-local [HLC25]. non-monotone [CLQY25]. non-probability [LSLL25]. Non-reflecting [ADLP25]. non-smooth [RM25, TLSW25]. Non-standard [CFCP26]. non-uniform [EGM25]. Non-uniformly [DG25]. nonconforming [CLWY25, DGMN25]. nonconvex [CLJX25, JWS^{+}25, Xie25]. nonhomogeneous [JMX25]. nonlinear [AIF26, AP25, CPSW25, CAN25, CLWY25, CP25b, CH25, CG25, CKL^{+}25, DZZ25, FSC25, FF25, FWL25, GD25, GLL^{+}25, Hag25, HNHR25, HRST25, JN25a, JJM25, KL25, LCZ25, LWLW25, LRG25, LTL^{+}25, LLBX25, LLXB25, LJY25, MYL25, MMD25, MO25, PK25, SQLZ25, TZ25b, TLSW25, VKGM25, WLCM25, WCW25, XGX25, YTT25, Yao25a, Yu25, YYJZ25, ZMG25, ZGL^{+}25, ZC25a]. Nonlocal [HU25, ADJL25, LP25b, Sai25]. nonnegative [ZNL25]. Nonparametric [WSL25]. nonseparable [WLL25]. nonsmooth [ABC^{+}25, CLY25b, CLQY25, EYJKL26, FLC^{+}25]. nonstandard [FG25]. Nonsymmetric [NT25a]. nonzero [HY25b]. nonzero-sum [HY25b]. norm [LSPV25, Xie25, ZGWH25]. normalizing [LLLL25]. note [DdF25]. Novel [GHZ25, CA25a, FF25, GCT25, HWZ25, HC25, JCCZ25, MXW^{+}25, MZL25, SSW^{+}25, XGX25]. November [Ano25-37]. nuclear [ZGWH25]. nudging [DLR25]. numbers [GJ25]. Numerical [AD25, BFQ25, BPT25b, CK25, CdlHG25, FADA25, HYZ^{+}25, KM25, LWA25, NSW25, YY25, ZBK25, AXF26, AFA25, BPT26, CJ25, DWZ25, DGYH25, EGM25, GNUV25, HNHR25, JVW25, JN25a, JN25b, KKHK25, KR25, MMD25, MRRÁY25, QSW25, RK25, YK25, ZMWD25, ZSYZ25, ZHX25, ZFCM25]. Nuttall [MPB25]. Nyström [MS25a].

objective [HC25, LG25]. observations [BC25, SW25, XQW25]. October [Ano25g, Ano25m]. ODE [LS25b]. off [DdF25]. off-diagonal [DdF25]. offset [El 26]. one [BACF25, NKN25, SSW^{+}25, ZZX25]. one-dimensional [NKN25]. one-parameter [SSW^{+}25]. one-sided [ZZX25]. online [SSM25]. only [WKQ25]. Open [KPK25, CS25, HY25b, LGZ25]. open-loop [HY25b, LGZ25]. Open-source [KPK25]. operational [SKM26]. Operator [YXM25, CK25, DKM^{+}25, KR25, LWZZ25, XLZ25, ZLS25]. operators [ASRL25, CdlHG25, JN25b, LSCZ25, Öza25, YC25]. oppositional [VKA25]. Optical [MO25]. Optimal [CLWY25, CP25b, JU25, MT25, PXZZ25, SHK25, VMRBG25, WWC25, ZYTH25, ZC25a, GHZ25, GCT25, GLTW25, GS25, JWZZ25, uDJKA26, LS25b, LSPV25, PP25, RNC25, SSLP25, SKM25, SGRS25, SKM26, SWWY25, YS25, ZLC^{+}25, ZFCM25]. Optimality [EYJKL26]. Optimally [AKKM25]. Optimization [TL25, CLY25a, CLQY25, Che25b, CDPT25, EYJKL26, GCT25, GLTW25, HHF25, HLZ25, HC25, JWS^{+}25, KS25b, LR25, MHWY25, SS25, SSW^{+}25, SMDW25, TTZL25, TSL25, ZYTH25, ZLZ25]. optimization-by-continuation [LR25]. optimized [HBG25, MCL25]. Optimizer [VKA25]. Optimizing [BPT26, LK25]. Option [MSR26, AARK25, AAN25, MWTW25, YL25b]. options [ACO25, GHMN25, HKY25, ZZ25b, Zho25]. Order [DMS^{+}25, AK25a, AFA25, AAKW25, dAB25, Buo25, CM25, CWX25, CPSW25, CA25b, CLWY25, CP25b, DG25, DWZ25, DZZ25, FADA25, GBFNR26, GLTW25, GWHY25, GP25b, HKH^{+}25, HLWZ25, HGZL25, HRST25, JWZZ25, uDJKA26, KPTS25, Kar25b, LB25, LLZH25, LRG25, LGDN25, LMZ25, LWF25a, LWF25b, LC25c, Moi25, PFYH25, Ran25, RK25, SGRS25, SZD25, SBAM25, TZ25a, WX25a, WLL25, Wan25b, WRBT25, XC25, XNL25, XDZ25, YS25, YAO^{+}25b, YCLS25, ZYTH25, ZLC^{+}25, ZZJJ25, ZDX25, ZHX25]. order-book [DG25]. ordinary [HB25]. oriented [BS25b]. Orr [PM25b]. orthogonal [BDFMM25, EG25, JS25, LLZ25b]. Orthomin [DS25]. oscillating [BHV25, SHK25]. oscillation [HFSK25]. oscillators [WX25b]. oscillatory [AKKM25, CLZ25, LWF25a, WX25b, ZuI25, ZC25b]. osteoporosis [YLC25]. outcomes [LU25]. outer [GC25, GLL^{+}25, HGZL25]. Outlier [CGR25]. output [HC25]. overfitting [CCGRR26].

padded [BS25c]. PageRank

[SHL^{+}25, ZWSC25]. pairwise [Wan25c]. parabolic [CLWY25, CJ25, HLC25, NPA^{+}25, QWZ25, SBAM25, WCW25, XCZ25, YTT25, YS25]. Parallel [BS25c, dSMSdO^{+}25, SSLP25, XCZ25, XCGZ25, ZS25]. parallel-batch [SSLP25]. parallelization [BPT25b]. parameter [FBM25, NAF25, SQLZ25, SSW^{+}25]. parameter-free [SQLZ25]. parameterization [CS25]. parameterized [AK25b]. parameters [DJ25, DLR25, LWLW25]. parametric [ABB25, Hun25]. Pareto [HLS25, SSLP25]. part [AXF26, CLZ25]. partial [HWZ25, LS25a, LW25, LCR25, MKE26, XLZ25, ZDX25]. Particle [Yin25, Gho25]. path [AK25a, MZL25]. paths [BAPPF25b]. pathway [BPT^{+}25c]. pattern [FBM25, KKV25]. patterns [IHA25, JMX25]. PDE [MAHL25, PP25, ZLC^{+}25]. PDE-constrained [PP25]. PDEs [ABB25, FBM25]. Peaceman [TYJH25]. Peano [JN25b]. Peer [LS25b]. penalized [XLLJ25]. penalty [GS25, HJ25]. Penrose [CQ25]. pentadiagonal [Fah25]. performance [dAB25, JFP25, SY25]. period [CDPT25]. periodic [LZZ^{+}25b, SHK25]. perishable [CY25]. permutation [IHA25]. person [HY25b]. perspectives [TL25]. perturbation [EMID26, ZGL^{+}25]. perturbations [DdF25]. perturbed [CJ25, FBZ25, LMZ25]. Petrov [PM25b, ZCH^{+}25]. Petrovsky [KKHK25]. phase [Che25a, DGYH25, Gho25, GWHY25, LLZH25, LWZZ25, PFYH25, WX25a, YL25a, ZZ25a, ZFCM25]. phase-field [Gho25, LWZZ25, WX25a]. PHoPH [LM25a]. physical [KPK25]. physics [LWH^{+}25, LGDN25, MMD25, TSM25]. physics-informed [LGDN25]. Picard [LTL^{+}25, NGU25a]. piece [MRRÁY25]. piece-wise [MRRÁY25]. piecewise [PT26]. PINN [LGDN25]. Piskunov [KKHK25]. planar [FKVZ25, PSF25, RV25]. Planck [RK25, ZCH^{+}25]. plaque [LLZ25a]. plasma [FA25, MMD25]. platform [KPK25]. player [LGZ25, XGH25]. plus [FBZ25]. POD [GBFNR26, LRG25]. POD-based [GBFNR26]. point [ADY25, AK25b, CGuRZ25, CXLP25, MY25, NGU25a, PSF25, YK25, ZGL^{+}25, uRGYZ25]. points [CJ25, LSCZ25, WX25b]. pointwise [RNC25]. Poisson [YPLH25, ZGWH25]. poles [FA25]. polygonal [WCW25]. polyharmonic [KC25]. polyhedral [SWD25, WZ25a]. Polynomial [SSLP25, ZYJ25, FC25, Sah25]. polynomials [BDFMM25, EG25, FKVZ25, NKN25, PM25b]. polytopal [Wan25a]. Population [LZFP25]. Population-based [LZFP25]. poroelasticity [LRG25]. porous [GCN25, GWHY25, LWF25b, PM25a, YL25a, You25]. port [BS25b]. port-Hamiltonian [BS25b]. portfolio [CNQY25, CDPT25, LG25, SS25, WKQ25]. posed [DKM^{+}25, FLC^{+}25, JL25, TLSW25, ZGL^{+}25]. posedness [EES25, XHLH25]. positive [LBM25]. Positivity [KKHK25, BAPPF25a, DWZ25, HCCC25, PFYH25, PY25, ZZ25a]. positivity-preserving [DWZ25, PFYH25, PY25, ZZ25a]. possibly [NBG25]. post [HWZ25, KZG25]. post-processed [HWZ25]. posteriori [MWTW25, MR23, MR25, YTT25, You25]. posteriori-based [You25]. Powell [GP25a]. power [AAN25, CLY25b, LLZ25a, TSL25, XNL25]. power-law [LLZ25a]. PR [CLJX25]. practical [YLW^{+}25]. precise [ZMWD25]. Preconditioned [GHRK25]. preconditioner [AK25b, BRS25, YWS25]. preconditioning [GP25b, MHO25]. predator [CP25b, DH25, FG25, JMX25]. predator-prey [DH25]. predator-taxis [JMX25]. Prediction [SSM25, DMS^{+}25, HBG25, ZDX25]. predictions [LB25]. predictor [AK25a]. predictor-corrector [AK25a]. prescribed [LS25a, Sah25]. preservation [BDvdHV25, HLQF25, NRS25]. preserve [LWF25a, PSF25]. preserving [ABP25, BPT25a, DWZ25, GWHY25, HCCC25, KKHK25, KS25a, LM25b, PFYH25, PY25, YL25a, YCLS25, ZZ25a]. pressure [FFDL25]. prey [CP25b, DH25, FG25]. prey-predator [CP25b]. price [CY25, DG25]. Pricing [HKY25, Zho25, AARK25, AAN25, ACO25, CY25, GHMN25, MWTW25, MSR26, YL25b, ZZ25b]. primal [AK25a, LR25, YLW^{+}25, ZHF25]. primal-dual [AK25a, LR25, YLW^{+}25, ZHF25]. principle [HLQF25, YCLS25]. prior [XL25]. priori [MR23, MR25]. priors [ABC^{+}25]. probability [LSLL25]. problem [AS25, ABH^{+}25, BFQ25, BLC25, BCD25, CK25, CLY25b, FWL25, GS25, GJWV25, HJ25, JU25, JLC25, LLZ25a, LYC25, LLG26, LMZ25, LF25, MHP25, MXW^{+}25, MR23, MR25, OTD^{+}25, Sah25, SS25, TC25, WYZ25, XLLJ25, You25, ZPGY25, ZYJ25]. problems [AK25a, AK25b, AP25, BRS25, BS25a, BCdR25, CPSW25, CFCP26, CLJX25, CXTL25, CLY25a, CLY25b, CLQY25, Che25b, CXLP25, EYJKL26, FBM25, FLC^{+}25, GC25, GCN25, GHZ25, GCT25, GLTW25, GHRK25, GLL^{+}25, HLWZ25, HHF25, HV25, HLZ25, HMZ25, IRS25, JN25a, JJM25, JCCZ25, JWS^{+}25, JL25, KC25, KS25b, KR26, LC25b, LGDN25, LJY25, LY25, MY25, MMK25, MG25, NKN25, NPA^{+}25, OAO25, SQLZ25, SSLP25, SS25, SKM25, SKM26, SWWY25, TSL25, TYJH25, TLSW25, VKA25, WLZZ25, WCS25, WLCM25, WAXM25, WCW25, XNL25, YLC25, YXM25, YLW^{+}25, YS25, YYJZ25, ZYTH25, ZGL^{+}25, ZH25, ZLLL25, uRGYZ25]. procedure [WZTL25]. process [ACO25, CP25a, HY25a, LTL^{+}25, VKGM25]. processed [HWZ25]. processes [Son25]. product [LC25c, MS25a, MHO25, NT25a]. production [HC25]. products [CY25]. profit [Kar25b]. programming [AGGM25, CLY25a, CLY25b, GBB25, HGZL25, JN25a, JU25, LG25, TYJH25, Wan25c, XLLJ25, YLW^{+}25, ZPGY25]. programs [CLJX25]. Projection [GXC25, AIF26, CXTL25, FLC^{+}25, LTL^{+}25, LJY25, MYL25, WRBT25, YYJZ25]. projection-based [MYL25, WRBT25, YYJZ25]. projection-type [CXTL25]. Projective [KS25c]. projectors [XK25]. propagation [BCdR25, LWL^{+}25, Sak25]. proper [JS25]. properties [ASRL25]. property [QWZ25]. proportional [Kar25a]. proportional-XL [Kar25a]. proposed [FG25]. proximal [CLQY25, TYJH25, YLW^{+}25]. PRP [Don25]. pseudo [LJY25, YYJZ25]. pseudo-monotone [LJY25, YYJZ25]. pseudomonotone [CGuRZ25]. pseudostress [FFDL25]. pseudostress-pressure-velocity [FFDL25]. pulmonary [HRO^{+}25]. pure [XGH25]. pursuit [LLZ25b, WYZ25, YLW^{+}25]. Pythagorean [FKVZ25, PSF25, RV25]. Pythagorean-Hodograph [RV25, FKVZ25, PSF25].

quad [WLZZ25]. quad-curl [WLZZ25]. quadratic [JVW25, TSL25, Wan25c, Xie25, Zha25]. quadratization [MCL25, ZSYZ25]. quadrature [AXF26, SHK25, SH25, ZBK25, ZC25b]. quality [CY25]. quantification [LLLL25]. quantile [SK25b, WSL25]. quantile-based [SK25b]. quantiles [GR25]. quantum [GJWV25, Shi26]. quartic [TSL25]. quartic-quadratic [TSL25]. Quasi [DP25, HKM25, JCCZ25]. quasi-localized [HKM25]. Quasi-Monte [DP25]. quasi-synchronization [JCCZ25]. quaternion [GJWV25]. Quaternionic [El 26]. Queue [ACO25, AGGPDRC25, BC25]. quintic [PSF25, RV25]. Quotient [Ngu25b]. quotients [JS25].

Rachford [TYJH25]. radial

[KCM25, MHO25]. radiation [PY25]. radii [KR25]. radiotherapy [CXTL25]. radius [PNL25]. railway [HGGPD^{+}25]. Random [BAPPF25a, BAPPF25b, BPT^{+}25c, CXTL25, HC25, SWWY25, Son25]. Randomized [LLXB25, LWH^{+}25, NZ25]. range [JVW25]. rank [BKST25, SZD25, XNL25, YT25]. rank- [XNL25]. ranked [NSW25, WZTL25]. ranking [LZFP25, NSW25]. rate [DS25, YC25, Zho25]. rates [BC25, ZHF25]. ratio [LY25, OAO25]. Rational [GP25a]. rationality [HY25b]. Raviart [LZZ25a]. Rayleigh [MHP25, Ngu25b, TZ25a]. Reaction [EES25, AAKW25, CFCP26, FBM25, PXZZ25]. reaction-advection-diffusion [FBM25]. reaction-diffusion [AAKW25]. reactions [UX25]. reactive [LWF25b]. real [WKQ25]. recipe [ADPP25]. Reciprocal [TSM25]. reciprocity [QWLF25]. Reconstructing [DGMN25]. reconstruction [CK25]. reconstructions [ABC^{+}25]. record [JFP25]. recovering [Fan25]. Recovery [LLZ25b, CPSW25, IRS25, ZLLL25]. rectangular [KIF^{+}25]. recurrence [CP25a]. recursive [LZZ^{+}25b]. Reduced [DMS^{+}25, GCN25, CKC25, GBFNR26, GS25, LRG25, WRBT25]. reduced-order [GBFNR26, LRG25]. Reducing [CCGRR26]. reduction [ABH^{+}25, BS25c, HGZL25, MAHL25, SZD25, XJQ25, Xie25]. reductions [PK25]. reference [CY25]. refined [CA25b]. refinement [WSKH26, You25]. reflected [IRS25]. reflected-forward-backward [IRS25]. reflecting [ADLP25]. refuge [DH25]. regime [AKKM25, ZZ25b]. regime-switching [ZZ25b]. region [QWZ25, TST25, Xie25]. regional [HGZL25]. regions [HU25, NAF25, ZMWD25]. regression [AIF26, BPT^{+}25c, HJC26, KS25d, LC25b, LJY25, NBG25, SK25b, Ton25, YYJZ25]. regularisation [PP25]. Regularity [CLYY25]. regularization [ANZ25, BCD25, CCGRR26, Fan25, LF25, RNC25]. Regularized [SMDW25, CLJX25, WYZ25, ZLLL25, ZHF25]. regularizers [YPLH25]. reinsurance [Kar25a, PW25, WZZ25]. reinsurance-investment [WZZ25]. related [GHRK25]. relation [LLZ25a, PNL25]. relative [BAPPF25b, Kar25b]. Relaxation [LWF25a, HGZL25]. relaxed [LJY25, MYL25, YLC25]. relaxed-inertial [LJY25]. Reliability [SDS25, SY25, YSP25, LZZ25a, Son25]. Remora [VKA25]. removal [YPLH25]. renewal [CP25a]. reordering [SHL^{+}25]. reorthogonalization [LTZ25]. repair [MHWY25, SDS25]. repairable [MHWY25]. Representation [Pho25, AS25, CKŚ25, GP25a]. representations [MPB25]. Reproducing [ZuI25, Gho25, HKM25]. repulsive [CJ25]. residual [LSPV25, WSL25]. ResNet [CCGRR26]. resolution [HRO^{+}25, MMD25]. resource [LY25]. respect [MHO25]. respiratory [LB25]. restart [AIF26, JWS^{+}25]. restarting [Don25]. restoration [CKC25, Che25b, CXLP25, JJM25, MYL25, TYJH25]. restricted [HLZ25, WCYL26]. results [JU25, MKE26, RM25, TSM25]. retrial [AGGPDRC25, MHWY25]. Revisiting [TST25]. Richards [SSM25]. Richardson [NDSZ25]. ridge [LC25b]. Riemannian [JWS^{+}25, TTZL25]. Riesz [CdlHG25]. right [Fah25]. right-hand [Fah25]. risk [AAN25, CLY25a, DMS^{+}25, HBG25, LK25, WZZ25]. risks [WZTL25]. risky [LG25]. Ritz [MR25, MR23]. RKN [LWF25a]. RKN-type [LWF25a]. RLW [ZZJJ25]. RNN [JN25a]. robotic [AP25]. robots [MZL25]. Robust [NBG25, SC25, YS25, AK25b, CLQY25, CDPT25, uDJKA26, KS25d, LSLL25, PW25, ZLLL25]. rock [YL25a]. role [CNQY25, ITM25, LK25, WRBT25]. roots [uDJKA26]. Rosenau [SRMGMD25, ZZJJ25]. Rosenbrock [CM25]. Rössler [YAO^{+}25b]. rotating [Sai25]. ruled [GP25a]. rules [MRRÁY25, NT25a]. Runge [DZZ25, KS25c, Moi25]. rupture [DMS^{+}25].

S [SC25]. SAA [CLY25b]. SAA-based [CLY25b]. Sabin [GP25a]. saddle [AK25b, MY25]. Samarskii [VN25]. Samarskii-type [VN25]. sample [HLF^{+}25, XLLJ25]. sample-based [HLF^{+}25]. sampled [DG25, XDZ25]. sampled-data [XDZ25]. samples [LSLL25, WZTL25]. Sampling [PS25, NSW25, WZTL25, WRBT25]. Sard [JN25b]. SAT [KM25]. satisfying [NGU25a]. SAV [ZMG25]. SBP [KM25]. SBP-SAT [KM25]. SCAD [ZLLL25]. SCAD-Schatten- [ZLLL25]. scale [Che25b, LGDN25, TGDC25]. Scaled [ADPP25]. scaling [HHF25, ZS25, ZHF25]. scattering [LFH25]. Schatten [ZLLL25]. schedules [SSLP25]. scheduling [SSLP25, ZS25]. scheme [AAKW25, BACF25, DWZ25, DGYH25, FG25, GD25, HLQF25, HRST25, KKHK25, KS25a, LLZH25, MMD25, PFYH25, PY25, Ran25, XC25, ZSYZ25, ZMG25, ZNL25]. schemes [CFCP26, FC25, GHZ25, HCCC25, uDJKA26, LWZZ25, RK25, VN25, WWC25, XCZ25, YS25, YY25, ZC25a, Zha25, ZZJJ25]. Schrödinger [AKKM25, CH25, CG25, GD25, LF25, MO25, PK25, XCZ25, XGX25, Yao25a, ZC25a]. Schrödinger-type [ZC25a]. Schur [TA25]. Schur-constant [TA25]. Schwarz [MHP25, WRBT25]. Schwarz-based [WRBT25]. sciences [COPPS25]. Scott [VS25]. SDEs [Hag25, YY25]. SDO [AK25a]. search [HLF^{+}25, HBG25]. seasonal [LP25a]. Second [AFA25, AK25a, CP25b, DWZ25, FADA25, HGZL25, JS25, LLZH25, LWF25a, LWF25b, Ngu25b, PFYH25, Ran25, TZ25a, YS25, ZYTH25, ZDX25]. Second-order [AFA25, AK25a, CP25b, DWZ25, FADA25, HGZL25, LLZH25, LWF25a, LWF25b, PFYH25, Ran25, TZ25a, YS25, ZYTH25, ZDX25]. sections [OTD^{+}25]. seeking [LSCZ25]. Segel [FBM25]. segmental [BE26]. Seidel [dSMSdO^{+}25]. selection [CDPT25, WKQ25, WBH^{+}26, YT25]. Self [AS25, ACO25, OAO25]. self-adaptive [OAO25]. self-exciting [ACO25]. Semi [ADJL25, DGYH25, Hun25, JU25, KR26, MHO25, WWC25, ZDD25]. semi-continuity [Hun25]. Semi-discrete [ADJL25, WWC25]. semi-implicit [DGYH25]. semi-infinite [JU25]. semi-inner [MHO25]. semi-simple [ZDD25]. semi-smooth [KR26]. semidefinite [EYJKL26]. semilinear [QWZ25, TZ25a, Yao25a]. semismooth [CLY25b]. sense [VMRBG25]. sensing [CLJX25, LWLW25, ZHLD25]. sensor [MT25]. separable [TYJH25, WLCM25, ZPGY25]. September [Ano25-38]. septic [LM25a]. sequence [BPT25b]. sequentially [JWZZ25]. sequentially-decoupled [JWZZ25]. serial [PJPW25]. series [MPB25]. service [AGGPDRC25, BC25]. set [NSW25, WZTL25, ZYTH25]. set-valued [ZYTH25]. sets [CXTL25, WKQ25]. several [ADJL25]. SGP [LWLW25]. shape [CK25]. share [LY25]. sharing [SY25]. Sharp [AXF26, DS25]. sharpening [ANZ25]. shear [CCT25]. shift [AK25b, PS25]. shift-invariant [PS25]. shift-splitting [AK25b]. shifted [XNL25]. Shishkin [LMZ25, VN25]. sided [GP25b, ZZX25]. sides [Fah25]. sigmoidal [YC25]. sign [CKŚ25]. signal [JJM25]. silver [BPT26]. simple [ZDD25]. simplex [ZC25b]. simply [CS25]. simulated [DG25, HRO^{+}25]. simulating [LLZ25a]. simulation [HLL25, KPK25, KM25, SDS25]. simulations [BHV25]. single [SSLP25]. singular [CKC25, CAN25, DKM^{+}25, EMID26, MS25a, WZ25b]. singularly [CJ25, LMZ25]. SIS [KS25a]. sixth [ZZJJ25]. sixth-order [ZZJJ25]. skew [BRS25, WCYL26]. skew- [WCYL26]. skew-Hermitian [BRS25]. skewed [NBG25]. Smart [AM25]. SMEs [HBG25]. smooth [KR26, LWA25, RM25, TLSW25]. smoothed [Gho25]. Sobolev [DZZ25, EG25]. SOC [CLY25b]. soils [MMK25]. soliton [GD25]. solitons [MO25]. Solution [DKM^{+}25, KIF^{+}25, BRS25, FSC25, GNUV25, GBB25, Hun25, KSC25, LWA25, MAHL25, TSM25, Wan25c, ZHX25]. solution-based [Wan25c]. solutions [BPT25b, FADA25, KC25, MS25b, PK25, RM25, SHL^{+}25]. solve [CPSW25, GLTW25, HHF25, HNHR25, LWH^{+}25, SRMGMD25]. solver [Fah25, HB25]. solvers [MHO25]. Solving [Yu25, uRGYZ25, AK25a, AIF26, AD25, ADLP25, CGuRZ25, CK25, CXTL25, CLQY25, CXLP25, EMID26, GHD25, GHRK25, HV25, HGZL25, HRST25, IRS25, JN25a, KC25, LC25b, LGDN25, LTL^{+}25, LJY25, LF25, MYL25, MXW^{+}25, NSW25, NZ25, NDSZ25, RK25, SQLZ25, SS25, SGRS25, TGDC25, TZ25b, WLCM25, WAXM25, WSKH26, XNL25, Yao25a, ZSYZ25]. Some [TA25, ASRL25, Yin25, ZLS25]. Sommerfeld [PM25b]. sound [HU25]. source [Fan25, KPK25]. sources [AAN25, CLYY25]. Space [HLC25, BS25a, CH25, GNUV25, GP25b, HGZL25, LWA25, LWF25b, SHK25, WLL25, YK25]. space-fractional [CH25, GP25b, WLL25]. Space-time [HLC25, LWA25, LWF25b]. spaced [ZYJ25]. spaces [HKM25, KIF^{+}25, OAO25, PS25, uRGYZ25]. sparrow [HBG25]. sparse [CPSW25, CDPT25, JJM25, LC25a, Sah25, WYZ25, YL25b, ZYJ25, ZGWH25]. spatial [ABC^{+}25, Fan25]. Spatially [JMX25]. Spatio [DCT^{+}25, HLF^{+}25]. Spatio-temporal [DCT^{+}25, HLF^{+}25]. Spearman [TST25]. Spectral [CAN25, TC25, AP25, Che25b, GHD25, HLWZ25, HRST25, LWLW25, LM25b, MXW^{+}25, PM25b, PNL25, TTZL25, VKGM25, WLZZ25, ZMG25]. speed [ZS25]. SPH [SBAM25]. spherical [TSL25]. spillover [LS25a]. spline [LWL^{+}25, RV25, RK25, XK25]. splines [GP25a, XK25]. split [CGuRZ25, CXTL25, YLC25, YY25]. split-step [YY25]. splitable [SSLP25]. splitting [AK25b, AAKW25, BRS25, CLJX25, FWL25, IRS25, LC25b, LWZZ25, MG25, PNL25, TYJH25, WCS25, Yin25]. square [Hag25]. squares [AP25, TL25, Ton25, WLCM25]. Stability [FSC25, FADA25, HY25b, Kar25b, CCLH25, KPTS25, SX25, SAP^{+}25]. Stabilization [LSPV25, CCLH25]. stabilized [LZZ25a, Wan25a]. Stable [Wan25b, WSKH26, CKL^{+}25, DWZ25, KKHK25, LLZH25, LWZZ25, PFYH25, QSW25, SAP^{+}25, XC25, ZZJJ25]. Stackelberg [LGZ25]. standard [CFCP26, LJY25]. standby [MHWY25]. state [CNQY25, FBM25, GHZ25, GJ25, GS25, RNC25, SY25]. state-dependent [CNQY25]. state-parameter [FBM25]. states [GXC25]. stationary [AKKM25, FFDL25, HJ25]. statistic [SAP^{+}25]. Statistical [KZG25, TL25]. steady [GXC25, LB25]. Stefan [NKN25]. Steklov [QWLF25]. step [LTL^{+}25, LJY25, Moi25, SRMGMD25, SHL^{+}25, YY25]. stepping [WX25a]. steps [PXZZ25, ZWZJ25]. stepsize [LS25b]. Stiefel [Che25b]. STLN [YT25]. STLN-based [YT25]. Stochastic [MHWY25, AARK25, CCLH25, CLY25b, CTZ25, HKY25, HCCC25, JL25, KS25a, LGZ25, LY25, SMDW25, XLLJ25, ZLZ25, ZZX25, Zho25]. Stokes [MR25, BRS25, FFDL25, GH25b, HJ25, KL25, MR23, TZ25a, Zha25]. Strang [Yin25]. strategic [Shi26]. strategies [PW25, WRBT25, XGH25]. strategy [AIF26, CGuRZ25, Don25, HGZL25, JWS^{+}25, LGZ25, WZZ25, XGH25, ZLC^{+}25]. stratification [KZG25]. stripe [LC25a]. strong [Yin25]. strongly [KS25a, SHK25, uRGYZ25]. structural [VKA25, ZLY25]. structure [CLY25a, YL25a]. structure-preserving [YL25a]. structured [AP25, LS25a, Sah25, YT25]. study [BPT26, PM25b, UX25]. Sturm [GC25]. sub [LWA25]. sub-diffusion [LWA25]. subdifferential [ZYTH25]. subdiffusion [YXM25]. subgradient [CXLP25, OAO25, uRGYZ25]. Subsampling [TL25]. subspace [WAXM25]. Sufficient [Xie25]. suitable [XK25]. sum [HY25b, ZYTH25, ZZM25]. superconvergence [PXZZ25]. Superconvergent [RM25, CAN25]. supply [EYJKL26, HBG25]. support [XGH25]. sure [CCLH25]. surfaces [AM25, CS25, GP25a]. surfactant [LWZZ25]. surrogate [WYZ25]. suspenders [CCT25]. SVM [AGGM25]. switching [SDS25, ZZ25b]. Sylvester [BKST25, DdF25]. Symmetric [LS25a, SAP^{+}25, ZLLL25, ZNL25]. symmetrical [FBZ25]. symmetry [MS25b]. synchronization [JCCZ25]. system [Buo25, DJ25, DL25, DWZ25, GJ25, KL25, LB25, MAHL25, PFYH25, SRMGMD25, SY25, WBH^{+}26, YSP25, ZMG25, ZC25a, ZHF25]. systems [AAKW25, BS25b, CCLH25, CJ25, DS25, FBZ25, Fah25, HGGPD^{+}25, HB25, HCCC25, KPTS25, LCR25, LWF25a, LZZ^{+}25b, MHWY25, NRS25, NDSZ25, SZD25, TGDC25, Wan25c, WSKH26, XJQ25, XDZ25, ZLY25, ZBN^{+}25].

T [SC25, XK25]. T-splines [XK25]. Tail [YWY25]. tailed [SK25b]. Takens [WJ25]. taxis [JMX25]. Taylor [CA25a, VS25]. Taylor-like [CA25a]. technique [AD25, FF25, LWZZ25, MCL25, MO25, OAO25, PNL25, SWWY25, TYJH25, TZ25b]. techniques [HLF^{+}25, HJC26, HRO^{+}25, TTZL25]. temperature [FGGdO25]. tempered [HuR25]. Temporal [TZ25a, DCT^{+}25, HLF^{+}25, LLZ25a]. tensile [SBAM25]. tensor [AS25, BKST25, CWX25, NDSZ25, TL25, XL25, XLLJ25, XGH25, ZLLL25]. tensors [HLS25, NDSZ25]. term [Fan25, FSC25, KSC25, LLZ25a]. terms [FLS25]. ternary [DWZ25]. test [LSPV25]. Testing [SAP^{+}25, Kar25b]. Tetrahedral [Zha25]. Thailand [LP25a]. their [ASRL25, WZ25b]. theorem [JN25b]. theorems [HLS25, XGH25, YK25]. theoretical [LC25a]. theory [LG25]. thermal [BPT26, BFQ25, GD25, Sai25, XCGZ25]. thermodynamical [YL25a]. thermoelastic [BFQ25, Sai25]. thermoelasticity [CCT25]. thermoporoelasticity [ABH^{+}25]. Third [Moi25, AD25, GHD25, UX25]. third-grade [UX25]. Thompson [BFQ25]. three [AK25b, Gho25, GJ25, KKHK25, LC25c, ZPGY25]. three-block [ZPGY25]. three-by-three [AK25b]. three-dimensional [KKHK25]. three-phase [Gho25]. three-state [GJ25]. Tikhonov [ZHF25]. till [CM25, DJ25]. time [ABH^{+}25, ABP25, AAKW25, AGGPDRC25, BS25b, CCLH25, CFCP26, CG25, FG25, FLLZ25, Fan25, GCT25, GJ25, HFSK25, HLC25, JWZZ25, JCCZ25, KPTS25, KM25, KS25c, KL25, LWA25, LGZ25, LSCZ25, LWF25b, LZZ^{+}25b, MHWY25, MZL25, PXZZ25, QSW25, QZG^{+}25, RK25, Sak25, SSLP25, SC25, SW25, TZ25a, WX25a, XJQ25, XGYH25, XDZ25, ZCH^{+}25, ZMG25, ZWZJ25, ZHF25]. time-delay [LZZ^{+}25b]. time-delayed [GCT25]. time-dependent [ABP25, FLLZ25, QSW25, QZG^{+}25]. time-fractional [CFCP26, CG25, Fan25, JWZZ25, RK25, TZ25a, ZCH^{+}25]. time-harmonic [Sak25]. time-limited [XJQ25]. time-stepping [WX25a]. time-varying [KPTS25, LSCZ25, SC25, SW25, XDZ25]. timer [HKY25]. times [AGGPDRC25, CP25a]. Timestep [OTD^{+}25]. Timestep-dependent [OTD^{+}25]. Toeplitz [FBZ25, Fah25]. Toeplitz-plus-Hankel [FBZ25]. tomography [ABC^{+}25, LW25]. Total [Fan25, BAPPF25a, XL25]. tracking [ABB25, DCT^{+}25, MZL25]. traffic [ZDX25]. trajectory [ZHF25]. transaction [AAN25, PW25]. transfer [ADY25, GD25, HYZ^{+}25]. transform [El 26, WuR25]. transformed [BS25a, CG25]. transmission [ITM25, LP25a, LLG26, TC25]. transport [AHLS26, GCN25, JLC25, LWF25b, MMK25, PFYH25]. trapezoidal [Wan25c]. treatment [ITM25]. triangles [ZBK25]. triangular [WLZZ25]. triggered [JCCZ25]. triplets [LS25b]. truncated [AKKM25, DKM^{+}25, Hag25]. truss [CLY25a]. trust [Xie25]. Tseng [CGuRZ25]. Tseng-type [CGuRZ25]. turbulent [LB25]. Turkey [ITM25]. turning [CJ25, WX25b]. Two [CH25, JWS^{+}25, LRG25, LGZ25, MYL25, ZZJJ25, ZZX25, CLJX25, Che25a, DGYH25, GCN25, GWHY25, GP25b, KR26, LWLW25, LWF25a, LJY25, Moi25, NAF25, PFYH25, SRMGMD25, TZ25a, WLL25, WZ25b, WZZ25, WCW25, YL25a, ZGL^{+}25, ZZM25, ZZ25a]. two-block [CLJX25]. two-body [KR26]. two-dimensional [WLL25, WZ25b]. Two-grid [CH25, LRG25, GCN25, TZ25a, WCW25]. two-parameter [NAF25]. two-phase [Che25a, DGYH25, GWHY25, PFYH25, YL25a, ZZ25a]. Two-player [LGZ25]. two-point [ZGL^{+}25]. two-sided [GP25b]. two-step [LJY25, Moi25, SRMGMD25]. type [BFQ25, BRPP25, CGuRZ25, CCT25, CLJX25, CXTL25, CKL^{+}25, DdF25, FBM25, HV25, KS25d, LBM25, LC25b, LWF25a, NAF25, SWWY25, SWD25, TZ25b, VN25, XC25, YLC25, YY25, ZLLL25, uRGYZ25, ZC25a]. type-II [NAF25].

ultrasound [ZLC^{+}25]. unbounded [LM25b, ZMG25]. Uncertain [LG25, SW25, XQW25, CLY25a, LLG25, XDZ25, ZYTH25]. uncertainty [EYJKL26, GHMN25, LLLL25, LG25]. Unconditional [CG25, CKL^{+}25]. unconditionally [KKHK25, LLZH25]. unconstrained [Che25b, HLZ25, SSW^{+}25]. undamped [LDZY25, ZLY25]. underlying [YL25b]. unequal [WZTL25]. uniform [EGM25, MT25]. uniformly [DG25]. uniqueness [FSC25]. univariate [EGM25, SAP^{+}25, SSW^{+}25]. unreliable [MHWY25]. unstructured [XCGZ25]. Unveiling [YAO^{+}25b]. update [TZ25b]. updates [AK25a]. updating [LS25a, LDZY25, SHL^{+}25, ZLY25]. upper [KR25]. Using [BPT25b, AD25, CA25a, FF25, GNUV25, Gho25, HRO^{+}25, HC25, HGZL25, KM25, KZG25, LC25a, LTL^{+}25, MSR26, MO25, SK25b, Son25, SSM25, VKGM25, Wan25c, WCYL26, XLLJ25, ZLLL25]. utility [Wan25c]. utilizing [PM25b].

vaccination [ITM25]. value [BCdR25, BCD25, CKC25, CLY25a, DKM^{+}25, LC25b, LL25, MMK25, Ran25, ZGL^{+}25]. value-at-risk [CLY25a]. valued [HJC26, HY25a, MG25, WX25b, ZYTH25]. values [JFP25, WCYL26]. vanishing [ZHF25]. Variable [LS25b, DL25, GP25b, HYZ^{+}25, JWZZ25, LWF25b, MS25a, PXZZ25, Sai25, WLL25, ZWZJ25]. variable-order [GP25b, JWZZ25, WLL25]. Variable-stepsize [LS25b]. Variably [ADPP25]. variance [PW25, WZZ25]. variant [HLZ25]. variation [Fan25, XL25]. variational [Buo25, JU25, LY25, OAO25, SS25, TSM25, XHLH25]. variational-hemivariational [XHLH25]. variational-like [SS25]. various [MAHL25]. varying [KPTS25, LWA25, LSCZ25, SC25, SW25, XDZ25]. vector [HLZ25, SS25, TSM25, WX25b]. vector-valued [WX25b]. vectors [Wan25c]. velocity [FFDL25]. vertical [FWL25, HV25]. via [ADPP25, BRPP25, CGR25, EYJKL26, EG25, GBB25, KSC25, KIF^{+}25, MMD25, MS25b, QWLF25, SAP^{+}25, SZD25, SSW^{+}25, XL25, XGH25, Yu25, ZYTH25, ZC25b]. vibration [WAXM25]. vibroacoustic [LDZY25]. video [HLF^{+}25, ZLLL25]. VIP [CGuRZ25]. virtual [CLWY25, CP25b, CG25, WCW25]. viscoelastic [LWL^{+}25]. Viscosity [LZZ25a, YLC25]. Viscosity-independent [LZZ25a]. viscosity-type [YLC25]. viscous [LM25b, Wan25b]. Vogelius [VS25]. volatility [AARK25, GHMN25, HKY25, Zho25]. volatility-barrier [HKY25]. Volterra [AFA25, AD25, CAN25, FADA25, GHD25, MS25a, PT26, RM25]. volume [GBFNR26, LFH25, LWF25b, PY25, XCGZ25, Zha25]. VSBDF2 [YS25].

WASI [YWS25]. wave [BCdR25, JS25, LWL^{+}25, LM25b, Sak25, Wan25b, XCZ25]. waveguides [ADLP25]. wavelet [ANZ25, SKM26, ZHX25]. wavelets [SKM25]. Weak [SHL^{+}25, ADJL25, LMZ25, QWZ25, Ton25, Wan25a, WZ25a]. weakly [CAN25, KIF^{+}25, MS25a, WZ25b, ZZ25a]. weight [NT25a]. weighted [BRPP25, CP25b, DGMN25, HJC26, ZGWH25]. Weir [JU25]. Well [GBFNR26, XHLH25, AK25a, EES25]. Well-balanced [GBFNR26]. well-known [AK25a]. Well-posedness [XHLH25, EES25]. wheeled [MZL25]. while [LS25a]. white [CTZ25, Ran25]. wise [MRRÁY25]. within [Don25]. without [JJM25, LU25, Wan25a, WZ25a]. WKB [AKKM25]. Wolfe [ZHLD25]. words [CKŚ25].

XL [Kar25a].

Yosida [RNC25].

zero [ZYTH25]. zero-sum [ZYTH25]. zeroing [MZL25]. zeros [EG25, FA25].

References

Asante-Asamani:2025:FOE

[AAKW25] E. O. Asante-Asamani, A. Kleefeld, and B. A. Wade. A fourthorder exponential time differencing scheme with dimensional splitting for nonlinear reaction-diffusion systems. Journal of Computational and Applied Mathematics, 465(??):??, September 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725000834.

Alsenafi:2025:GAP

[AAN25] Abdulaziz Alsenafi, Fares Alazemi, and Alireza Najafi. Geometric Asian power option pricing with transaction cost under the geometric fractional Brownian motion with w sources of risk in fuzzy environment. Journal of Computational and Applied Mathematics, 453(??):??, January 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704272400414X.

Agazzotti:2025:COP

[AARK25] Gaetano Agazzotti, Jean-Philippe Aguilar, Claudio Aglieri Rinella, and Justin Lars Kirkby. Calibration and option pricing with stochastic volatility and double exponential jumps. Journal of Computational and Applied Mathematics, 465(??):??, September 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725000780.

Alghamdi:2025:GMM

[ABB25] Moataz Alghamdi, Daniele Boffi, and Francesca Bonizzoni. A greedy MOR method for the tracking of eigensolutions to parametric elliptic PDEs. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724005193.

Alsaker:2025:BRN

[ABC^{+}25] Melody Alsaker, Benjamin Bladow, Scott E. Campbell, Nicholas Linthacum, Thomas M. McKenzie, Jennifer L. Mueller, and Talles Batista Rattis Santos. Dbar reconstructions with nonsmooth learned spatial priors in 2D electrical impedance tomography. Journal of Computational and Applied Mathematics, 463(??):??, August 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725000275.

Alikhanov:2025:MMR

[ABH^{+}25] Anatoly Alikhanov, Huiran Bai, Jian Huang, Aleksei Tyrylgin, and Yin Yang. Multiscale model reduction for the time fractional thermoporoelasticity problem in fractured and heterogeneous media. Journal of Computational and Applied Mathematics, 455(??):??, February 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724004060.

Amiri:2025:NBP

[ABP25] Abdolreza Amiri, Gabriel R. Barrenechea, and Tristan Pryer. A nodally boundpreserving finite element method for time-dependent convection-diffusion equations. Journal of Computational and Applied Mathematics, 470(??):??, December 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725002055.

Arias:2025:HQH

[ACO25] Luis A. Souto Arias, Pasquale Cirillo, and Cornelis W. Oosterlee. The Heston– Queue–Hawkes process: a new self-exciting jump-diffusion model for options pricing, and an extension of the COS method for discrete distributions. Journal of Computational and Applied Mathematics, 454(??):??, January 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724004266.

Aourir:2025:NCT

[AD25] E. Aourir and H. Laeli Dastjerdi. Numerical computational technique for solving Volterra integro-differential equations of the third kind using meshless collocation method. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724005429.

Abreu:2025:SDL

[ADJL25] Eduardo Abreu, Richard De la cruz, Juan Juajibioy, and Wanderson Lambert. Semidiscrete Lagrangian–Eulerian approach based on the weak asymptotic method for nonlocal conservation laws in several dimensions. Journal of Computational and Applied Mathematics, 458 (??):??, April 2025. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724005739.

Aregba-Driollet:2025:NRB

[ADLP25] Denise Aregba-Driollet, Patrick Lacoste, and Corentin Prigent. Non-reflecting boundary conditions for solving Maxwell’s equations in waveguides. Journal of Computational and Applied Mathematics, 466(??):??, October 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704272500113X.

Audone:2025:RLV

[ADPP25] G. Audone, F. Della Santa, E. Perracchione, and S. Pieraccini. A recipe for learning Variably Scaled Kernels via Discontinuous Neural Networks. Journal of Computational and Applied Mathematics, 469(??):??, December 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042725001670.

Ahmad:2025:FPA

[ADY25] Haroon Ahmad, Fahim Ud Din, and Mudasir Younis. A fixed point analysis of fractional dynamics of heat transfer in chaotic fluid layers. Journal of Computational and Applied Mathematics, 453(??):??, January 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042724003935.

Amirali:2025:SON

[AFA25] Ilhame Amirali, Burcu Fedakar, and Gabil M. Amiraliyev. Second-order numerical method for a neutral Volterra integrodifferential equation. Journal of Computational and Applied Mathematics, 453 (??):??, January 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724004096.

Astorino:2025:DCP

[AGGM25] Annabella Astorino, Manlio Gaudioso, Enrico Gorgone, and Benedetto Manca. Difference of convex programming in adversarial SVM. Journal of Computational and Applied Mathematics, 454 (??):??, January 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724004503.

Atencia:2025:DTQ

[AGGPDRC25] Iván Atencia, José Luis Galán-García, Yolanda Padilla-Domínguez, and Pedro Rodríguez-Cielos. A discrete-time queue with service time adjustments and general retrial times. Journal of Computational and Applied Mathematics, 467(??):??, October 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725001207.

Ammosov:2026:MHC

[AHLS26] Dmitry Ammosov, Jian Huang, Wing Tat Leung, and Buzheng Shan. Multicontinuum homogenization for coupled flow and transport equations. Journal of Computational and Applied Mathematics, 471(??):??, January 1, 2026. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704272500250X.

Abubakar:2026:DFP

[AIF26] Auwal Bala Abubakar, Abdulkarim Hassan Ibrahim, and Yuming Feng. Derivativefree projection CG-based algorithm with restart strategy for solving convexconstrained nonlinear monotone equations and its application to logistic regression. Journal of Computational and Applied Mathematics, 471(??):??, January 1, 2026. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725001906.

Abedi:2025:ESO

[AK25a] Hadis Abedi and Behrouz Kheirfam. An efficient second-order predictor-corrector infeasible primal-dual IPM algorithm with large iteration path updates for solving well-known SDO problems. Journal of Computational and Applied Mathematics, 459(??):??, May 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724006277.

Ahmad:2025:RPE

[AK25b] Sk. Safique Ahmad and Pinki Khatun. A robust parameterized enhanced shift-splitting preconditioner for three-by-three block saddle point problems. Journal of Computational and Applied Mathematics, 459(??):??, May 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704272400606X.

Arnold:2025:OTW

[AKKM25] Anton Arnold, Christian Klein, Jannis Körner, and Jens Markus Melenk. Optimally truncated WKB approximation for the 1D stationary Schrödinger equation in the highly oscillatory regime. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724004898.

Arnal:2025:BSS

[AM25] A. Arnal and J. Monterde. Bézier-Smart surfaces of arbitrary degree. Journal of Computational and Applied Mathematics, 457 (??):??, March 15, 2025.

CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042724005028.

Almutairi:2025:NEE

[AMRS25] Hanan Almutairi, Gérard Meurant, Lothar Reichel, and Miodrag M. Spalević. New error estimates for the conjugate gradient method. Journal of Computational and Applied Mathematics, 459(??):??, May 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042724006058.

Anonymous:2025:Ab

[Ano25a] Anonymous. 1 August 2025. Journal of Computational and Applied Mathematics, 463(??):??, August 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Anonymous:2025:Ac

[Ano25b] Anonymous. 1 August 2025. Journal of Computational and Applied Mathematics, 463(??):??, August 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Anonymous:2025:Da

[Ano25c] Anonymous. 1 December 2025. Journal of Computational and Applied Mathematics, 469(??):??, December 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Anonymous:2025:Ja

[Ano25d]  Anonymous.  1 January

2025. Journal of Computational and Applied Mathematics, 453(??):??, January 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Anonymous:2025:Ma

[Ano25e] Anonymous. 1 March 2025.

Journal of Computational and Applied Mathematics, 456(??):??, March 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Anonymous:2025:Md

[Ano25f] Anonymous. 1 May 2025.

Journal of Computational and Applied Mathematics, 460(??):??, May 1, 2025. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Anonymous:2025:Oa

[Ano25g]  Anonymous.  1 October

2025. Journal of Computational and Applied Mathematics, 466(??):??, October 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Anonymous:2025:Ad

[Ano25h]  Anonymous.  15 August

2025. Journal of Computational and Applied Mathematics, 464(??):??, August 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Anonymous:2025:Db

[Ano25i] Anonymous. 15 December 2025. Journal of Computational and Applied Mathematics, 470(??):??, December 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Anonymous:2025:Jb

[Ano25j] Anonymous. 15 January 2025. Journal of Computational and Applied Mathematics, 454(??):??, January 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Anonymous:2025:Mb

[Ano25k] Anonymous. 15 March 2025. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Anonymous:2025:Mc

[Ano25l] Anonymous. 15 May 2025. Journal of Computational and Applied Mathematics, 459(??):??, May 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Anonymous:2025:Ob

[Ano25m] Anonymous. 15 October 2025. Journal of Computational and Applied Mathematics, 467(??):??, October 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Anonymous:2025:Aa

[Ano25n] Anonymous. April 2025.

Journal of Computational and Applied Mathematics, 458(??):??, April 2025. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Anonymous:2025:EBa

[Ano25o] Anonymous. Editorial Board.

Journal of Computational and Applied Mathematics, 453(??):??, January 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704272400462X.

Anonymous:2025:EBb

[Ano25p] Anonymous. Editorial Board.

Journal of Computational and Applied Mathematics, 454(??):??, January 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724004801.

Anonymous:2025:EBc

[Ano25q] Anonymous. Editorial Board. Journal of Computational and Applied Mathematics, 455(??):??, February 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704272400493X.

Anonymous:2025:EBd

[Ano25r] Anonymous. Editorial Board. Journal of Computational and Applied Mathematics, 456(??):??, March 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724005156.

Anonymous:2025:EBe

[Ano25s] Anonymous. Editorial Board. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724005995.

Anonymous:2025:EBf

[Ano25t] Anonymous. Editorial Board. Journal of Computational and Applied Mathematics, 458(??):??, April 2025. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724006381.

Anonymous:2025:EBg

[Ano25u] Anonymous. Editorial Board.

Journal of Computational and Applied Mathematics, 459(??):??, May 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724006824.

Anonymous:2025:EBh

[Ano25v] Anonymous. Editorial Board.

Journal of Computational and Applied Mathematics, 460(??):??, May 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725000081.

Anonymous:2025:EBi

[Ano25w] Anonymous. Editorial Board.

Journal of Computational and Applied Mathematics, 461(??):??, June 2025. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725000512.

Anonymous:2025:EBj

[Ano25x] Anonymous. Editorial Board.

Journal of Computational and Applied Mathematics, 462(??):??, July 2025. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725000676.

Anonymous:2025:EBk

[Ano25y] Anonymous. Editorial Board. Journal of Computational and Applied Mathematics, 463(??):??, August 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042725000755.

Anonymous:2025:EBl

[Ano25z] Anonymous. Editorial Board. Journal of Computational and Applied Mathematics, 463(??):??, August 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042725000755.

Anonymous:2025:EBm

[Ano25-27] Anonymous. Editorial Board. Journal of Computational and Applied Mathematics, 464(??):??, August 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042725001268.

Anonymous:2025:EBn

[Ano25-28] Anonymous. Editorial Board. Journal of Computational and Applied Mathematics, 465(??):??, September 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042725001578.

Anonymous:2025:EBo

[Ano25-29] Anonymous. Editorial Board.

Journal of Computational and Applied Mathematics, 466(??):??, October 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725001943.

Anonymous:2025:EBp

[Ano25-30] Anonymous. Editorial Board.

Journal of Computational and Applied Mathematics, 467(??):??, October 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725002274.

Anonymous:2025:EBq

[Ano25-31] Anonymous. Editorial Board.

Journal of Computational and Applied Mathematics, 468(??):??, November 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725002547.

Anonymous:2025:EBr

[Ano25-32] Anonymous. Editorial Board.

Journal of Computational and Applied Mathematics, 469(??):??, December 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725002778.

Anonymous:2025:EBs

[Ano25-33] Anonymous. Editorial Board. Journal of Computational and Applied Mathematics, 470(??):??, December 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704272500336X.

Anonymous:2025:F

[Ano25-34] Anonymous. February 2025. Journal of Computational and Applied Mathematics, 455(??):??, February 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Anonymous:2025:Jd

[Ano25-35] Anonymous. July 2025. Journal of Computational and Applied Mathematics, 462 (??):??, July 2025. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Anonymous:2025:Jc

[Ano25-36]  Anonymous.  June 2025.

Journal of Computational and Applied Mathematics, 461(??):??, June 2025. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Anonymous:2025:N

[Ano25-37] Anonymous. November 2025. Journal of Computational and Applied Mathematics, 468(??):??, November 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Anonymous:2025:S

[Ano25-38]  Anonymous.  September

2025. Journal of Computational and Applied Mathematics, 465(??):??, September 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Anonymous:2026:J

[Ano26a]  Anonymous.  1 January

2026. Journal of Computational and Applied Mathematics, 471(??):??, January 1, 2026. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Anonymous:2026:EBa

[Ano26b] Anonymous. Editorial Board.

Journal of Computational and Applied Mathematics, 471(??):??, January 1, 2026. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725003425.

Ai:2025:SMF

[ANZ25] Xiao Ai, Guoxi Ni, and Tieyong Zeng. A sharpening median filter for Cauchy noise with wavelet based regularization. Journal of Computational and Applied Mathematics, 467(??):??, October 15, 2025. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725001402.

Awwal:2025:NSS

[AP25] Aliyu Muhammed Awwal and Nuttapol Pakkaranang. New structured spectral gradient methods for nonlinear least squares with application in robotic motion control problems. Journal of Computational and Applied Mathematics, 469(??):??, December 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725001852.

Aghamohammadi:2025:SRB

[AS25] Faezeh Aghamohammadi and Fatemeh Shakeri. Self representation based methods for tensor completion problem. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724005454.

Al-Shdaifat:2025:NCG

[ASRL25] Hamza Al-Shdaifat and Rosana Rodríguez-López. Non-convolutional general fractional operators and some of their properties. Journal of Computational and Applied Mathematics, 464 (??):??, August 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725000421.

Ali:2026:SBH

[AXF26] Muhammad Aamir Ali, Jianqiang Xie, and Shigeru Furuichi. Sharp bounds of HH–Mercer inequalities for quadrature formulas in fractional calculus as part of numerical analysis. Journal of Computational and Applied Mathematics, 471 (??):??, January 1, 2026. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725002626.

Ben-Artzi:2025:CSM

[BACF25] M. Ben-Artzi, J.-P. Croisille, and D. Fishelov. A compact scheme for the munk boundary-layer equation in one dimension. Journal of Computational and Applied Mathematics, 466(??): ??, October 1, 2025. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725001104.

Baz:2025:ECM

[BAPPF25a] Juan Baz, Pedro Alonso, Juan Manuel Peña, and Raúl Pérez-Fernández. Estimation of the covariance matrix of a Gaussian Markov Random Field under a total positivity constraint. Journal of Computational and Applied Mathematics, 464(??):??, August 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725000585.

Baz:2025:GMR

[BAPPF25b] Juan Baz, Pedro Alonso, Juan Manuel Peña, and Raúl Pérez-Fernández. Gaussian Markov Random Fields over graphs of paths and high relative accuracy. Journal of Computational and Applied Mathematics, 453 (??):??, January 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724003911.

Basak:2025:EAS

[BC25] Arpita Basak and Amit Choudhury. Estimating arrival and service rates from queue length observations. Journal of Computational and Applied Mathematics, 455(??):??, February 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704272400445X.

Boukraa:2025:FRM

[BCD25] Mohamed Aziz Boukraa, Laëtitia Caillé, and Franck Delvare. Fading regularization method for an inverse boundary value problem associated with the biharmonic equation. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724005338.

Bernal:2025:IMH

[BCdR25] Francisco Bernal, Xingyuan Chen, and Gonçalo dos Reis. An iterative method for Helmholtz boundary value problems arising in wave propagation. Journal of Computational and Applied Mathematics, 466 (??):??, October 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725000962.

Branquinho:2025:FMC

[BDFMM25] Amílcar Branquinho, Juan E. F. Díaz, Ana Foulquié-Moreno, and Manuel Mañas. Finite Markov chains and multiple orthogonal polynomials. Journal of Computational and Applied Mathematics, 463(??):??, August 1, 2025. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042724007337.

Bauwelinckx:2025:CPC

[BDvdHV25] Yves-Cédric Bauwelinckx, Jan Dhaene, Milan van den Heuvel, and Tim Verdonck. On the causality-preservation capabilities of generative modelling. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042724005600.

Bruno:2026:AES

[BE26] Ludovico Bruni Bruno and Giacomo Elefante. An algorithm for the estimation of the segmental Lebesgue constant. Journal of Computational and Applied Mathematics, 471(??):??, January 1, 2026. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042725002596.

Bazarra:2025:NAT

[BFQ25] N. Bazarra, J. R. Fernández, and R. Quintanilla. Numerical analysis of a thermoelastic problem of Moore– Gibson–Thompson type with history dependence in the thermal displacement. Journal of Computational and Applied Mathematics, 457 (??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704272400565X.

Bublik:2025:CNN

[BHV25] Ondrej Bublík, Václav Heidler, and Jan Vimmr. Convolution neural network for fluid flow simulations in cascade with oscillating blades. Journal of Computational and Applied Mathematics, 462(??):??, July 2025. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704272400726X.

Bagheri:2025:MBH

[BKST25] Mahsa Bagheri, Faranges Kyanfar, Abbas Salemi, and Azita Tajaddini. A modified block Hessenberg method for low-rank tensor Sylvester equation. Journal of Computational and Applied Mathematics, 456(??):??, March 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724004588.

Bian:2025:DGM

[BLC25] Jialiang Bian, Rui Li, and Zhangxin Chen. A discontinuous Galerkin method for a coupled Brinkman–Biot problem. Journal of Computational and Applied Mathematics, 469(??):??, December 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042725001736.

Barrenechea:2025:NBP

[BPT25a] Gabriel R. Barrenechea, Tristan Pryer, and Alex Trenam. A nodally bound-preserving discontinuous Galerkin method for the drift-diffusion equation. Journal of Computational and Applied Mathematics, 470(??):??, December 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042725001840.

Boroujeni:2025:NSK

[BPT25b] Ahmad Aliyari Boroujeni, Reza Pourgholi, and Seyed Hashem Tabasi. Numerical solutions of KDV and mKDV equations: Using sequence and multi-core parallelization implementation. Journal of Computational and Applied Mathematics, 454 (??):??, January 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042724004333.

Brown:2025:RFR

[BPT^{+}25c] Meredith G. L. Brown, Matt G. Peterson, Irina K. Tezaur, Kara J. Peterson, and Diana L. Bull. Random forest regression feature importance for climate impact pathway detection. Journal of Computational and Applied Mathematics, 464 (??):??, August 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724007271.

B:2026:OTE

[BPT26] Vinoth Kumar B., Sreenivasulu P., and Poornima T. Optimizing the thermal efficiency and drug delivery to narrowing arteries with silver–gold nanoparticles: a numerical study. Journal of Computational and Applied Mathematics, 471 (??):??, January 1, 2026. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725002377.

Butt:2025:NCJ

[BRPP25] S. I. Butt, T. Rasheed, D. Pecarić, and J. Pecarić. New cyclic Jensen type bounds via weighted HH inequalities and Montgomery identity with application to information inequalities. Journal of Computational and Applied Mathematics, 453(??):??, January 1, 2025.

CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042724003868.

Badahmane:2025:GAH

[BRS25] A. Badahmane, A. Ratnani, and H. Sadok. Global accelerated Hermitian and skew-Hermitian splitting preconditioner for the solution of discrete Stokes problems. Journal of Computational and Applied Mathematics, 467 (??):??, October 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042725001359.

Bandha:2025:NST

[BS25a] Raghunath Bandha and Rajen Kumar Sinha. A new space transformed finite element method for elliptic interface problems in R^{n}. Journal of Computational and Applied Mathematics, 457 (??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042724005260.

Bartel:2025:GOT

[BS25b] Andreas Bartel and Manuel Schaller. Goal-oriented time adaptivity for port-Hamiltonian systems. Journal of Computational and Applied Mathematics, 461 (??):??, June 2025. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724006988.

Bertolazzi:2025:PCR

[BS25c] Enrico Bertolazzi and Davide Stocco. Parallel cyclic reduction of padded bordered almost block diagonal matrices. Journal of Computational and Applied Mathematics, 458(??):??, April 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704272400579X.

Buong:2025:FOD

[Buo25] Nguyen Buong. A first order dynamical system and its discretization for a class of variational inequalities. Journal of Computational and Applied Mathematics, 458 (??):??, April 2025. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724005892.

Bilgil:2025:CCC

[BY25] Halis Bilgil and Simge Yüksel. Comparison of Conformable and Caputo fractional grey models. Journal of Computational and Applied Mathematics, 463 (??):??, August 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042725000159.

Chaskalovic:2025:EIA

[CA25a] Joël Chaskalovic and Franck Assous. Enhancing interpolation and approximation error estimates using a novel Taylor-like formula. Journal of Computational and Applied Mathematics, 453 (??):??, January 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042724004023.

Chaskalovic:2025:RFO

[CA25b] Joël Chaskalovic and Franck Assous. A refined first-order expansion formula in R^{n}: Application to interpolation and finite element error estimates. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042724005235.

Chakraborty:2025:SAS

[CAN25] Samiran Chakraborty, Shivam Kumar Agrawal, and Gnaneshwar Nelakanti. Spectral approximated superconvergent method for nonlinear Volterra Hammerstein integral equations with weakly singular kernels. Journal of Computational and Applied Mathematics, 467 (??):??, October 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725001165.

Chacon-Chamorro:2026:ROR

[CCGRR26] Manuela Chacon-Chamorro, Fernando A. Gallego, and J. C. Riano-Rojas. Reducing overfitting in ResNet with Adaptive Lipschitz regularization. Journal of Computational and Applied Mathematics, 471(??):??, January 1, 2026. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725002614.

Cai:2025:ASE

[CCLH25] Ting Cai, Pei Cheng, Xing Liu, and Mingang Hua. Almost sure exponential stability and stochastic stabilization of discrete-time stochastic systems with impulses. Journal of Computational and Applied Mathematics, 453(??):??, January 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724004011.

Chabekh:2025:ASB

[CCT25] Meriem Chabekh, Nadhir Chougui, and Delfim F. M. Torres. Analysis of a shear beam model with suspenders in thermoelasticity of type III. Journal of Computational and Applied Mathematics, 463(??):??, August 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042724007192.

Cuesta:2025:NAR

[CdlHG25] Carlota M. Cuesta, Francisco de la Hoz, and Ivan Girona. Numerical approximation of Riesz–Feller operators on R. Journal of Computational and Applied Mathematics, 454(??):??, January 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042724004436.

Crisci:2025:BRO

[CDPT25] Serena Crisci, Valentina De Simone, Monica Pragliola, and Gerardo Toraldo. Bilevel robust optimization approach for multi-period sparse portfolio selection. Journal of Computational and Applied Mathematics, 470(??): ??, December 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725002432.

Cardone:2026:NSS

[CFCP26] Angelamaria Cardone, Gianluca Frasca-Caccia, and Beatrice Paternoster. Nonstandard schemes for timefractional reaction–advection– diffusion problems. Journal of Computational and Applied Mathematics, 471 (??):??, January 1, 2026. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725002717.

Chen:2025:UEA

[CG25] Yanping Chen and Jixiao Guo. Unconditional error analysis of the linearized transformed L1 virtual element method for nonlinear coupled time-fractional Schrödinger equations. Journal of Computational and Applied Mathematics, 457 (??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724005272.

Cesarone:2025:ODM

[CGR25] Francesco Cesarone, Rosella Giacometti, and Jacopo Maria Ricci. Outlier detection of multivariate data via the maximization of the cumulant generating function. Journal of Computational and Applied Mathematics, 461(??):??, June 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724007052.

Ceng:2025:CTT

[CGuRZ25] Lu-Chuan Ceng, Debdas Ghosh, Habib ur Rehman, and Xiaopeng Zhao. Composite Tseng-type extragradient algorithms with adaptive inertial correction strategy for solving bilevel split pseudomonotone VIP under split common fixed-point constraint. Journal of Computational and Applied Mathematics, 470(??):??, December 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725001979.

Chen:2025:TGF

[CH25] Yanping Chen and Hanzhang Hu. Two-grid finite element methods for space-fractional nonlinear Schrödinger equations. Journal of Computational and Applied Mathematics, 459(??):??, May 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724006186.

Chen:2025:CMF

[Che25a]  Shuangshuang Chen.  A

combined mixed finite element method and discontinuous Galerkin method for hybrid-dimensional fracture models of two-phase flow. Journal of Computational and Applied Mathematics, 459(??):??, May 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724006216.

Chen:2025:SHS

[Che25b]  Yuting Chen.  A spec-

tral Hestenes–Stiefel CG algorithm for large-scale unconstrained optimization in image restoration problems. Journal of Computational and Applied Mathematics, 470(??):??, December 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725002237.

Clavero:2025:ENM

[CJ25] Carmelo Clavero and Juan Carlos Jorge. An efficient numerical method for 1D singularly perturbed parabolic convection-diffusion systems with repulsive interior turning points. Journal of Computational and Applied Mathematics, 470(??):??, December 15, 2025. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042725002420.

Chakib:2025:NSR

[CK25] Abdelkrim Chakib and Ibrahim Khalil. Numerical shape reconstruction approaches for solving free boundary inverse problem for the p-Laplacian operator. Journal of Computational and Applied Mathematics, 463(??):??, August 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042725000032.

Cai:2025:CIR

[CKC25] Xiao-Min Cai, Yi-Fen Ke, and Lin-Jie Chen. Color image restoration based on reduced biquaternion matrix singular value decomposition. Journal of Computational and Applied Mathematics, 460(??):??, May 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042724006769.

Chen:2025:USM

[CKL^{+}25] Meng Chen, Linghua Kong, Yiru Liang, Wei Wang, and Wei Yang. An unconditional stable modified finite element methods for Maxwell’s equation in Kerr-type nonlinear media. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724004965.

Czajka:2025:ERW

[CKŚ25] Marcin Mateusz Czajka, Daria Kubacka, and Aleksandra Świetlicka. Embedding representation of words in sign language. Journal of Computational and Applied Mathematics, 465 (??):??, September 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725001050.

Chao:2025:ERP

[CLJX25] Miantao Chao, Yongzi Lu, Jinbao Jian, and Xiao Xu. An efficient regularized PR splitting type algorithm for twoblock nonconvex linear constrained programs in l_{1/2} regularized compressed sensing problems. Journal of Computational and Applied Mathematics, 453(??):??, January 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724003947.

Chen:2025:NMP

[CLQY25] Jiawei Chen, Jie Liu, Xiaolong Qin, and Jen-Chih Yao. A non-monotone proximal gradient algorithm for solving nonsmooth multiobjective optimization problems with an extending application to robust multiobjective optimization. Journal of Computational and Applied Mathematics, 460(??):??, May 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724006708.

Chen:2025:OCA

[CLWY25] Yanping Chen, Wanxiang Liu, Yang Wang, and Huaming Yi. Optimal convergence analysis of the lowestorder nonconforming virtual element method for general nonlinear parabolic equations. Journal of Computational and Applied Mathematics, 465(??):??, September 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725000913.

Chen:2025:EUC

[CLY25a] Jie Chen, Haoxuan Li, and Xiangfeng Yang. An efficient uncertain chance constrained geometric programming model based on valueat-risk for truss structure optimization problems. Journal of Computational and Applied Mathematics, 458 (??):??, April 2025. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724005958.

Chen:2025:ISN

[CLY25b] Pin-Bo Chen, Gui-Hua Lin, and Zhen-Ping Yang. An inexact semismooth Newton SAA-based algorithm for stochastic nonsmooth SOC complementarity problems with application to a stochastic power flow programming problem. Journal of Computational and Applied Mathematics, 458(??):??, April 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724006095.

Cao:2025:RAF

[CLYY25] Huihui Cao, Hengguang Li, Nianyu Yi, and Peimeng Yin. Regularity and an adaptive finite element method for elliptic equations with Dirac sources on line cracks. Journal of Computational and Applied Mathematics, 462 (??):??, July 2025. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724007143.

Chen:2025:CFP

[CLZ25] Ruyun Chen, Yu Li, and Yongxiong Zhou. On computation of finite-part integrals of highly oscillatory functions. Journal of Computational and Applied Mathematics, 460(??):??, May 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704272400582X.

Cano:2025:EER

[CM25] B. Cano and M. J. Moreta. Efficient exponential Rosenbrock methods till order four. Journal of Computational and Applied Mathematics, 453(??):??, January 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042724004072.

Chen:2025:RHC

[CNQY25] An Chen, Thai Nguyen, Linyi Qian, and Zhixin Yang. The role of health in consumption and portfolio decisionmaking: Insights from statedependent models. Journal of Computational and Applied Mathematics, 457 (??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042724005387.

Caballero:2025:CMF

[COPPS25] Josefa Caballero, Hanna Okrasińska-Płociniczak, Lukasz Płociniczak, and Kishin Sadarangani. Collocation method for a functional equation arising in behavioral sciences. Journal of Computational and Applied Mathematics, 458(??):??, April 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724005910.

Chadjiconstantinidis:2025:BDM

[CP25a] Stathis Chadjiconstantinidis and Konstadinos Politis. Bounds for the distribution and moments of the forward and backward recurrence times in a renewal process. Journal of Computational and Applied Mathematics, 454(??):??, January 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724004151.

Chen:2025:OEE

[CP25b] Yanping Chen and Shanshan Peng. Optimal error estimates of secondorder weighted virtual element method for nonlinear coupled prey-predator equation. Journal of Computational and Applied Mathematics, 467(??):??, October 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042725001323.

Carderera:2025:EFO

[CPSW25] Alejandro Carderera, Sebastian Pokutta, Christof Schütte, and Martin Weiser. An efficient first-order conditional gradient algorithm in data-driven sparse identification of nonlinear dynamics to solve sparse recovery problems under noise. Journal of Computational and Applied Mathematics, 470 (??):??, December 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704272500189X.

Cui:2025:GEM

[CQ25] Chunfeng Cui and Liqun Qi. A genuine extension of the Moore–Penrose inverse to dual matrices. Journal of Computational and Applied Mathematics, 454 (??):??, January 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042724004345.

Choi:2025:HPH

[CS25] Gary P. T. Choi and Mahmoud Shaqfa. Hemispheroidal parameterization and harmonic decomposition of simply connected open surfaces. Journal of Computational and Applied Mathematics, 461(??):??, June 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724007039.

Cui:2025:FDM

[CTZ25] Yanzhen Cui, Shibing Tang, and Chao Zhang. Finite difference methods for stochastic Helmholtz equation driven by white noise. Journal of Computational and Applied Mathematics, 457 (??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704272400534X.

Cao:2025:DIH

[CWX25] Zhengbang Cao, Yimin Wei, and Pengpeng Xie. An L-DEIM induced high order tensor interpolatory decomposition. Journal of Computational and Applied Mathematics, 453(??):??, January 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724003923.

Cholamjiak:2025:DIS

[CXLP25] Prasit Cholamjiak, Zhongbing Xie, Min Li, and Papinwich Paimsang. Double inertial subgradient extragradient algorithm for solving equilibrium problems and common fixed point problems with application to image restoration. Journal of Computational and Applied Mathematics, 460(??):??, May 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724006447.

Chao:2025:IRP

[CXTL25] Wen-Jing Chao, Fu-Quan Xia, Kai Tu, and Jian Lu. An improved random projectiontype algorithm for solving multiple-sets split feasibility problems with application to medical radiotherapy. Journal of Computational and Applied Mathematics, 463(??):??, August 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725000020.

Chen:2025:DPI

[CY25] Xuexue Chen and Cuilian You. Dynamic pricing and inventory model for perishable products with reference price and reference quality. Journal of Computational and Applied Mathematics, 458(??):??, April 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724005867.

Bortoloti:2025:CPG

[dAB25] Marcio Antônio de Andrade Bortoloti. Computational performance of a generalized descent gradient method based algorithm with conformable fractional-order derivatives. Journal of Computational and Applied Mathematics, 462(??):??, July 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724007283.

Davis:2025:STM

[DCT^{+}25] Warren L. Davis IV, Max L.

Carlson, Irina K. Tezaur, Diana L. Bull, Kara J. Peterson, and Laura P. Swiler. Spatiotemporal multivariate cluster evolution analysis for detecting and tracking climate impacts. Journal of Computational and Applied Mathematics, 465(??):??, September 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725000986.

Du:2025:NES

[DdF25] Zhibin Du and Carlos M.

da Fonseca. A note on the eigenvalues of a Sylvester– Kac type matrix with offdiagonal biperiodic perturbations. Journal of Computational and Applied Mathematics, 461(??):??, June 2025. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042724006770.

Diana:2025:NUS

[DG25] Derick Diana and Tim Gebbie. Non-uniformly sampled simulated price impact of an order-book. Journal of Computational and Applied Mathematics, 456(??):??, March 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042724004515.

DellAccio:2025:RAF

[DGMN25] Francesco Dell’Accio, Allal Guessab, Gradimir V. Milovanović, and Federico Nudo. Reconstructing algebraic functions from a nonconforming exponential weighted enriched finite element. Journal of Computational and Applied Mathematics, 466(??):??, October 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042725001189.

Duan:2025:EES

[DGYH25] Dongmei Duan, Fuzheng Gao, Jinjin Yang, and Xiaoming He. Error estimates of semi-implicit numerical scheme for a diffuse interface model of twophase magnetohydrodynamic flows. Journal of Computational and Applied Mathematics, 465(??):??, September 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725000950.

Du:2025:DAD

[DH25] Xiongxiong Du and Xiaoling Han. Dynamics analysis of a discrete diffusion predatorprey model with prey refuge. Journal of Computational and Applied Mathematics, 470(??):??, December 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725002201.

Dembinska:2025:MLI

[DJ25] Anna Dembińska and Krzysztof Jasiński. Maximum likelihood inference about parameters of geometric lifetimes of heterogeneous components from data collected till failure of a k-out-of-n: G system. Journal of Computational and Applied Mathematics, 454(??):??, January 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724004448.

Dykes:2025:SLI

[DKM^{+}25] Laura Dykes, Mykhailo Kuian, Thomas Mach, Silvia Noschese, and Lothar Reichel. Solution of linear ill-posed operator equations by modified truncated singular value expansion. Journal of Computational and Applied Mathematics, 467 (??):??, October 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725001360.

Ding:2025:CAF

[DL25] Qianqian Ding and Mingxia Li. Convergence analysis of finite element method for incompressible magnetohydrodynamics system with variable density. Journal of Computational and Applied Mathematics, 462(??):??, July 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724007180.

Diegel:2025:ACD

[DLR25] Amanda E. Diegel, Xuejian Li, and Leo G. Rebholz. Analysis of continuous data assimilation with large (or even infinite) nudging parameters. Journal of Computational and Applied Mathematics, 456(??):??, March 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724004709.

DInverno:2025:MIR

[DMS^{+}25] Giuseppe Alessio D’Inverno, Saeid Moradizadeh, Sajad Salavatidezfouli, Pasquale Claudio Africa, and Gianluigi Rozza. Mesh-informed Reduced Order Models for aneurysm rupture risk prediction. Journal of Computational and Applied Mathematics, 470(??):??, December 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725002419.

Dong:2025:IPC

[Don25] Xiao Liang Dong. An improved PRP conjugate gradient algorithm with a damping factor within an automatic restarting strategy. Journal of Computational and Applied Mathematics, 468(??):??, November 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725001529.

Dick:2025:QMC

[DP25] Josef Dick and Friedrich Pillichshammer. Quasi-Monte Carlo integration over R^{s} based on digital nets. Journal of Computational and Applied Mathematics, 462 (??):??, July 2025. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704272400699X.

Draganescu:2025:SEC

[DS25] Andrei Draganescu and Florin Spinu. Sharp estimates for the convergence rate of Orthomin(k) for a class of linear systems. Journal of Computational and Applied Mathematics, 470 (??):??, December 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042725002134.

Menezes:2025:PIH

[dSMSdO^{+}25] Matheus da Silva Menezes, Paulo Henrique Lopes Silva, João Paulo Caraú de Oliveira, Raimundo Leandro Andrade Marques, and Ivan Mezzomo. A parallel iterative hybrid Gauss–Jacobi–Seidel method. Journal of Computational and Applied Mathematics, 468(??):??, November 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042725001426.

Dong:2025:PPS

[DWZ25] Lixiu Dong, Cheng Wang, and Zhengru Zhang. A positivity-preserving, secondorder energy stable and convergent numerical scheme for a ternary system of macromolecular microsphere composite hydrogels. Journal of Computational and Applied Mathematics, 462 (??):??, July 2025. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724007118.

Du:2025:EEH

[DZZ25] Caiyue Du, Di Zhao, and Qiang Zhang. Error estimate of high order Runge– Kutta local discontinuous Galerkin method for nonlinear convection-dominated Sobolev equation. Journal of Computational and Applied Mathematics, 469 (??):??, December 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725001712.

Elghandouri:2025:EWP

[EES25]  Mohammed  Elghandouri,

Khalil Ezzinbi, and Lamiae Saidi. Exploring wellposedness and asymptotic behavior in an Advection– Diffusion–Reaction (ADR) model. Journal of Computational and Applied Mathematics, 462(??):??, July 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724007131.

Escribano:2025:BZS

[EG25] C. Escribano and R. Gonzalo. Boundedness of zeros of Sobolev orthogonal polynomials via generalised eigenvalues. Journal of Computational and Applied Mathematics, 465(??):??, September 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042725000718.

Egidi:2025:FND

[EGM25] Nadaniela Egidi, Josephin Giacomini, and Pierluigi Maponi. Fast numerical derivatives of univariate functions on non-uniform grids. Journal of Computational and Applied Mathematics, 467(??):??, October 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042725001347.

ElHaoui:2026:QHO

[El 26] Youssef El Haoui. Quaternionic hyperbolic offset linear canonical transform. Journal of Computational and Applied Mathematics, 471 (??):??, January 1, 2026. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042725002699.

Eshkuvatov:2026:MHP

[EMID26] Z. K. Eshkuvatov, H. X.

Mamatova, Sh. Ismail, and F. Deraman. Modified homotopy perturbation method for solving singular integral equations of the first kind. Journal of Computational and Applied Mathematics, 471(??):??, January 1, 2026. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725002675.

El-Yahyaoui:2026:OCC

[EYJKL26] Youness El-Yahyaoui, Mohsine Jennane, El Mostafa Kalmoun, and Lahoussine Lafhim. Optimality conditions for a class of nonsmooth semidefinite bilevel optimization problems via convexifications applied to bilevel supply chain under uncertainty. Journal of Computational and Applied Mathematics, 471(??):??, January 1, 2026. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725002481.

Frantsuzov:2025:GAB

[FA25] Viktor A. Frantsuzov and Anton V. Artemyev. A global argument-based algorithm for finding complex zeros and poles to investigate plasma kinetic instabilities. Journal of Computational and Applied Mathematics, 456(??):??, March 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042724004667.

Fedakar:2025:NSS

[FADA25] Burcu Fedakar, Ilhame Amirali, Muhammet Enes Durmaz, and Gabil M. Amiraliyev. Numerical solutions for second-order neutral Volterra integro-differential equations: Stability analysis and finite difference method. Journal of Computational and Applied Mathematics, 459(??):??, May 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042724006198.

Fahd:2025:FSP

[Fah25] Hcini Fahd. A fast solver for pentadiagonal Toeplitz systems with multiple righthand sides. Journal of Computational and Applied Mathematics, 470(??):??, December 15, 2025. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042725002110.

Fan:2025:TVR

[Fan25] Bin Fan. Total variation regularization for recovering the spatial source term in a time-fractional diffusion equation. Journal of Computational and Applied Mathematics, 463(??):??, August 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725000329.

Flavien:2025:JSP

[FBM25] Alonzo Flavien, Dia Ben Mansour, and Saad Mazen. Joint state-parameter estimation and inverse problems governed by reactionadvection-diffusion type PDEs with application to biological Keller–Segel equations and pattern formation. Journal of Computational and Applied Mathematics, 461(??):??, June 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724007027.

Fahd:2025:EAP

[FBZ25] Hcini Fahd, Skander Belhaj, and Yulin Zhang. Efficient algorithms for perturbed symmetrical Toeplitzplus-Hankel systems. Journal of Computational and Applied Mathematics, 457 (??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042724005818.

Frasca-Caccia:2025:FDS

[FC25] Gianluca Frasca-Caccia. Finite difference schemes with non polynomial local conservation laws. Journal of Computational and Applied Mathematics, 458(??):??, April 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042724005788.

Fatone:2025:LCD

[FF25] Lorella Fatone and Daniele Funaro. Low-cost denoising and deblurring using a novel nonlinear diffusion technique. Journal of Computational and Applied Mathematics, 461(??):??, June 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704272400671X.

Farhloul:2025:FEA

[FFDL25] M. Farhloul, N. Fall, I. Dione, and S. Léger. Finite element analysis of a pseudostresspressure-velocity formulation of the stationary Navier– Stokes equations. Journal of Computational and Applied Mathematics, 469 (??):??, December 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725001797.

Faheem:2025:DDD

[FG25] Mo Faheem and Bapan Ghosh. Dynamics of a delayed discrete-time predator prey model proposed from a nonstandard finite difference scheme. Journal of Computational and Applied Mathematics, 458(??):??, April 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724005946.

Ferreira:2025:NFD

[FGGdO25] José A. Ferreira, Mario Grassi, Elías Gudiño, and Paula de Oliveira. Non-Fickian diffusion enhanced by temperature. Journal of Computational and Applied Mathematics, 457 (??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724005624.

Farouki:2025:AMC

[FKVZ25] Rida T. Farouki, Marjeta Knez, Vito Vitrih, and Emil Zagar. Application of a metric for complex polynomials to bounded modification of planar Pythagoreanhodograph curves. Journal of Computational and Applied Mathematics, 456 (??):??, March 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042724004849.

Fu:2025:ABL

[FLC^{+}25] Zhenwu Fu, Yang Li, Yong Chen, Bo Han, and Hao Tian. An accelerated Bouligand– Landweber method based on projection and Nesterov acceleration for nonsmooth ill-posed problems. Journal of Computational and Applied Mathematics, 462 (??):??, July 2025. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042724004576.

Fan:2025:IAT

[FLLZ25] Enyu Fan, Jichun Li, Yang Liu, and Yangpeng Zhang. Isogeometric analysis for time-dependent Maxwell’s equations in complex media. Journal of Computational and Applied Mathematics, 465(??):??, September 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042725000810.

Ferreira:2025:CAE

[FLS25] Chelo Ferreira, José L. López, and Ester Pérez Sinusía. Convergent and asymptotic expansions of the displacement elastodynamic integral in terms of known functions. Journal of Computational and Applied Mathematics, 460(??):??, May 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724006435.

Fang:2025:SUC

[FSC25] Jun Fang, Zhijun Shen, and Xia Cui. Stability and uniqueness of coupled nonlinear finite element solution for anisotropic diffusion equation with nonlinear capacity term. Journal of Computational and Applied Mathematics, 469(??):??, December 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725001785.

Fu:2025:NMB

[FWL25] Zu-Feng Fu, Shi-Liang Wu, and Liang Li. New modulusbased matrix splitting method for the vertical nonlinear complementarity problem. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724005004.

Ghanem:2025:SDD

[GBB25] Aicha Ghanem, Touffik Bouremani, and Djamel Benterki. On the solution of Dolichobrachistochrone differential game via dynamic programming approach. Journal of Computational and Applied Mathematics, 461(??):??, June 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042724007088.

Gomez-Bueno:2026:WBP

[GBFNR26] I. Gómez-Bueno, E. D. Fernández-Nieto, and S. Rubino. Well-balanced PODbased reduced-order models for finite volume approximation of hyperbolic balance laws. Journal of Computational and Applied Mathematics, 471(??):??, January 1, 2026. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042725002493.

Gao:2025:IOI

[GC25] Qin Gao and Minhong Chen. An inner–outer iterative method for inverse Sturm– Liouville problems. Journal of Computational and Applied Mathematics, 463 (??):??, August 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725000299.

Gao:2025:NIR

[GCN25] Wansheng Gao, Ludovic Chamoin, and Insa Neuweiler. Non-intrusive reduced basis two-grid method for flow and transport problems in heterogeneous porous media. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724005697.

Ghosh:2025:NGB

[GCT25] Indranil Ghosh, Huey Tyng Cheong, and Kok Lay Teo. A novel gradient-based discrete time-delayed optimization algorithm for optimal control problems with Caputo– Fabrizio fractional derivative. Journal of Computational and Applied Mathematics, 464(??):??, August 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704272500041X.

Guo:2025:LMC

[GD25] Feng Guo and Weizhong Dai.

A linear and mass conservative scheme for the thermal soliton model based on nonlinear Schrödinger and heat transfer equations. Journal of Computational and Applied Mathematics, 464 (??):??, August 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042725000445.

Gidisu:2025:BDE

[GH25a] Perfect Y. Gidisu and Michiel E. Hochstenbach. Block discrete empirical interpolation methods. Journal of Computational and Applied Mathematics, 454(??):??, January 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042724004357.

Greif:2025:MMM

[GH25b] Chen Greif and Yunhui He. Monolithic multigrid for the marker-and-cell discretization of the Stokes–Darcy equations. Journal of Computational and Applied Mathematics, 463(??):??, August 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042725000330.

Ghalini:2025:ESC

[GHD25] Rohollah Ghaedi Ghalini, Esmail Hesameddini, and Hojatollah Laeli Dastjerdi. An efficient spectral collocation method for solving Volterra delay integral equations of the third kind. Journal of Computational and Applied Mathematics, 454 (??):??, January 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704272400387X.

Gao:2025:NNB

[GHMN25] Jinwu Gao, Haomiao Hu, Farshid Mehrdoust, and Idin Noorani. A neural networkbased method for pricing American options and assessing implied volatility under uncertainty. Journal of Computational and Applied Mathematics, 470(??): ??, December 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704272500233X.

Ghoneim:2025:MPF

[Gho25] Adam Y. Ghoneim. Meshfree phase-field modeling of threephase flow using smoothed particle hydrodynamics with differential reproducing kernels and artificial compressibility. Journal of Computational and Applied Mathematics, 469(??):??, December 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725001682.

Griese:2025:PFB

[GHRK25] Franziska Griese, Fabian Hoppe, Alexander Rüttgers, and Philipp Knechtges. Preconditioned FEM-based neural networks for solving incompressible fluid flows and related inverse problems. Journal of Computational and Applied Mathematics, 469(??):??, December 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042725001773.

Ge:2025:NEI

[GHZ25] Liang Ge, Jian Huang, and Jianwei Zhou. Novel efficient iterative schemes for linear finite element approximations of elliptic optimal control problems with integral constraint on the state. Journal of Computational and Applied Mathematics, 456(??):??, March 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042724004746.

Goroncy:2025:DTT

[GJ25] Agnieszka Goroncy and Krzysztof Jasiński. Discrete time three-state k-out-of-n system’s failure and numbers of components in each state. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724005041.

Guo:2025:EAE

[GJWV25] Zhenwei Guo, Tongsong Jiang, Gang Wang, and V. I. Vasil’ev. An efficient algorithm for the eigenvalue problem of a Hermitian quaternion matrix in quantum chemistry. Journal of Computational and Applied Mathematics, 463(??):??, August 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725000317.

Guo:2025:MBI

[GLL^{+}25] Wenxiu Guo, Jianqing Li, Yongyao Li, Xiaoping Lu, and Hua Zheng. A modulusbased inner-outer iteration method for nonlinear complementarity problems. Journal of Computational and Applied Mathematics, 463 (??):??, August 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725000044.

Gong:2025:GBO

[GLTW25] Zhaohua Gong, Chongyang Liu, Kok Lay Teo, and Yonghong Wu. A gradientbased optimization algorithm to solve optimal control problems with conformable fractional-order derivatives. Journal of Computational and Applied Mathematics, 454(??):??, January 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724004187.

Garcia:2025:NSS

[GNUV25] A. García, M. Negreanu, F. Ureña, and A. M. Vargas. On the numerical solution to space fractional differential equations using meshless finite differences. Journal of Computational and Applied Mathematics, 457 (??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724005703.

Groselj:2025:RCP

[GP25a] Jan Groselj and Ada Sadl Praprotnik. Rational C^{1} cubic Powell–Sabin B-splines with application to representation of ruled surfaces. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724005405.

Guo:2025:IFA

[GP25b] Xiaofeng Guo and Jianyu Pan. An improved fast approximation to two-sided variable-order space-fractional diffusion equation and its preconditioning. Journal of Computational and Applied Mathematics, 465(??): ??, September 2025. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725000706.

Gu:2025:DCQ

[GR25] Yurun Gu and Clément Rey.

Deterministic computation of quantiles in a Lipschitz framework. Journal of Computational and Applied Mathematics, 458(??):??, April 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724005922.

Gudi:2025:EEB

[GS25] Thirupathi Gudi and Pratibha Shakya. Error estimates for a bubble-enriched C^{0} interior penalty method to a reduced state constrained elliptic optimal control problem. Journal of Computational and Applied Mathematics, 460(??):??, May 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042724006459.

Guo:2025:HOB

[GWHY25] Hui Guo, Kaixuan Wang, Jian Huang, and Yang Yang. High-order bound-preserving finite difference methods for incompressible two-phase flow in porous media. Journal of Computational and Applied Mathematics, 469 (??):??, December 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042725001724.

Gu:2025:PMS

[GXC25] Shuting Gu, Ming Xiao, and Rui Chen. Projection method for steady states of Cahn– Hilliard equation with the dynamic boundary condition. Journal of Computational and Applied Mathematics, 469(??):??, December 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042725001888.

Haghighi:2025:MSC

[Hag25]  Amir Haghighi.  Mean-

square convergence of an explicit derivative-free truncated method for nonlinear SDEs covering the noncommutative noise case. Journal of Computational and Applied Mathematics, 460(??):??, May 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724006642.

Hu:2025:FIS

[HB25] Tony Hu and James Bremer.

A frequency-independent solver for systems of linear ordinary differential equations. Journal of Computational and Applied Mathematics, 470 (??):??, December 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725002109.

Hou:2025:ISS

[HBG25] Liangliang Hou, Gongbing Bi, and Qianqian Guo. An improved sparrow search algorithm optimized Light-GBM approach for credit risk prediction of SMEs in supply chain finance. Journal of Computational and Applied Mathematics, 454 (??):??, January 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724004461.

Huang:2025:FRM

[HC25] Lizhen Huang and Lei Chen. Fuzzy random multiobjective optimization using a novel mixed fuzzy random inverse DEA model in inputoutput production. Journal of Computational and Applied Mathematics, 470 (??):??, December 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725002316.

Huang:2025:EPP

[HCCC25] Can Huang, Huangxin Chen, Qing Cheng, and Lijun Chen. Efficient positivity preserving schemes for stochastic complex systems. Journal of Computational and Applied Mathematics, 462 (??):??, July 2025. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704272400712X.

Honain:2025:GET

[HFSK25] Aljowhara H. Honain, Khaled M. Furati, Ibrahim O. Sarumi, and Abdul Q. M. Khaliq. Generalized exponential time differencing for fractional oscillation models. Journal of Computational and Applied Mathematics, 461 (??):??, June 2025. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724007040.

Hernando:2025:LCI

[HGGPD^{+}25] Antonio Hernando, José Luis Galán-García, Yolanda Padilla-Domínguez, María Ángeles Galán-García, and Gabriel Aguilera-Venegas. A library in CoCoA for implementing railway interlocking systems. Journal of Computational and Applied Mathematics, 466(??):??, October 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725001098.

Huang:2025:OSB

[HGZL25] Xiaoli Huang, Yuelin Gao, Bo Zhang, and Xia Liu. An outer space branchreduction-bound algorithm using second-order cone relaxation with regional reduction strategy for solving equivalent generalized linear multiplicative programming. Journal of Computational and Applied Mathematics, 462(??):??, July 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724007295.

He:2025:IAA

[HHF25] Xin He, Rong Hu, and Ya-Ping Fang. Inertial accelerated augmented Lagrangian algorithms with scaling coefficients to solve exactly and inexactly linearly constrained convex optimization problems. Journal of Computational and Applied Mathematics, 460(??):??, May 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042724006733.

Han:2025:PMA

[HJ25] Wei-Wei Han and Yao-Lin Jiang. A penalty method for approximation of the stationary Stokes–Darcy problem. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042724005211.

Hesamian:2026:FMR

[HJC26] Gholamreza Hesamian, Arne Johannssen, and Nataliya Chukhrova. A fuzzy multiple regression model adopted with locally weighted and interval-valued techniques. Journal of Computational and Applied Mathematics, 471(??):??, January 1, 2026. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042725002651.

Ham:2025:FOF

[HKH^{+}25] Seokjun Ham, Seungyoon Kang, Youngjin Hwang, Gyeonggyu Lee, Soobin Kwak, Jyoti, and Junseok Kim. A fourth-order finite difference method for the Allen–Cahn equation. Journal of Computational and Applied Mathematics, 453 (??):??, January 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724004084.

Harbrecht:2025:CQL

[HKM25] Helmut Harbrecht, Rüdiger Kempf, and Michael Multerer. Construction of quasilocalized dual bases in reproducing kernel Hilbert spaces. Journal of Computational and Applied Mathematics, 465(??):??, September 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725000603.

Ha:2025:PTV

[HKY25] Mijin Ha, Donghyun Kim, and Ji-Hun Yoon. Pricing of timer volatility-barrier options under Heston’s stochastic volatility model. Journal of Computational and Applied Mathematics, 457 (??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724005582.

Hu:2025:STN

[HLC25] Jiuhua Hu, Wing Tat Leung, and Eric Chung. Space-time non-local multi-continua multiscale method for channelizedmedia parabolic equations. Journal of Computational and Applied Mathematics, 469(??):??, December 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042725001839.

Han:2025:IFF

[HLF^{+}25] Ruyi Han, Shenghai Liao, Shujun Fu, Yuanfeng Zhou, Yuliang Li, and Hongbin Han. Inter-frame feature fusion enhanced spatiotemporal consistent video inpainting with sample-based techniques and adaptive local search. Journal of Computational and Applied Mathematics, 464(??):??, August 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704272500038X.

Huang:2025:DAF

[HLL25] Yunqing Huang, Chanjie Li, and Jichun Li. Developing and analyzing a FDTD method for simulation of metasurfaces. Journal of Computational and Applied Mathematics, 460(??): ??, May 1, 2025. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724006125.

Hou:2025:ECN

[HLQF25] Yabin Hou, Jingwei Li, Yuanyang Qiao, and Xinlong Feng. An efficient Crank–Nicolson scheme with preservation of the maximum bound principle for the high-dimensional Allen–Cahn equation. Journal of Computational and Applied Mathematics, 465(??):??, September 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725001013.

He:2025:PEI

[HLS25] Jun He, Yanmin Liu, and Xiaowei Shen. Pareto eigenvalue inclusion theorems and copositive tensors. Journal of Computational and Applied Mathematics, 454 (??):??, January 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724004424.

Han:2025:GCS

[HLWZ25] Yang Han, Ping Lin, Lixiu Wang, and Qian Zhang. Graddiv-conforming spectral element method for fourthorder div problems. Journal of Computational and Applied Mathematics, 465 (??):??, September 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042725001141.

Hu:2025:VRV

[HLZ25] Qingjie Hu, Ruyun Li, and Yanyan Zhang. A vector restricted variant MVHS+ CG method based algorithm for unconstrained vector optimization problems. Journal of Computational and Applied Mathematics, 463 (??):??, August 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042725000019.

Huo:2025:LFC

[HMZ25] Fuchang Huo, Weilong Mo, and Yulin Zhang. A lockingfree conforming discontinuous Galerkin finite element method for linear elasticity problems. Journal of Computational and Applied Mathematics, 465(??):??, September 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042725000974.

Heydari:2025:NMB

[HNHR25] M. H. Heydari, J. Navari, M. Hosseininia, and M. Razzaghi. A numerical method based on the hat functions to solve a category of nonlinear fractional integro-differential equations involving Caputo– Hadamard derivative. Journal of Computational and Applied Mathematics, 463 (??):??, August 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725000214.

Howard:2025:CTI

[HRO^{+}25] Kyler Howard, Chris Rocheleau, Trevor Overton, Joel Barraza Nava, Mason Faldet, Kristina Moen, Summer Soller, Tyler Stephens, Esther van de Lagemaat, Natalie Wijesinghe, Kaylee Wong Dolloff, Nilton Barbosa da Rosa, and Jennifer L. Mueller. A comparison of techniques to improve pulmonary EIT image resolution using a database of simulated EIT images. Journal of Computational and Applied Mathematics, 460(??):??, May 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724006630.

Huang:2025:SCS

[HRST25] Yu Huang, Narges Tohidi Rad, Mohammad Hadi Noori Skandari, and Emran Tohidi. A spectral collocation scheme for solving nonlinear delay distributed-order fractional equations. Journal of Computational and Applied Mathematics, 456(??):??, March 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704272400476X.

Hu:2025:NAS

[HU25] N. Hu and S. Utyuzhnikov. Nonlocal active sound control for composite regions with common boundaries. Journal of Computational and Applied Mathematics, 465(??):??, September 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042725000925.

Hung:2025:SCC

[Hun25] Nguyen Van Hung. On semicontinuity and continuity of solution maps of parametric generalized multiobjective games in fuzzy environments. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704272400551X.

Hussain:2025:GTF

[HuR25] Furqan Hussain and Mujeeb ur Rehman. On general tempered fractional calculus with Luchko kernels. Journal of Computational and Applied Mathematics, 458(??):??, April 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724005879.

He:2025:NTM

[HV25] Jiewen He and Seakweng Vong. Newton-type methods for solving vertical linear complementarity problems. Journal of Computational and Applied Mathematics, 460(??):??, May 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724006666.

He:2025:NPP

[HWZ25] Wenming He, Jiming Wu, and Zhimin Zhang. A novel post-processed finite element method and its convergence for partial differential equations. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724005673.

Hu:2025:CIV

[HY25a] Xiang Hu and Jing Yao. A combined integer-valued autoregressive process with actuarial applications. Journal of Computational and Applied Mathematics, 459(??): ??, May 15, 2025. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042724006320.

Hu:2025:SOL

[HY25b] Zuopeng Hu and Yanlong Yang. Stability of open-loop Nash equilibria for n-person nonzero-sum bounded rationality generalized differential games. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042724005090.

Huang:2025:NDD

[HYZ^{+}25] Jian Huang, Zhen Yue, Jing Zhang, Kai Fu, and Xilin Feng. Numerical discretization of a Darcy–Forchheimer flow with variable density and heat transfer. Journal of Computational and Applied Mathematics, 457 (??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042724005016.

Islam:2025:DCF

[IHA25] Md. Asraful Islam, Ivna Ratul Hassan, and Payer Ahmed. Dynamic complexity of fifthdimensional Henon map with Lyapunov exponent, permutation entropy, bifurcation patterns and chaos. Journal of Computational and Applied Mathematics, 466 (??):??, October 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725000627.

Izuchukwu:2025:CIR

[IRS25] Chinedu Izuchukwu, Simeon Reich, and Yekini Shehu. Convergence of an inertial reflected-forward-backward splitting algorithm for solving monotone inclusion problems with application to image recovery. Journal of Computational and Applied Mathematics, 460(??):??, May 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724006538.

Isik:2025:MMR

[ITM25] Osman Rasit Isik, Necibe Tuncer, and Maia Martcheva. A mathematical model for the role of vaccination and treatment in measles transmission in Turkey. Journal of Computational and Applied Mathematics, 457 (??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042724005569.

Jiang:2025:NAE

[JCCZ25] Ping Jiang, Jiejie Chen, Boshan Chen, and Zhigang Zeng. A novel adaptive eventtriggered controller for exponential quasi-synchronization and synchronization problems to coupled inertial memristive neural networks with time delays. Journal of Computational and Applied Mathematics, 460(??):??, May 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042724006745.

Jafari:2025:ALP

[JFP25] A. A. Jafari, S. Firoozi, and A. Pak. Analysis of the lifetime performance index in Gompertz distribution based on k-record values. Journal of Computational and Applied Mathematics, 453(??):??, January 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704272400400X.

Jian:2025:IBA

[JJM25] Jinbao Jian, Jiachen Jin, and Guodong Ma. An IDFPMbased algorithm without Lipschitz continuity to constrained nonlinear equations for sparse signal and blurred image restoration problems. Journal of Computational and Applied Mathematics, 456(??):??, March 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724004837.

Jin:2025:CAS

[JL25] Qinian Jin and Yanjun Liu.

Convergence analysis of a stochastic heavy-ball method for linear ill-posed problems. Journal of Computational and Applied Mathematics, 470(??):??, December 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704272500216X.

Jiang:2025:DGM

[JLC25] Xia Jiang, Rui Li, and Zhangxin Chen. A discontinuous Galerkin method for the Brinkman–Darcy-transport problem. Journal of Computational and Applied Mathematics, 453(??):??, January 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724004047.

Jia:2025:SNP

[JMX25] Caijuan Jia, Yan Meng, and Jiaxin Xiao. Spatially nonhomogeneous patterns for a modified Leslie–Gower model with predator-taxis. Journal of Computational and Applied Mathematics, 464 (??):??, August 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725000573.

Jahangiri:2025:ERB

[JN25a] Mohammadreza Jahangiri and Alireza Nazemi. An efficient RNN based algorithm for solving fuzzy nonlinear constrained programming problems with numerical experiments. Journal of Computational and Applied Mathematics, 463(??):??, August 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724006964.

Jornet:2025:PST

[JN25b] Marc Jornet and Juan J. Nieto. The Peano–Sard theorem for fractional operators with Mittag-Leffler kernel and application in classical numerical approximation. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724005119.

Janes:2025:NPO

[JS25] Andrew Janes and John R.

Singler. A new proper orthogonal decomposition method with second difference quotients for the wave equation. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724005284.

Jayswal:2025:OCD

[JU25] Anurag Jayswal and Gaurav Uniyal. Optimal conditions and duality results for a semiinfinite variational programming problem and its Mond– Weir dual involving Caputo– Fabrizio fractional derivatives. Journal of Computational and Applied Mathematics, 468(??):??, November 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725001438.

Jacob:2025:CQN

[JVW25] Birgit Jacob, Lukas Vorberg, and Christian Wyss. Computing the quadratic numerical range. Journal of Computational and Applied Mathematics, 453(??):??, January 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042724002991.

Jiang:2025:THC

[JWS^{+}25] Meixuan Jiang, Yun Wang, Hu Shao, Ting Wu, and Weiwei Sun. Two hybrid conjugate gradient based algorithms on Riemannian manifolds with adaptive restart strategy for nonconvex optimization problems. Journal of Computational and Applied Mathematics, 461 (??):??, June 2025. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042724007003.

Jia:2025:FSD

[JWZZ25] Jinhong Jia, Hong Wang, Zhaojie Zhou, and Xiangcheng Zheng. A fast sequentially-decoupled matrixdecomposed algorithm for variable-order time-fractional optimal control and error estimate. Journal of Computational and Applied Mathematics, 469(??):??, December 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042725001815.

Karageyik:2025:IFI

[Kar25a] Basak Bulut Karageyi̇k. An integrated framework for indicator-based decision analysis in proportional-XL reinsurance. Journal of Computational and Applied Mathematics, 461(??):??, June 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724006897.

Kartal:2025:DFO

[Kar25b] Senol Kartal. A discrete fractional order Cournot duopoly game model with relative profit delegation: Stability, bifurcation, chaos, 0–1 testing and control. Journal of Computational and Applied Mathematics, 457 (??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724005326.

Karageorghis:2025:MLM

[KC25] Andreas Karageorghis and C. S. Chen. Multi-level method of fundamental solutions for solving polyharmonic problems. Journal of Computational and Applied Mathematics, 456(??): ??, March 1, 2025. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724004692.

Kumar:2025:AFR

[KCM25] D. Kumar, A. K. B. Chand, and P. R. Massopust. Approximation with fractal radial basis functions. Journal of Computational and Applied Mathematics, 454 (??):??, January 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724004497.

Kattan:2025:SIE

[KIF^{+}25] Doha A. Kattan, Umar Ishtiaq, Muhammad Farhan, Ioannis K. Argyros, Mohammad Akram, and Samundra Regmi. Solution of an integral equation in controlled rectangular metric spaces via weakly contractive mappings. Journal of Computational and Applied Mathematics, 454(??):??, January 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724004369.

Kang:2025:PPU

[KKHK25] Seungyoon Kang, Soobin Kwak, Youngjin Hwang, and Junseok Kim. Positivity preserving and unconditionally stable numerical scheme for the three-dimensional modified Fisher–Kolmogorov– Petrovsky–Piskunov equation. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724005223.

Kalantarov:2025:BBD

[KKV25] V. K. Kalantarov, H. V.

Kalantarova, and O. Vantzos. Backward behavior and determining functionals for chevron pattern equations. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724005314.

Kunwar:2025:DDL

[KL25] Hemanta Kunwar and Hyesuk Lee. Domain decomposition with local time discretization for the nonlinear Stokes–Biot system. Journal of Computational and Applied Mathematics, 457 (??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724005594.

Kjelldahl:2025:NSG

[KM25] Vilma Kjelldahl and Ken Mattsson. Numerical simulation of the generalized modified Benjamin–Bona–Mahony equation using SBP-SAT in time. Journal of Computational and Applied Mathematics, 459(??):??, May 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042724006253.

Karban:2025:OSP

[KPK25] Pavel Karban, David Pánek, and Jan Kaska. Opensource platform for simulation of physical fields: Agros. Journal of Computational and Applied Mathematics, 465(??):??, September 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042725001049.

Kaliraj:2025:FTS

[KPTS25] K. Kaliraj, P. K. Lakshmi Priya, V. Tamilarasan, and S. Suresh. Finite time stability of neutral multiterm fractional order time-varying delay systems. Journal of Computational and Applied Mathematics, 461(??):??, June 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042724007076.

Kittaneh:2025:NUB

[KR25] Fuad Kittaneh and M. H. M. Rashid. New upper bounds for the q-numerical radii of 2×2 operator matrices. Journal of Computational and Applied Mathematics, 467 (??):??, October 15, 2025.

CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725001335.

Kovtunenko:2026:CAS

[KR26] Victor A. Kovtunenko and Yves Renard. Convergence analysis of semi-smooth Newton method for mixed FEM approximations of dynamic two-body contact and crack problems. Journal of Computational and Applied Mathematics, 471(??):??, January 1, 2026. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725002365.

Kiouvrekis:2025:DPS

[KS25a] Yiannis Kiouvrekis and Ioannis S. Stamatiou. Domain preserving and strongly converging explicit scheme for the stochastic SIS epidemic model. Journal of Computational and Applied Mathematics, 456(??):??, March 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724004680.

Knobloch:2025:AAM

[KS25b] Petr Knobloch and René Schneider. On adaptive anisotropic mesh optimization for convection–diffusion problems. Journal of Computational and Applied Mathematics, 469(??):??, December 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704272500175X.

Koellermeier:2025:PIM

[KS25c] Julian Koellermeier and Giovanni Samaey. Projective integration methods in the Runge–Kutta framework and the extension to adaptivity in time. Journal of Computational and Applied Mathematics, 454(??):??, January 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042724003960.

Kong:2025:FRR

[KS25d] Lingtao Kong and Chenwei Song. Fuzzy robust regression based on exponentialtype kernel functions. Journal of Computational and Applied Mathematics, 457 (??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042724005430.

Katani:2025:ASM

[KSC25] Roghayeh Katani, Sedaghat Shahmorad, and Dajana Conte. Approximate solution of multi-term fractional differential equations via a block-by-block method. Journal of Computational and Applied Mathematics, 453 (??):??, January 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724003844.

Kouhanestani:2025:SIC

[KZG25] Mina Azizi Kouhanestani, Ehsan Zamanzade, and Sareh Goli. Statistical inference on the cumulative distribution function using judgment post stratification. Journal of Computational and Applied Mathematics, 458 (??):??, April 2025. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724005880.

Lancmanova:2025:EHO

[LB25] A. Lancmanová and T. Bodnár.

Effects of higher order discretization on the predictions of steady turbulent airflow in human respiratory system. Journal of Computational and Applied Mathematics, 466(??):??, October 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725001086.

Lakhdari:2025:FMN

[LBM25] Abdelghani Lakhdari, Djaber Chemseddine Benchettah, and Badreddine Meftah. Fractional multiplicative Newton-type inequalities for multiplicative s-convex positive functions with application. Journal of Computational and Applied Mathematics, 465 (??):??, September 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725001153.

Li:2025:ITA

[LC25a] Haifeng Li and Wengu Chen. An improved theoretical analysis of convolutional sparse coding using stripe coherence. Journal of Computational and Applied Mathematics, 469(??):??, December 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725001451.

Li:2025:ENT

[LC25b] Xuehua Li and Cairong Chen. An efficient Newton-type matrix splitting algorithm for solving generalized absolute value equations with application to ridge regression problems. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724005776.

Lovat:2025:LDE

[LC25c] Giampiero Lovat and Salvatore Celozzi. Lucas decomposition and extrapolation methods for the evaluation of infinite integrals involving the product of three Bessel functions of arbitrary order. Journal of Computational and Applied Mathematics, 453(??):??, January 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704272400390X.

Liu:2025:DNN

[LCR25] Min Liu, Zhiqiang Cai, and Karthik Ramani. Dual Neural Network (DuNN) method for elliptic partial differential equations and systems. Journal of Computational and Applied Mathematics, 467 (??):??, October 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725001116.

Li:2025:ACN

[LCZ25] Xiaolin Li, Xiyong Cui, and Shougui Zhang. Analysis of a Crank–Nicolson fast elementfree Galerkin method for the nonlinear complex Ginzburg– Landau equation. Journal of Computational and Applied Mathematics, 457 (??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042724005715.

Luo:2025:DUM

[LDZY25] Jiajie Luo, Wenting Duan, Min Zeng, and Yongxin Yuan. A direct updating method for undamped vibroacoustic FE models. Journal of Computational and Applied Mathematics, 463 (??):??, August 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042725000391.

Lv:2025:KRM

[LF25] Xianli Lv and Xiufang Feng. A kind regularization method for solving Cauchy problem of the Schrödinger equation. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042724004552.

Labarca-Figueroa:2025:CBV

[LFH25] Ignacio Labarca-Figueroa and Ralf Hiptmair. Coupled boundary and volume integral equations for electromagnetic scattering. Journal of Computational and Applied Mathematics, 461 (??):??, June 2025. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724006915.

Li:2025:UBO

[LG25] Bo Li and Qinglong Gao.

Uncertain bi-objective portfolio programming models of risky assets with liquidity and entropy constraints under uncertainty theory based DEA efficiency measures. Journal of Computational and Applied Mathematics, 461(??):??, June 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724006903.

Linghu:2025:HOM

[LGDN25] Jiale Linghu, Weifeng Gao, Hao Dong, and Yufeng Nie. Higher-order multi-scale physics-informed neural network (HOMS-PINN) method and its convergence analysis for solving elastic problems of authentic composite materials. Journal of Computational and Applied Mathematics, 456(??):??, March 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724004722.

Lin:2025:TPL

[LGZ25] Yaning Lin, Shuaichen Geng, and Weihai Zhang. Twoplayer LQ Stackelberg difference game with open-loop strategy under stochastic discrete-time dynamics in finite horizon. Journal of Computational and Applied Mathematics, 463(??):??, August 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042725000251.

Liu:2025:TSR

[LJY25] Wenli Liu, Jinbao Jian, and Jianghua Yin. A two-step relaxed-inertial derivativefree projection based algorithm for solving standard nonlinear pseudo-monotone equations and logistic regression problems. Journal of Computational and Applied Mathematics, 457 (??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042724005752.

Lazoglu:2025:OII

[LK25]  Çigdem Lazoglu and Ugur

Karabey. Optimizing insurance investments: the role of liquidity risk in assetliability management. Journal of Computational and Applied Mathematics, 470 (??):??, December 15, 2025.

CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725001505.

Li:2025:IMU

[LL25] Suhua Li and Chaoqian Li.

Invertible matrices under non-Archimedean absolute value and its inverse. Journal of Computational and Applied Mathematics, 470 (??):??, December 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725002146.

Liu:2025:GKM

[LLBX25] Li Liu, Wei-Guo Li, Wen-Di Bao, and Li-Li Xing. Greedy Kaczmarz methods for nonlinear equation. Journal of Computational and Applied Mathematics, 467 (??):??, October 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704272500144X.

Liu:2025:MLE

[LLG25] Zhe Liu, Yang Liu, and Rong Gao. Maximum likelihood estimation for multi-factor uncertain differential equations. Journal of Computational and Applied Mathematics, 454(??):??, January 15, 2025. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042724004540.

Lin:2026:AFE

[LLG26] Lei Lin, Junliang Lv, and Jiahui Gao. An adaptive finite element DtN method for the acoustic transmission problem. Journal of Computational and Applied Mathematics, 471(??):??, January 1, 2026. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042725002390.

Li:2025:CNF

[LLLL25] Sen Li, Ke Li, Yu Liu, and Qifeng Liao. A conditional normalizing flow for domain decomposed uncertainty quantification. Journal of Computational and Applied Mathematics, 465 (??):??, September 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704272500086X.

Liu:2025:GRK

[LLXB25] Li Liu, Wei-Guo Li, Li-Li Xing, and Wen-Di Bao. Greedy Randomized Kaczmarz with momentum method for nonlinear equation. Journal of Computational and Applied Mathematics, 459 (??):??, May 15, 2025. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724006071.

Li:2025:ETM

[LLZ25a] Xinyu Li, Ping Lin, and Weifeng Zhao. An efficient temporal multiscale algorithm for simulating a longterm plaque growth problem in relation to powerlaw blood flows. Journal of Computational and Applied Mathematics, 469 (??):??, December 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725001803.

Liu:2025:RCG

[LLZ25b] Hanbing Liu, Chongjun Li, and Yijun Zhong. Recovery conditions for generalized orthogonal matching pursuit based coherence. Journal of Computational and Applied Mathematics, 469 (??):??, December 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725001621.

Li:2025:SOL

[LLZH25] Yaoda Li, Min Li, Boya Zhou, and Zhibin Han. A second-order linear unconditionally energy-stable scheme for the phase field crystal equation. Journal of Computational and Applied Mathematics, 455(??):??, February 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042724004527.

Li:2025:HIM

[LM25a] Jingxuan Li and Hwan Pyo Moon. C^{1} Hermite interpolation method for septic PHoPH curves. Journal of Computational and Applied Mathematics, 464 (??):??, August 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042725000639.

Ling:2025:EDP

[LM25b] Dan Ling and Zhiping Mao. Efficient and dissipationpreserving Hermite spectral Galerkin methods for diffusive-viscous wave equations in unbounded domains. Journal of Computational and Applied Mathematics, 469(??):??, December 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042725001669.

Liu:2025:CAW

[LMZ25] Shicheng Liu, Xiangyun Meng, and Qilong Zhai.

Convergence analysis of a weak Galerkin finite element method on a Shishkin mesh for a singularly perturbed fourth-order problem in 2D. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724005727.

Lamwong:2025:FDM

[LP25a] Jiraporn Lamwong and Puntani Pongsumpun. A fractional derivative model of the dynamic of dengue transmission based on seasonal factors in Thailand. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724005053.

Li:2025:CAA

[LP25b] Hongru Li and Miltiadis V.

Papalexandris. Convergence analysis and applicability of a domain decomposition method with nonlocal interface boundary conditions. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042724005259.

Loris:2025:CAP

[LR25] Ignace Loris and Simone Rebegoldi. Convergence analysis of a primal-dual optimization-by-continuation algorithm. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042724005478.

Li:2025:TGR

[LRG25] Hongpeng Li, Hongxing Rui, and Mengqi Gao. Twogrid reduced-order method based on POD for a nonlinear poroelasticity model. Journal of Computational and Applied Mathematics, 460(??):??, May 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042724006617.

Lalhriatpuia:2025:SSF

[LS25a] H. Lalhriatpuia and Tanay Saha. Symmetric structured finite element model updating with prescribed partial eigenvalues while maintaining no spillover. Journal of Computational and Applied Mathematics, 470 (??):??, December 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725002122.

Lang:2025:VSI

[LS25b] Jens Lang and Bernhard A.

Schmitt. Variable-stepsize implicit Peer triplets in ODE constrained optimal control. Journal of Computational and Applied Mathematics, 460(??):??, May 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724006654.

Liu:2025:DIA

[LSCZ25] Xin Liu, Xianhua Song, Lili Chen, and Yanfeng Zhao. Distributed Ishikawa algorithms for seeking the fixed points of multi-agent global operators over time-varying communication graphs. Journal of Computational and Applied Mathematics, 457 (??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724004990.

Liu:2025:DRE

[LSLL25] Zhan Liu, Yi Sun, Yong Li, and Yuanmeng Li. Doubly robust estimation for nonprobability samples with heterogeneity. Journal of Computational and Applied Mathematics, 465(??):??, September 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725000822.

Los:2025:SIF

[LSPV25] Marcin Loś, Tomasz Słuzalec, Maciej Paszyński, and Eirik Valseth. Stabilization of isogeometric finite element method with optimal test functions computed from L_{2} norm residual minimization. Journal of Computational and Applied Mathematics, 460(??):??, May 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724006587.

Liu:2025:ADS

[LTL^{+}25] J. K. Liu, B. Tang, T. Liu, Z. T. Yang, and S. Liang. An accelerated double-step derivative-free projection method based algorithm using Picard– Mann iterative process for solving convex constrained nonlinear equations. Journal of Computational and Applied Mathematics, 464 (??):??, August 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725000561.

Li:2025:BEA

[LTZ25] Haibo Li, Guangming Tan, and Tong Zhao. Backward error analysis of the Lanczos bidiagonalization with reorthogonalization. Journal of Computational and Applied Mathematics, 460(??): ??, May 1, 2025. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724006629.

Lopez-Urena:2025:AFE

[LU25] Sergio López-Ureña. Activation functions enabling the addition of neurons and layers without altering outcomes. Journal of Computational and Applied Mathematics, 470(??):??, December 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725002444.

Li:2025:ADG

[LW25] Xiaosheng Li and Wei Wang.

An analysis of discontinuous Galerkin method for Electrical Impedance Tomography with partial data. Journal of Computational and Applied Mathematics, 459(??):??, May 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724006241.

Li:2025:NMF

[LWA25] Xuhao Li, Patricia J. Y. Wong, and Anatoly A. Alikhanov. Numerical method for fractional sub-diffusion equation with space-time varying diffusivity and smooth solution. Journal of Computational and Applied Mathematics, 464(??):??, August 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042724007210.

Liu:2025:RRT

[LWF25a] Kai Liu, Bin Wang, and Ting Fu. Relaxation RKN-type integrators that preserve two invariants for second-order (oscillatory) systems. Journal of Computational and Applied Mathematics, 457 (??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704272400548X.

Liu:2025:STS

[LWF25b] Wei Liu, Pengshan Wang, and Gexian Fan. A spacetime second-order algorithm based on finite volume method for Brinkman flow and reactive transport model in porous media with variable fractures. Journal of Computational and Applied Mathematics, 462(??):??, July 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724007167.

Li:2025:FFI

[LWH^{+}25] Xi’an Li, Jinran Wu, Yujia Huang, Zhe Ding, Xin Tai, Liang Liu, and You-Gan Wang. Fourier-feature induced physics informed randomized neural network method to solve the biharmonic equation. Journal of Computational and Applied Mathematics, 468(??): ??, November 2025. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725001499.

Li:2025:CBS

[LWL^{+}25] Yaomeng Li, Feng Wang, Qiao Li, Chao Fu, and Xu Guo. Cubic B-spline based elastic and viscoelastic wave propagation method. Journal of Computational and Applied Mathematics, 456(??):??, March 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724004850.

Li:2025:MSM

[LWLW25] Dandan Li, Songhua Wang, Yong Li, and Jiaqi Wu. A modified SGP method with two alternative spectral parameters for constrained nonlinear equations with a compressive sensing application. Journal of Computational and Applied Mathematics, 463(??):??, August 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725000184.

Lu:2025:DES

[LWZZ25] Nan Lu, Chenxi Wang, Lun Zhang, and Zhen Zhang. Decoupled and energy stable schemes for phase-field surfactant model based on mobility operator splitting technique. Journal of Computational and Applied Mathematics, 459(??):??, May 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724006137.

Long:2025:SBG

[LY25] Xian-Jun Long and Jing Yang. A stochastic Bregman golden ratio algorithm for non-Lipschitz stochastic mixed variational inequalities with application to resource share problems. Journal of Computational and Applied Mathematics, 459(??):??, May 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724006290.

Li:2025:ICE

[LYC25] Zishang Li, Changqing Ye, and Eric T. Chung. An iterative constraint energy minimizing generalized multiscale finite element method for contact problem. Journal of Computational and Applied Mathematics, 461 (??):??, June 2025. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724006460.

Lancinskas:2025:PBA

[LZFP25] Algirdas Lancinskas, Julius Zilinskas, Pascual Fernández, and Blas Pelegrín. Populationbased algorithm for discrete facility location with ranking of candidate locations. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724005521.

Li:2025:VIR

[LZZ25a] Jingshi Li, Jiachuan Zhang, and Ran Zhang. Viscosityindependent reliability and efficiency constants for Brinkman equations with stabilized Crouzeix–Raviart element. Journal of Computational and Applied Mathematics, 460(??):??, May 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042724006599.

Lv:2025:RIA

[LZZ^{+}25b] Lingling Lv, Jiali Zhao, Bingqian Zheng, Jianwei Shen, and Huaicheng Yan. A recursive identification algorithm for discrete timedelay periodic linear systems. Journal of Computational and Applied Mathematics, 461(??):??, June 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042724006952.

Mohssine:2025:ESC

[MAHL25] A. Mohssine, L. Afraites, A. Hadri, and A. Laghrib. Existence of solution for a coupled diffusion PDE system for various noise reduction. Journal of Computational and Applied Mathematics, 463(??):??, August 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042725000366.

Meng:2025:IAC

[MCL25] Xiaoqing Meng, Aijie Cheng, and Zhengguang Liu. Improving the accuracy and consistency of the energy quadratization method with an energy-optimized technique. Journal of Computational and Applied Mathematics, 459(??):??, May 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724006162.

Mosic:2025:NEC

[MF25] D. Mosić and D. E. Ferreyra.

A new extension of the core inverse. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724005533.

Mezzadri:2025:MBM

[MG25] Francesco Mezzadri and Emanuele Galligani. Modulusbased matrix splitting algorithms for generalized complex-valued horizontal linear complementarity problems. Journal of Computational and Applied Mathematics, 460(??):??, May 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724006885.

Martin:2025:GPB

[MHO25] Dirk Martin, Gundolf Haase, and Günter Offner. Generic preconditioning based on the inverse with respect to the semi-inner product for iterative solvers for radial basis function interpolation. Journal of Computational and Applied Mathematics, 459(??):??, May 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724006034.

Martinez:2025:ISD

[MHP25] D. Martínez, H. Herrero, and F. Pla. Implicit Schwarz domain decomposition method for a Rayleigh–Bénard problem. Journal of Computational and Applied Mathematics, 470(??):??, December 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725002328.

Ma:2025:SMO

[MHWY25] Mengrao Ma, Linmin Hu, Lin Wang, and Xiaoyun Yu. Stochastic modeling and optimization of discrete-time cold standby repairable systems with unreliable repair facility and retrial mechanism. Journal of Computational and Applied Mathematics, 453(??):??, January 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724003832.

Moghaddamfar:2026:ERG

[MKE26] Maryam Moghaddamfar, Manochehr Kazemi, and Reza Ezzati. Existence results for generalized 2D fractional partial integro-differential equations. Journal of Computational and Applied Mathematics, 471(??):??, January 1, 2026. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725002195.

Macias:2025:FGN

[MMD25] Siegfried Macías and Jorge E.

Macías-Díaz. On a fractional generalization of a nonlinear model in plasma physics and its numerical resolution via a multi-conservative and efficient scheme. Journal of Computational and Applied Mathematics, 462 (??):??, July 2025. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724007222.

Marynets:2025:BVP

[MMK25] Vasyl Marynets, Kateryna Marynets, and Oksana Kohutych. Boundary value problems modeling moisture transport in soils. Journal of Computational and Applied Mathematics, 465 (??):??, September 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042725001128.

Murad:2025:OSD

[MO25] Muhammad Amin S. Murad and Faraj M. Omar. Optical solitons, dynamics of bifurcation, and chaos in the generalized integrable (2+1)- dimensional nonlinear conformable Schrödinger equations using a new Kudryashov technique. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042724005466.

Mohammed:2025:DML

[Moh25] Pshtiwan Othman Mohammed. On the delta Mittag-Leffler functions and its application in monotonic analysis. Journal of Computational and Applied Mathematics, 465(??):??, September 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042725000809.

Moisa:2025:TOT

[Moi25] Andrew Moisa. Third order two-step Runge–Kutta– Chebyshev methods. Journal of Computational and Applied Mathematics, 457 (??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724005399.

Masirevic:2025:ISR

[MPB25] Dragana Jankov Masirević, Tibor K. Pogány, and Tomislav Burić. Integral and series representations of the Nuttall function. Journal of Computational and Applied Mathematics, 453(??):??, January 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724003881.

Minakowski:2023:PPE

[MR23] P. Minakowski and T. Richter.

A priori and a posteriori error estimates for the Deep Ritz method applied to the Laplace and Stokes problem. Journal of Computational and Applied Mathematics, 421(??):??, March 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042722004435.

See erratum [MR25].

Minakowski:2025:EPP

[MR25] P. Minakowski and T. Richter.

Erratum to “A priori and a posteriori error estimates for the Deep Ritz method applied to the Laplace and Stokes problem” [J. Comput. Appl. Math. 421 (2023) 114845]. Journal of Computational and Applied Mathematics, 460(??):??, May 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704272400654X.

See [MR23].

Mahata:2025:GCN

[MRRÁY25] Shipra Mahata, Samala Rathan, Juan Ruiz-Álvarez, and Dionisio F. Yáñez. A general correction for numerical integration rules over piece-wise continuous functions. Journal of Computational and Applied Mathematics, 459(??):??, May 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724006265.

Ma:2025:NMB

[MS25a] Zheng Ma and Martin Stynes. A Nyström method based on product integration for weakly singular Volterra integral equations with variable exponent. Journal of Computational and Applied Mathematics, 454(??):??, January 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724004138.

Muduli:2025:ASG

[MS25b] Tapan Kumar Muduli and Purnima Satapathy. Analytical solutions for the generalized (2 + 1)-dimensional Konopelchenko–Dubrovsky equation via Lie symmetry analysis. Journal of Computational and Applied Mathematics, 465(??):??, September 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725000792.

Maurya:2026:OPU

[MSR26] Vikas Maurya, Ankit Singh, and Manoj K. Rajpoot. Option pricing using the IMEX-AVF method with high jump intensity. Journal of Computational and Applied Mathematics, 471(??):??, January 1, 2026. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725002730.

Masoodi:2025:OUS

[MT25] Mehdi Masoodi and Haniyeh Taromideh. Optimal and uniform sensor arrangement in near-field imaging. Journal of Computational and Applied Mathematics, 454 (??):??, January 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042724004370.

Mao:2025:AOP

[MWTW25] Mengli Mao, Wansheng Wang, Tianhai Tian, and Lehan Wang. Adaptive option pricing based on a posteriori error estimates for fully discrete finite difference methods. Journal of Computational and Applied Mathematics, 460(??):??, May 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042724006551.

Meng:2025:NMS

[MXW^{+}25] Pinchao Meng, Zhaobin Xu, Xianchao Wang, Weishi Yin, and Hongyu Liu. A novel method for solving the inverse spectral problem with incomplete data. Journal of Computational and Applied Mathematics, 463 (??):??, August 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042725000408.

Ma:2025:AGM

[MY25] Changfeng Ma and Xiaojuan Yu. Analysis of a generalized MASSOR method for saddle point problems. Journal of Computational and Applied Mathematics, 467 (??):??, October 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725001414.

Ma:2025:TIR

[MYL25] Xuejie Ma, Sixing Yang, and Pengjie Liu. Two inertial-relaxed hybridized CG projection-based algorithms for solving nonlinear monotone equations applied in image restoration. Journal of Computational and Applied Mathematics, 464 (??):??, August 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725000615.

Miao:2025:NFT

[MZL25] Peng Miao, Daoyuan Zhang, and Shuai Li. A novel fixed-time zeroing neural network and its application to path tracking control of wheeled mobile robots. Journal of Computational and Applied Mathematics, 460(??):??, May 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724006502.

Nik:2025:BCR

[NAF25] A. Saadati Nik, A. Asgharzadeh, and A. J. Fernández.

Bayesian credible regions for two-parameter exponential distributions under type-II censoring. Journal of Computational and Applied Mathematics, 470(??):??, December 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725002353.

Nandy:2025:RIL

[NBG25] Amarnath Nandy, Ayanendranath Basu, and Abhik Ghosh. Robust inference for linear regression models with possibly skewed error distribution. Journal of Computational and Applied Mathematics, 463(??):??, August 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725000172.

Niu:2025:TAA

[NDSZ25] Jing Niu, Lei Du, Tomohiro Sogabe, and Shao-Liang Zhang. A tensor Alternating Anderson–Richardson method for solving multilinear systems with Mtensors. Journal of Computational and Applied Mathematics, 461(??):??, June 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724006678.

Nawaz:2025:EPA

[NGU25a] Bashir Nawaz, Krzysztof Gdawiec, and Kifayat Ullah. Effectiveness of Picard– Abbas iteration for fixed point approximation of mappings satisfying condition (E). Journal of Computational and Applied Mathematics, 469(??):??, December 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725001633.

Nguyen:2025:RQI

[Ngu25b] Du Nguyen. Rayleigh Quotient Iteration, cubic convergence, and second covariant derivative. Journal of Computational and Applied Mathematics, 463(??):??, August 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725000378.

Narbek:2025:CHP

[NKN25] Orazbek Narbek, Samat A.

Kassabek, and Targyn Nauryz. A collocation heat polynomials method for onedimensional inverse Stefan problems. Journal of Computational and Applied Mathematics, 460(??):??, May 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724006046.

Nawa:2025:LMM

[NN25] Victor Nawa and Saralees Nadarajah. Logarithmic method of moments estimators for the Fréchet distribution. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042724005417.

Nikiforov:2025:MGM

[NPA^{+}25] Djulustan Nikiforov, Leonardo A. Poveda, Dmitry Ammosov, Yesy Sarmiento, and Juan Galvis. Meshfree generalized multiscale exponential integration method for parabolic problems. Journal of Computational and Applied Mathematics, 459(??):??, May 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042724006150.

Navarro:2025:KDP

[NRS25] Jorge Navarro, Tomasz Rychlik, and Magdalena Szymkowiak. Key distributions in the preservation of aging classes under the construction of systems. Journal of Computational and Applied Mathematics, 469(??):??, December 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725001645.

Niu:2025:NMS

[NSW25] Qiang Niu, Binrui Shen, and Yenan Wang. Numerical methods for solving a class of matrix equations arising from inference for ranked set sampling on imperfect ranking. Journal of Computational and Applied Mathematics, 463(??):??, August 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725000342.

Notaris:2025:NPI

[NT25a] Sotirios E. Notaris and Nikolaos J. Theodorakopoulos. Nonsymmetric product integration rules for Chebyshev weight functions with Chebyshev abscissae. Journal of Computational and Applied Mathematics, 469 (??):??, December 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725001827.

Novo:2025:CNN

[NT25b] Julia Novo and Eduardo Terrés. Can neural networks learn finite elements? Journal of Computational and Applied Mathematics, 453 (??):??, January 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042724004175.

Niu:2025:RBK

[NZ25] Yu-Qi Niu and Bing Zheng. On the randomized block Kaczmarz algorithms for solving matrix equation AXB = C. Journal of Computational and Applied Mathematics, 460(??):??, May 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042724006691.

Oyewole:2025:ISE

[OAO25] Olawale K. Oyewole, Hammed A. Abass, and Olufemi J. Ogunsola. An improved subgradient extragradient selfadaptive algorithm based on the golden ratio technique for variational inequality problems in Banach spaces. Journal of Computational and Applied Mathematics, 460(??):??, May 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704272400668X.

Orynyak:2025:TDE

[OTD^{+}25] Igor Orynyak, Anton Tsybulnyk, Kirill Danylenko, Andrii Oryniak, and Sergii Radchenko. Timestep-dependent element interpolation functions in the method of matched sections on the example of heat conduction problem. Journal of Computational and Applied Mathematics, 456(??):??, March 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724004710.

Ozarslan:2025:AFC

[Öza25] Mehmet Ali Özarslan. On the approximation to fractional calculus operators with multivariate Mittag-Leffler function in the kernel. Journal of Computational and Applied Mathematics, 454 (??):??, January 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724003972.

Pecile:2025:DDD

[PDT^{+}25] Alessandro Pecile, Nicola Demo, Marco Tezzele, Gianluigi Rozza, and Dimitri Breda. Data-driven discovery of delay differential equations with discrete delays. Journal of Computational and Applied Mathematics, 461(??):??, June 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724006873.

Pan:2025:FDS

[PFYH25] Mingyang Pan, Chengxing Fu, Liu Yang, and Dongdong He. A fullydecoupled, second-order accurate, positivity-preserving and energy stable scheme for a two-phase flow system with ions transport. Journal of Computational and Applied Mathematics, 465(??):??, September 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042725000883.

Phong:2025:RCH

[Pho25] Duong Thanh Phong. Representation computation for the hypergeometric function of a Hermitian matrix argument. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042724005077.

Paczek:2025:CCE

[PJPW25] Kewin Paczek, Damian Jelito, Marcin Pitera, and Agnieszka Wyłomańska. Conditional correlation estimation and serial dependence identification. Journal of Computational and Applied Mathematics, 468 (??):??, November 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725001475.

Polyanin:2025:ESR

[PK25] Andrei D. Polyanin and Nikolay A. Kudryashov. Exact solutions and reductions of nonlinear Schrödinger equations with delay. Journal of Computational and Applied Mathematics, 462(??):??, July 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724007258.

Pu:2025:FAB

[PLW^{+}25] Rong-Fang Pu, Liang Li, Qin Wang, Zhao-Yu Lu, and Li-Hong Liao. FastAGM-Gar: an aggregation-based algebraic multigrid method. Journal of Computational and Applied Mathematics, 463(??):??, August 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725000305.

Park:2025:DGM

[PM25a] Do Yang Park and Minam Moon. Detective generalized multiscale hybridizable discontinuous Galerkin (GMsHDG) method for porous media. Journal of Computational and Applied Mathematics, 457(??):

??, March 15, 2025. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042724005685.

Piterskaya:2025:SOS

[PM25b] Anna Piterskaya and Mikael Mortensen. A study of the Orr–Sommerfeld and induction equations by Galerkin and Petrov–Galerkin spectral methods utilizing Chebyshev polynomials. Journal of Computational and Applied Mathematics, 459(??): ??, May 15, 2025. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042724006228.

Prokopev:2025:SRI

[PNL25] Sergei Prokopev, Alexander Nepomnyashchy, and Tatyana Lyubimova. The spectral radius of iterative methods for the Cahn– Hilliard equation and its relation to the splitting technique. Journal of Computational and Applied Mathematics, 470(??):??, December 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042725001876.

Power:2025:ARP

[PP25] Jenny Power and Tristan Pryer. Adaptive regularisation for PDE-constrained optimal control. Journal of Computational and Applied Mathematics, 470(??): ??, December 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725001657.

Priyanka:2025:SDF

[PS25] Kumari Priyanka and A. Antony Selvan. Sampling with derivatives in finitely generated shift-invariant spaces. Journal of Computational and Applied Mathematics, 470 (??):??, December 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725002067.

Pelosi:2025:CPM

[PSF25] Francesca Pelosi, Maria Lucia Sampoli, and Rida T. Farouki. Control point modifications that preserve the Pythagorean-hodograph nature of planar quintic curves. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724005491.

Peyrovan:2026:CMV

[PT26] P. Peyrovan and A. Tari. Collocation method for Volterra integro-differential equations with piecewise delays. Journal of Computational and Applied Mathematics, 471 (??):??, January 1, 2026. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042725002687.

Peng:2025:ARM

[PW25] Xingchun Peng and Yankai Wang. Alpha-robust meanvariance reinsurance and investment strategies with transaction costs. Journal of Computational and Applied Mathematics, 457 (??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042724005065.

Pei:2025:OEE

[PXZZ25] Lifang Pei, Chao Xu, Jiwei Zhang, and Yanmin Zhao. Optimal error estimations and superconvergence analysis of anisotropic FEMs with variable time steps for reaction–diffusion equations. Journal of Computational and Applied Mathematics, 469(??):??, December 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725001700.

Peng:2025:ALE

[PY25] Gang Peng and Di Yang.

An arbitrary Lagrangian– Eulerian positivity-preserving finite volume scheme for radiation hydrodynamics equations in the equilibriumdiffusion limit. Journal of Computational and Applied Mathematics, 455(??): ??, February 2025. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724004059.

Qiao:2025:SHA

[QSW25] Yan Qiao, Lin Sang, and Hua Wu. A stable and highaccuracy numerical method for determining the timedependent coefficient in the bioheat equation. Journal of Computational and Applied Mathematics, 464 (??):??, August 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725000433.

Qiu:2025:DME

[QWLF25] Wensong Qiu, Hongyan Wang, Yuan Li, and Lixin Feng. Determination of the modified exterior Steklov eigenvalues via the reciprocity gap method. Journal of Computational and Applied Mathematics, 459(??): ??, May 15, 2025. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724006083.

Qin:2025:IRP

[QWZ25] Mingze Qin, Xiuli Wang, and Huifang Zhou. Invariant region property of weak Galerkin method for semilinear parabolic equations. Journal of Computational and Applied Mathematics, 460(??):??, May 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724006605.

Qin:2025:HMD

[QZG^{+}25] Rong Qin, Jiansong Zhang, Xinxin Guo, Hui Guo, and Jiang Zhu. Hybrid mixed discontinuous Galerkin finite element analysis of time-dependent convectiondiffusion equations and its application in chemotaxis model. Journal of Computational and Applied Mathematics, 465(??):??, September 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725000858.

Ranjbar:2025:EES

[Ran25] Hassan Ranjbar. An explicit Euler scheme for generalized n-dimensional secondorder differential equations with initial value conditions driven by additive Gaussian white noises. Journal of Computational and Applied Mathematics, 454(??):??, January 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704272400428X.

Roul:2025:HON

[RK25] Pradip Roul and Trishna Kumari. High-order numerical schemes based on B-spline for solving a time-fractional Fokker–Planck equation. Journal of Computational and Applied Mathematics, 460 (??):??, May 1, 2025. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724006344.

Ruby:2025:SRF

[RM25] Ruby and Moumita Mandal.

Superconvergent results for fractional Volterra integrodifferential equations with non-smooth solutions. Journal of Computational and Applied Mathematics, 458 (??):??, April 2025. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042724005855.

Robert:2025:MYR

[RNC25] Maria Robert, Suresh Kumar Nadupuri, and Nagaiah Chamakuri. Moreau– Yosida regularization to optimal control of the monodomain model with pointwise control and state constraints in cardiac electrophysiology. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042724005545.

Romani:2025:ACP

[RV25] Lucia Romani and Alberto Viscardi. Algebraic characterization of planar cubic and quintic Pythagorean-Hodograph B-spline curves. Journal of Computational and Applied Mathematics, 465(??):??, September 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042725001074.

Saha:2025:IEP

[Sah25] Tanay Saha. Inverse eigenvalue problem for affinely structured or sparse polynomial matrices with prescribed eigenvalues. Journal of Computational and Applied Mathematics, 454(??):??, January 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724004564.

Said:2025:IGR

[Sai25] Samia M. Said. Influence of gravitational on a rotating nonlocal thermoelastic medium with thermal variable conductivity. Journal of Computational and Applied Mathematics, 454 (??):??, January 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724004291.

Saksa:2025:CFE

[Sak25] Tytti Saksa. Comparison of finite element and discrete exterior calculus in computation of time-harmonic wave propagation with controllability. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724004035.

Skowronek:2025:TEI

[SAP^{+}25] Katarzyna Skowronek, Marek Arendarczyk, Anna K. Panorska, Tomasz J. Kozubowski, and Agnieszka Wyłomańska.

Testing and estimation of the index of stability of univariate and bivariate symmetric α-stable distributions via modified Greenwood statistic. Journal of Computational and Applied Mathematics, 467(??):??, October 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725001025.

Stoyanovskaya:2025:DAS

[SBAM25] O. P. Stoyanovskaya, O. A. Burmistrova, M. S. Arendarenko, and T. V. Markelova. Dispersion analysis of SPH for parabolic equations: High-order kernels against tensile instability. Journal of Computational and Applied Mathematics, 457 (??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724005648.

Shen:2025:RCL

[SC25] Zixiang Shen and Guo Chen. Robust H_{∞} control for LFC of discrete T–S fuzzy MAPS with DFIG and time-varying delays. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704272400520X.

Shekhar:2025:RAI

[SDS25] Chandra Shekhar, Mahendra Devanda, and Keshav Sharma. Reliability analysis of imperfect repair and switching failures: a Bayesian inference and Monte Carlo simulation approach. Journal of Computational and Applied Mathematics, 461 (??):??, June 2025. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724007064.

Singh:2025:HEO

[SGRS25] Randhir Singh, Vandana Guleria, Higinio Ramos, and Mehakpreet Singh. Highly efficient optimal decomposition approach and its mathematical analysis for solving fourth-order Lane–Emden– Fowler equations. Journal of Computational and Applied Mathematics, 456 (??):??, March 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724004874.

Sugiura:2025:EBG

[SH25] Hiroshi Sugiura and Takemitsu Hasegawa. Error bounds for Gauss–Lobatto quadrature of analytic functions on an ellipse. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042724005740.

Shi:2026:DQC

[Shi26]  Jianxin Shi.  Dynamics

of quantum Cournot model under strategic delegation. Journal of Computational and Applied Mathematics, 471(??):??, January 1, 2026. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042725002663.

Shadimetov:2025:OQF

[SHK25] Kholmat Shadimetov, Abdullo Hayotov, and Umedjon Khayriev. Optimal quadrature formulas for approximating strongly oscillating integrals in the Hilbert space W̃_{2}(m, m − 1) of periodic functions. Journal of Computational and Applied Mathematics, 453(??):??, January 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042724003820.

Shen:2025:WDB

[SHL^{+}25] Zhao-Li Shen, Guo-Liang Han, Yu-Tong Liu, Bruno Carpentieri, Chun Wen, and Jian-Jun Wang. Weak dangling block reordering and multi-step block compression for efficiently computing and updating PageRank solutions. Journal of Computational and Applied Mathematics, 458(??):??, April 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724005806.

Saah:2025:NCE

[SK25a] David K. Saah and Tomasz J.

Kozubowski. A new class of extended Laplace distributions with applications to modeling contaminated Laplace data. Journal of Computational and Applied Mathematics, 465(??): ??, September 2025. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725001037.

Sabetrasekh:2025:LQB

[SK25b] Maryam Sabetrasekh and Iraj Kazemi. Longitudinal quantile-based regression models using multivariate asymmetric heavy-tailed distributions and leapfrog HMC algorithm. Journal of Computational and Applied Mathematics, 470 (??):??, December 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042725002043.

Singh:2025:EWM

[SKM25] Akanksha Singh, Ankur Kanaujiya, and Jugal Mohapatra. Euler wavelets method for optimal control problems of fractional integrodifferential equations. Journal of Computational and Applied Mathematics, 454 (??):??, January 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042724004278.

Singh:2026:EGL

[SKM26] Akanksha Singh, Ankur Kanaujiya, and Jugal Mohapatra. An efficient generalized Lucas wavelet operational matrix method for optimal control problems. Journal of Computational and Applied Mathematics, 471 (??):??, January 1, 2026. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042725002638.

Suppalap:2025:RNA

[SMDW25] Siwakon Suppalap, Dawrawee Makmuang, Vipavee Damminsed, and Rabian Wangkeeree. Regularized Nesterov’s accelerated damped BFGS method for stochastic optimization. Journal of Computational and Applied Mathematics, 467(??):??, October 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725001311.

Shen:2025:LGM

[SNZ25] Binrui Shen, Qiang Niu, and Shengxin Zhu. Lightning graph matching. Journal of Computational and Applied Mathematics, 454 (??):??, January 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724004382.

Song:2025:MDM

[Son25]  Kai Song.  Multivariate

degradation modeling and reliability evaluation using gamma processes with hierarchical random effects. Journal of Computational and Applied Mathematics, 465(??):??, September 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725001062.

Shang:2025:NPF

[SQLZ25] Youlin Shang, Deqiang Qu, Junxiang Li, and Roxin Zhang. A new parameterfree continuously differentiable filled function algorithm for solving nonlinear equations and data fitting problems. Journal of Computational and Applied Mathematics, 454(??):??, January 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724004473.

Serna-Reyes:2025:CTS

[SRMGMD25] Adán J. Serna-Reyes, Siegfried Macías, Armando Gallegos, and Jorge E. Macías-Díaz. A convergent two-step method to solve a fractional extension of the Rosenau–Kawahara system. Journal of Computational and Applied Mathematics, 460(??):??, May 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724006721.

Singh:2025:EFO

[SS25] Shubham Singh and Shalini. An equivalent fractional optimization problem under invexity conditions for solving fractional vector variationallike inequality problems to portfolio optimization in finance. Journal of Computational and Applied Mathematics, 470(??):??, December 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725001980.

Shen:2025:PTA

[SSLP25] Haoxuan Shen, Vladimir Simic, Shuguang Li, and Dragan Pamucar. Polynomial time algorithms to find Pareto optimal schedules of bicriteria lot scheduling problems with splitable jobs on a single parallel-batch machine. Journal of Computational and Applied Mathematics, 459(??):??, May 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724006289.

Spiridonov:2025:PDO

[SSM25] Denis Spiridonov, Sergei Stepanov, and Tina Mai. Prediction of discretization of online GMsFEM using deep learning for Richards equation. Journal of Computational and Applied Mathematics, 454(??):??, January 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724004163.

Sun:2025:EAN

[SSW^{+}25] Guanglei Sun, Youlin Shang, Xiaoqiang Wang, Roxin Zhang, and Deqiang Qu. An efficient algorithm via a novel one-parameter filled function based on general univariate functions for unconstrained global optimization. Journal of Computational and Applied Mathematics, 468(??):??, November 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042725001463.

Shi:2025:UTV

[SW25] Yuxin Shi and Lu Wang. Uncertain time-varying autoregressive model with imprecise observations and its applications. Journal of Computational and Applied Mathematics, 463(??):??, August 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042725000238.

Sun:2025:GEB

[SWD25] Hongchun Sun, Yiju Wang, and Jiakang Du. Global error bound estimates algorithm for an R_{0}-type generalized LCP over polyhedral cone and its applications. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042724005363.

Song:2025:EAT

[SWWY25] Haiming Song, Hao Wang, Jiageng Wu, and Jinda Yang. An efficient ADAMtype algorithm with finite elements discretization technique for random elliptic optimal control problems. Journal of Computational and Applied Mathematics, 454 (??):??, January 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724004485.

Shah:2025:ASC

[SX25] Syed Omar Shah and Yonghui Xia. An analysis on the stability and controllability of non-linear measure differential equations. Journal of Computational and Applied Mathematics, 453 (??):??, January 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724003893.

Shi:2025:REM

[SY25] Yanjie Shi and Zaizai Yan.

Reliability evaluation of multi-state (k1, k2, …, km)- out-of-(n1, n2, …, nm) system with common bus performance sharing. Journal of Computational and Applied Mathematics, 470 (??):??, December 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725002183.

Song:2025:FFM

[SZD25] Qiu-Yan Song, Umair Zulfiqar, and Xin Du. Finitefrequency model order reduction of linear and bilinear systems via low-rank approximation. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042724005351.

Tuncel:2025:SBS

[TA25] Altan Tuncel and Tugba Aktas Aslan. Some bivariate Schur-constant distributions and application to life insurance. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042724005442.

Tan:2025:SAE

[TC25] Ting Tan and Waixiang Cao. Spectral approximation and error analysis for the transmission eigenvalue problem with an isotropic inhomogeneous medium. Journal of Computational and Applied Mathematics, 453 (??):??, January 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724004126.

Tan:2025:ADB

[TGDC25] Longze Tan, Xueping Guo, Mingyu Deng, and Jingrun Chen. On the adaptive deterministic block Kaczmarz method with momentum for solving large-scale consistent linear systems. Journal of Computational and Applied Mathematics, 457 (??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724005764.

Tang:2025:STL

[TL25] Ling Tang and Hanyu Li.

Subsampling for tensor least squares: Optimization and statistical perspectives. Journal of Computational and Applied Mathematics, 470 (??):??, December 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725002080.

Tong:2025:ALM

[TLSW25] Shanshan Tong, Muyi Liu, Hongying Shu, and PanPan Wang. Adaptive Levenberg– Marquardt–Kaczmarz method for nonlinear ill-posed problems with non-smooth constraints. Journal of Computational and Applied Mathematics, 463(??):??, August 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042725000196.

Tong:2025:LSR

[Ton25] Hongzhi Tong. Least squares regression under weak moment conditions. Journal of Computational and Applied Mathematics, 458 (??):??, April 2025. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042724005843.

Tang:2025:GCP

[TSL25] Jie Tang, Zhou Sheng, and Xin Li. A generalized complex power iteration-based algorithm for quartic-quadratic optimization problems over a spherical constraint and its applications. Journal of Computational and Applied Mathematics, 470(??): ??, December 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042725002304.

Treanta:2025:RSE

[TSM25] Savin Treanta, Vinay Singh, and Shashi Kant Mishra. Reciprocal solution existence results for a class of vector variational control inequalities with application in physics. Journal of Computational and Applied Mathematics, 461(??):??, June 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704272400709X.

Tschimpke:2025:RRD

[TST25] Marco Tschimpke, Manuela Schreyer, and Wolfgang Trutschnig. Revisiting the region determined by Spearman’s ρ and Spearman’s footrule φ. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724005089.

Tang:2025:ASC

[TTZL25] Chunming Tang, Wancheng Tan, Yongshen Zhang, and Zhixian Liu. An accelerated spectral CG based algorithm for optimization techniques on Riemannian manifolds and its comparative evaluation. Journal of Computational and Applied Mathematics, 462(??):??, July 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724007301.

Tang:2025:IPP

[TYJH25] Xianke Tang, Jianghua Yin, Dan Jian, and Daolan Han. An indefinite proximal Peaceman–Rachford splitting method-based algorithm integrating the generalization acceleration technique for separable convex programming problems in image restoration. Journal of Computational and Applied Mathematics, 470(??):??, December 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042725002079.

Tan:2025:TSO

[TZ25a] Zhijun Tan and Yunhua Zeng. Temporal secondorder two-grid finite element method for semilinear timefractional Rayleigh–Stokes equations. Journal of Computational and Applied Mathematics, 459(??):??, May 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704272400623X.

Tang:2025:LMT

[TZ25b] Jingyong Tang and Jinchuan Zhou. A Levenberg–Marquardt type algorithm with a Broydenlike update technique for solving nonlinear equations. Journal of Computational and Applied Mathematics, 460(??):??, May 1, 2025.

CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724006496.

Junjua:2026:RFO

[uDJKA26] Moin ud Din Junjua, Sunil Kumar, and Rashid Ali. A robust family of optimal fourthorder iteration schemes for multiple roots with applications. Journal of Computational and Applied Mathematics, 471(??):??, January 1, 2026. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725002729.

urRehman:2025:SEF

[uRGYZ25] Habib ur Rehman, Debdas Ghosh, Jen-Chih Yao, and Xiaopeng Zhao. Solving equilibrium and fixed-point problems in Hilbert spaces: a class of strongly convergent Mann-type dual-inertial subgradient extragradient methods. Journal of Computational and Applied Mathematics, 464(??):??, August 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704272500024X.

Ullah:2025:CSH

[UX25] Inayat Ullah and Yingxiang Xu. Computational study of homogeneous–heterogeneous reactions in bioconvective third-grade material flow. Journal of Computational and Applied Mathematics, 454(??):??, January 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724004308.

Vasilyeva:2025:GMF

[Vas25]  Maria Vasilyeva.  Gener-

alized multiscale finite element method for discrete network (graph) models. Journal of Computational and Applied Mathematics, 457 (??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724005247.

Varshney:2025:HRA

[VKA25] Megha Varshney, Pravesh Kumar, and Laith Abualigah. Hybridizing Remora and Aquila Optimizer with dynamic oppositional learning for structural engineering design problems. Journal of Computational and Applied Mathematics, 462 (??):??, July 2025. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724007234.

Vasilyeva:2025:MMI

[VKGM25] Maria Vasilyeva, Aleksei Krasnikov, Kelum Gajamannage, and Mehrube Mehrubeoglu. Multiscale method for image denoising using nonlinear diffusion process: Local denoising and spectral multiscale basis functions. Journal of Computational and Applied Mathematics, 470(??):??, December 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704272500247X.

Villa-Morales:2025:OCL

[VMRBG25] José Villa-Morales, Luis Rincón, and Gerardo Becerra-Guzmán. Optimal control of linear fractional differential equations in the Caputo sense. Journal of Computational and Applied Mathematics, 470(??):??, December 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725002213.

Vulanovic:2025:AST

[VN25] Relja Vulanović and Thái Anh Nhan. Advantages of the Samarskii-type schemes on the Shishkin mesh. Journal of Computational and Applied Mathematics, 470 (??):??, December 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704272500202X.

Vacek:2025:FEM

[VS25] Karel Vacek and Petr Svácek. Finite element method analysis of flutter: Comparing Scott–Vogelius and Taylor– Hood elements. Journal of Computational and Applied Mathematics, 469 (??):??, December 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042725001761.

Wang:2025:ASW

[Wan25a]  Chunmei Wang.  Auto-

stabilized weak Galerkin finite element methods on polytopal meshes without convexity constraints. Journal of Computational and Applied Mathematics, 466 (??):??, October 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042725000871.

Wang:2025:SHO

[Wan25b] Siyang Wang. Stable and high-order accurate finite difference methods for the diffusive viscous wave equation. Journal of Computational and Applied Mathematics, 463(??):??, August 1, 2025. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724007246.

Wang:2025:CFS

[Wan25c] Zhou-Jing Wang. Closedform solution-based fuzzy utility vectors acquired from trapezoidal fuzzy pairwise comparison matrices using logarithmic quadratic programming for improving fuzzy AHP decision-making systems. Journal of Computational and Applied Mathematics, 468(??):??, November 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704272500161X.

Wang:2025:NSI

[WAXM25] Biyi Wang, Hengbin An, Hehu Xie, and Zeyao Mo. A new subspace iteration algorithm for solving generalized eigenvalue problems in vibration analysis. Journal of Computational and Applied Mathematics, 468(??):??, November 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725001372.

Watkins:2026:EBF

[WBH^{+}26] Jerry Watkins, Luca Bertagna, Graham Harper, Andrew Steyer, Irina Tezaur, and Diana Bull. Entropy-based feature selection for capturing impacts in Earth system models with abrupt forcing. Journal of Computational and Applied Mathematics, 471(??):??, January 1, 2026. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725002389.

Wang:2025:IMB

[WCS25] Lu-Xin Wang, Yang Cao, and Qin-Qin Shen. Improved modulus-based matrix splitting iteration methods for a class of horizontal implicit complementarity problems. Journal of Computational and Applied Mathematics, 454(??):??, January 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724004321.

Wu:2025:EAT

[WCW25] Xiaohui Wu, Yanping Chen, and Yang Wang. Error analysis of two-grid virtual element method for nonlinear parabolic problems on general polygonal meshes. Journal of Computational and Applied Mathematics, 459(??):??, May 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724006174.

Wang:2026:MCF

[WCYL26] Wan-Lun Wang, Luis M. Castro, Shi-Xiu Yu, and Tsung-I Lin. Mixtures of common factor analyzers using the restricted multivariate skew-t distribution for clustering high-dimensional data with missing values. Journal of Computational and Applied Mathematics, 471 (??):??, January 1, 2026. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725002225.

Wang:2025:MFD

[WJ25] Ziyi Wang and Lijian Jiang.

Model free data assimilation with Takens embedding. Journal of Computational and Applied Mathematics, 460(??):??, May 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724006472.

Wang:2025:ELN

[WKQ25] Yingxiao Wang, Lingchen Kong, and Houduo Qi. An efficient Lagrange–Newton algorithm for long-only cardinality constrained portfolio selection on real data sets. Journal of Computational and Applied Mathematics, 461(??):??, June 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042724007015.

Wang:2025:MSI

[WLCM25] Luyao Wang, Guolin Liu, Yang Chen, and Huadong Ma. A method for solving ill-conditioned separable nonlinear least squares problems and its application. Journal of Computational and Applied Mathematics, 466(??):??, October 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042725001396.

Wang:2025:CNA

[WLL25] Qiu-Ya Wang, Cui-Yun Lin, and Cheng-Xue Lao. Crank– Nicolson alternative direction implicit method for two-dimensional variableorder space-fractional diffusion equations with nonseparable coefficients. Journal of Computational and Applied Mathematics, 469 (??):??, December 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042725001694.

Wang:2025:CTS

[WLZZ25] Lixiu Wang, Huiyuan Li, Qian Zhang, and Zhimin Zhang. H(curl^{2})-conforming triangular spectral element method for quad-curl problems. Journal of Computational and Applied Mathematics, 459(??):??, May 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724006101.

Wentland:2025:RIB

[WRBT25] Christopher R. Wentland, Francesco Rizzi, Joshua L. Barnett, and Irina K. Tezaur. The role of interface boundary conditions and sampling strategies for Schwarzbased coupling of projectionbased reduced order models. Journal of Computational and Applied Mathematics, 465(??):??, September 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725000998.

Wu:2026:SIR

[WSKH26] Chai Wah Wu, Mark S. Squillante, Vassilis Kalantzis, and Lior Horesh. Stable iterative refinement for solving linear systems with inaccurate computation. Journal of Computational and Applied Mathematics, 471(??):??, January 1, 2026. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042725002602.

Wu:2025:NEQ

[WSL25] Hongping Wu, Ang Shan, and Xiaosha Li. Nonparametric estimations of quantile residual life function with censored length-biased data. Journal of Computational and Applied Mathematics, 466(??):??, October 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042725001219.

Waheed:2025:GLT

[WuR25] Imtiaz Waheed and Mujeeb ur Rehman. A generalized Legendre transform method and its applications. Journal of Computational and Applied Mathematics, 468(??):??, November 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042725001608.

Wang:2025:OEE

[WWC25] Danxia Wang, Xuliang Wei, and Chuanjun Chen. Optimal error estimates for the semidiscrete and fully-discrete finite element schemes of the Allen–Cahn equation. Journal of Computational and Applied Mathematics, 462 (??):??, July 2025. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724007179.

Wang:2025:CEH

[WX25a] Weiwen Wang and Chuanju Xu. A class of efficient highorder time-stepping methods for the anisotropic phasefield dendritic crystal growth model. Journal of Computational and Applied Mathematics, 453(??):??, January 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724004102.

Wang:2025:ALM

[WX25b] Yinkun Wang and Shuhuang Xiang. Augmented Levin methods for vector-valued highly oscillatory integrals with exotic oscillators and turning points. Journal of Computational and Applied Mathematics, 470(??): ??, December 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725002018.

Wang:2025:SKM

[WYZ25] Ze Wang, Jun-Feng Yin, and Ji-Chen Zhao. The sparse Kaczmarz method with surrogate hyperplane for the regularized basis pursuit problem. Journal of Computational and Applied Mathematics, 454(??):??, January 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704272400431X.

Wang:2025:EWG

[WZ25a] Chunmei Wang and Shangyou Zhang. Efficient weak Galerkin finite element methods for Maxwell equations on polyhedral meshes without convexity constraints. Journal of Computational and Applied Mathematics, 465(??):??, September 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042725000901.

Wang:2025:MHF

[WZ25b] Min Wang and Zhimin Zhang. Mapped Hermite functions and their applications to two-dimensional weakly singular Fredholm-Hammerstein integral equations. Journal of Computational and Applied Mathematics, 465(??):??, September 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042725001001.

Wang:2025:ADC

[WZTL25] Liang Wang, Chunfang Zhang, Yogesh Mani Tripathi, and Yuhlong Lio. Analysis of dependent complementary competing risks data from a generalized inverted family of lifetime distributions under a maximum ranked set sampling procedure with unequal samples. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724005570.

Wu:2025:AMM

[WZZ25] Hongping Wu, Liming Zhang, and Qian Zhao. Alphamaxmin mean-variance reinsuranceinvestment strategy under negative risk dependence between two markets. Journal of Computational and Applied Mathematics, 463 (??):??, August 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725000160.

Xiao:2025:FOE

[XC25] Ming Xiao and Rui Chen.

A first-order energy stable scheme for the Allen–Cahn equation with the Allen– Cahn type dynamic boundary condition. Journal of Computational and Applied Mathematics, 460(??): ??, May 1, 2025. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042724006575.

Xu:2025:PDD

[XCGZ25] Lei Xu, Rongliang Chen, Linyan Gu, and Wu Zhang. A parallel domain decompositionbased implicit finite volume lattice Boltzmann method for incompressible thermal convection flows on unstructured grids. Journal of Computational and Applied Mathematics, 465(??):??, September 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042725000937.

Xiong:2025:EEP

[XCZ25] Wenzhuo Xiong, Xiujun Cheng, and Qifeng Zhang. Error estimates and parallel evaluation of hybrid schemes for parabolic, wave, and Schrödinger equations. Journal of Computational and Applied Mathematics, 465(??):??, September 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042725000949.

Xiong:2025:NFS

[XDZ25] Lianglin Xiong, Junzhou Dai, and Haiyang Zhang. Nonfragile sampled-data control for uncertain fractional-order systems with time-varying delay. Journal of Computational and Applied Mathematics, 461(??):??, June 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724006861.

Xu:2025:ETG

[XGH25] Chenli Xu, Kaixin Gao, and Zheng-Hai Huang. Existence theorems of generalized K-lateral support equilibria via local cooperative strategies for multi-player pure strategy games in tensor form. Journal of Computational and Applied Mathematics, 468(??):??, November 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725001487.

Xu:2025:NMF

[XGX25] Fei Xu, Yasai Guo, and Manting Xie. A novel multilevel finite element method for a generalized nonlinear Schrödinger equation. Journal of Computational and Applied Mathematics, 457 (??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724005296.

Xie:2025:TIG

[XGYH25] Wei Xie, Juan Galvis, Yin Yang, and Yunqing Huang. On time integrators for Generalized Multiscale Finite Element Methods applied to advection–diffusion in highcontrast multiscale media. Journal of Computational and Applied Mathematics, 460(??):??, May 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042724006113.

Xu:2025:WPC

[XHLH25] Wei Xu, Weimin Han, Ting Li, and Ziping Huang. Well-posedness of a class of evolutionary variationalhemivariational inequalities in contact mechanics. Journal of Computational and Applied Mathematics, 459(??): ??, May 15, 2025. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042724006149.

Xie:2025:SCE

[Xie25] Pengcheng Xie. Sufficient conditions for error distance reduction in the l^{2}-norm trust region between minimizers of local nonconvex multivariate quadratic approximates. Journal of Computational and Applied Mathematics, 453(??):??, January 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724003959.

Xiao:2025:DRB

[XJQ25] Zhi-Hua Xiao, Yao-Lin Jiang, and Zhen-Zhong Qi. Dimension reduction based on time-limited cross Gramians for bilinear systems. Journal of Computational and Applied Mathematics, 457 (??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724005508.

Xu:2025:LSP

[XK25] Hailun Xu and Hongmei Kang. Local spline projectors of analysis-suitable Tsplines. Journal of Computational and Applied Mathematics, 462(??):??, July 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724007325.

Xu:2025:BDT

[XL25] Haobo Xu and Fang Li.

Blind deconvolution via a tensor adaptive non-convex total variation prior. Journal of Computational and Applied Mathematics, 463 (??):??, August 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042725000263.

Xu:2025:CSP

[XLLJ25] Yuncheng Xu, Sanyang Liu, Lixia Liu, and Kewei Jie. CVaR stochastic programming model for monotone stochastic tensor complementarity problem by using its penalized sample average approximation algorithm. Journal of Computational and Applied Mathematics, 460(??):??, May 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042724006757.

Xu:2025:DCN

[XLZ25] Bo Xu, Xinliang Liu, and Lei Zhang. Dilated convolution neural operator for multiscale partial differential equations. Journal of Computational and Applied Mathematics, 461(??):??, June 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042724006563.

Xin:2025:ASH

[XNL25] Xiaofang Xin, Guyan Ni, and Ying Li. An alternating shifted higher order power method based algorithm for rank-R Hermitian approximation and solving Hermitian CP-decomposition problems. Journal of Computational and Applied Mathematics, 459(??):??, May 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724006332.

Xu:2025:UMC

[XQW25] Min Xu, Zhongfeng Qin, and Junbin Wang. Uncertain c-means clustering method with application to imprecise observations. Journal of Computational and Applied Mathematics, 459(??): ??, May 15, 2025. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724005934.

Yao:2025:MNM

[Yao25a]  Xudong Yao.  An MP-

Newton method for computing nonlinear eigenpairs and its application for solving a semilinear Schrödinger equation. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724005636.

Younis:2025:UFO

[YAO^{+}25b] Mudasir Younis, Haroon Ahmad, Mahpeyker Ozturk, Fahim Ud Din, and Muhammad Qasim. Unveiling fractional-order dynamics: a new method for analyzing Rössler chaos. Journal of Computational and Applied Mathematics, 468 (??):??, November 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725001530.

Yu:2025:CAR

[YC25] Dansheng Yu and Feilong Cao. Construction and approximation rate for feedforward neural network operators with sigmoidal functions. Journal of Computational and Applied Mathematics, 453(??):??, January 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724003996.

Yu:2025:HOI

[YCLS25] Kaichang Yu, Juan Cheng, Yuanyuan Liu, and Chi-Wang Shu. High-order implicit maximum-principle-preserving local discontinuous Galerkin methods for convection-diffusion equations. Journal of Computational and Applied Mathematics, 470(??):??, December 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725001748.

Yin:2025:CSC

[Yin25] Ruijie Yin. Convergence of some conservative Strang splitting methods for chargedparticle dynamics under a strong magnetic field. Journal of Computational and Applied Mathematics, 460 (??):??, May 1, 2025. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724006782.

Yadav:2025:CFP

[YK25] Kapil Yadav and Deepak Kumar. Common fixed point theorems for discontinuous mappings in M-metric space and numerical approximations. Journal of Computational and Applied Mathematics, 470(??):??, December 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725002341.

Yan:2025:LES

[YL25a] Yujing Yan and Xiaoli Li.

The linear and efficient structure-preserving method for the thermodynamical consistent model of two-phase flow in porous media with rock compressibility. Journal of Computational and Applied Mathematics, 465 (??):??, September 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042725000846.

Yang:2025:SGI

[YL25b] Jiefei Yang and Guanglian Li. On sparse grid interpolation for American option pricing with multiple underlying assets. Journal of Computational and Applied Mathematics, 464(??):??, August 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042725000597.

Yajai:2025:NDR

[YLC25] Watcharaporn Yajai, Wongthawat Liawrungrueang, and Watcharaporn Cholamjiak. A new double relaxed inertial viscosity-type algorithm for split equilibrium problems and its application to detecting osteoporosis health problems. Journal of Computational and Applied Mathematics, 466(??):??, October 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042725001177.

Yan:2025:IPP

[YLW^{+}25] Xihong Yan, Hao Li, Chuanlong Wang, Danqing Zhou, and Junfeng Yang. An improved proximal primal-dual ALM-based algorithm with convex combination proximal centers for equalityconstrained convex programming in basis pursuit practical problems. Journal of Computational and Applied Mathematics, 464(??): ??, August 15, 2025. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725000469.

Yousef:2025:PBL

[You25] Soleiman Yousef. A posterioribased, local multilevel mesh refinement for the Darcy porous media flow problem. Journal of Computational and Applied Mathematics, 454(??):??, January 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724004114.

Yu:2025:PNR

[YPLH25] Xiang Yu, Yehui Peng, Penglin Lou, and Bozhong Huang. Poisson noise removal based on non-convex hybrid regularizers. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042724005375.

Yang:2025:RSO

[YS25] Caijie Yang and Tongjun Sun. Robust secondorder VSBDF2 finite element schemes for parabolic distributed optimal control problems. Journal of Computational and Applied Mathematics, 469(??):??, December 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042725001864.

Yan:2025:RAD

[YSP25] Zaizai Yan, Yanjie Shi, and Xiuyun Peng. Reliability analysis of k/n(G) degradation system under dependent competing failures. Journal of Computational and Applied Mathematics, 461(??):??, June 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042724006927.

Yoshino:2025:CES

[YT25] Natsuki Yoshino and Akira Tanaka. Computationally efficient selection criterion of initial equation in STLNbased structured low-rank approximation. Journal of Computational and Applied Mathematics, 470(??): ??, December 15, 2025.

CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725002171.

Yang:2025:PEE

[YTT25] Shuo Yang, Liutao Tian, and Hongjiong Tian. A posteriori error estimates and adaptivity for the IMEX BDF2 method for nonlinear parabolic equations. Journal of Computational and Applied Mathematics, 457 (??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724005661.

Yu:2025:SNN

[Yu25] Yuexin Yu. Solving nonlinear neutral delay integrodifferential equations via general linear methods. Journal of Computational and Applied Mathematics, 458 (??):??, April 2025. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724005909.

Yang:2025:HWP

[YWS25] Ronghua Yang, Junxian Wang, and Shi Shu. HTL-WASI preconditioner for finite element discretization of complex ADR equation and its application. Journal of Computational and Applied Mathematics, 470 (??):??, December 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042725002407.

Yao:2025:EAM

[YWW25] Guoqing Yao, Zicheng Wang, and Zhongqing Wang. An efficient and accurate mapping method for elliptic equations in irregular annular domains. Journal of Computational and Applied Mathematics, 456(??):??, March 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042724004862.

Yang:2025:TMT

[YWY25] Yang Yang, Guojing Wang, and Jing Yao. Tail moments and tail joint moments for multivariate generalized hyperbolic distribution. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042724005557.

Yan:2025:BIN

[YXM25] Xiong-Bin Yan, Zhi-Qin John Xu, and Zheng Ma. Bayesian Inversion with Neural Operator (BINO) for modeling subdiffusion: Forward and inverse problems. Journal of Computational and Applied Mathematics, 454 (??):??, January 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724004400.

Yao:2025:NCS

[YY25] Jinran Yao and Zhengwei Yin. Numerical contractivity of split-step backward Milstein-type schemes for commutative SDEs with nonglobally Lipschitz continuous coefficients. Journal of Computational and Applied Mathematics, 461(??):??, June 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724006976.

Yue:2025:EIP

[YYJZ25] Yong-Yan Yue, Teng-Teng Yao, Xiao-Qing Jin, and Zhi Zhao. An efficient inertial projection-based algorithm for constrained nonlinear pseudo-monotone equations and its application to logistic regression problems. Journal of Computational and Applied Mathematics, 464(??):??, August 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042725000470.

Zhou:2025:NQG

[ZBK25] Jun Zhou, Pieter J. Barendrecht, and Jirí Kosinka. Numerical quadrature for Gregory triangles. Journal of Computational and Applied Mathematics, 453 (??):??, January 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042724003984.

Zhang:2025:MFE

[ZBN^{+}25] Zhizhuo Zhang, Mikaël Barboteu, Xiaobing Nie, Serge Dumont, Mahmoud Abdel-Aty, and Jinde Cao. Mixed finite element method for multi-layer elastic contact systems. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042724005302.

Zhang:2025:OEE

[ZC25a] Zhuoyue Zhang and Wentao Cai. Optimal L^{2} error estimates of mass- and energyconserved FE schemes for a nonlinear Schrödinger-type system. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724005612.

Zhou:2025:QHO

[ZC25b] Yongxiong Zhou and Ruyun Chen. On quadrature of highly oscillatory Bessel function via asymptotic analysis of simplex integrals. Journal of Computational and Applied Mathematics, 456(??):??, March 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724004886.

Zhang:2025:CPG

[ZCH^{+}25] Zemian Zhang, Yanping Chen, Yunqing Huang, Jian Huang, and Yanping Zhou. A continuous Petrov–Galerkin method for time-fractional Fokker–Planck equation. Journal of Computational and Applied Mathematics, 470 (??):??, December 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725002031.

Zhu:2025:MEE

[ZDD25] Huijian Zhu, Jiawen Ding, and Jiu Ding. A matrix eigenvalue-eigenvector equality for semi-simple eigenvalues. Journal of Computational and Applied Mathematics, 464(??):??, August 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042725000354.

Zhou:2025:MPG

[ZDX25] Qiqi Zhou, Huiming Duan, and Derong Xie. A multivariate partial grey prediction model based on secondorder traffic flow kinematics equation and its application. Journal of Computational and Applied Mathematics, 463(??):??, August 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042725000202.

Zhukova:2025:ANA

[ZFCM25] Aleksandra Zhukova, Anna Flerova, Alexey Chernov, and Georgy Milyuchikhin. Approaches to numerical analysis of optimal control with linear phase constraints on the example of the assets and liabilities management by a bank. Journal of Computational and Applied Mathematics, 453(??):??, January 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042724003790.

Zhang:2025:FGT

[ZGL^{+}25] Xiaoyan Zhang, Guangyu Gao, Yang Li, Zhenwu Fu, and Bo Han. A fast generalized two-point homotopy perturbation iteration with a learned initial value for nonlinear ill-posed problems. Journal of Computational and Applied Mathematics, 464(??):??, August 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725000287.

Zhu:2025:JGS

[ZGWH25] Jianguang Zhu, Wen Gao, Ying Wei, and Binbin Hao. Joint group sparse coding and weighted nuclear norm for Poisson denoising. Journal of Computational and Applied Mathematics, 470 (??):??, December 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725002092.

Zhang:2025:IIN

[ZH25] Xinru Zhang and Qiaolin He.

An immersed interface neural network for elliptic interface problems. Journal of Computational and Applied Mathematics, 459(??):??, May 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724006204.

Zhang:2025:TQF

[Zha25]  Jiehua Zhang.  Tetrahe-

dral quadratic finite volume method schemes for the Stokes equation. Journal of Computational and Applied Mathematics, 463 (??):??, August 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042724007209.

Zhu:2025:FCR

[ZHF25] Ting Ting Zhu, Rong Hu, and Ya Ping Fang. Fast convergence rates and trajectory convergence of a Tikhonov regularized inertial primaldual dynamical system with time scaling and vanishing damping. Journal of Computational and Applied Mathematics, 460(??):??, May 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042724006423.

Zhu:2025:EMC

[ZHLD25] Zhibin Zhu, Jiaqi Huang, Ying Liu, and Yuehong Ding. An efficient modified conjugate gradient algorithm under Wolfe conditions with applications in compressive sensing. Journal of Computational and Applied Mathematics, 458(??):??, April 2025. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724005831.

Zhou:2025:PEO

[Zho25] Ke Zhou. Pricing exchange options under hybrid stochastic volatility and interest rate models. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724005107.

Zhu:2025:GOF

[ZHX25] Jiayi Zhu, Kang Huang, and Yuanjie Xian. General K-order Franklin wavelet method for numerical solution of integral equations. Journal of Computational and Applied Mathematics, 466(??):??, October 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725001220.

Zhang:2025:EFO

[ZLC^{+}25] Yanzhu Zhang, Tingting Liu, Yangquan Chen, Jing Wang, and Mingyu Shi. An efficient fractional-order PDE based image denoising algorithm with optimal adaptive strategy for ultrasound medical image-based diagnostics. Journal of Computational and Applied Mathematics, 460(??):??, May 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042724006484.

Zhang:2025:SAT

[ZLLL25] Zhechen Zhang, Sanyang Liu, Lixia Liu, and Zhiping Lin. A symmetric ADMM-type algorithm for robust tensor completion problems using a regularized SCAD-Schatten-p model with application in color image and video recovery. Journal of Computational and Applied Mathematics, 466(??):??, October 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042725001190.

Zhang:2025:DIE

[ZLS25] Zecheng Zhang, Wing Tat Leung, and Hayden Schaeffer. A discretization-invariant extension and analysis of some deep operator networks. Journal of Computational and Applied Mathematics, 456(??):??, March 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042724004758.

Zeng:2025:IMU

[ZLY25] Min Zeng, Xianlu Liao, and Yongxin Yuan. An iterative method for updating undamped structural systems with connectivity constraints. Journal of Computational and Applied Mathematics, 454(??):??, January 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724004394.

Zhang:2025:ADS

[ZLZ25] Xuexue Zhang, Sanyang Liu, and Nannan Zhao. An accelerated decentralized stochastic optimization algorithm with inexact model. Journal of Computational and Applied Mathematics, 459(??): ??, May 15, 2025. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724006319.

Zhang:2025:ECS

[ZMG25] Xiaohao Zhang, Liquan Mei, and Shimin Guo. Energyconserving SAV–Hermite– Galerkin spectral scheme with time adaptive method for coupled nonlinear Klein– Gordon system in multidimensional unbounded domains. Journal of Computational and Applied Mathematics, 454(??):??, January 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042724004539.

Zhang:2025:BPM

[ZMWD25] Xuejing Zhang, Huanhuan Ma, Xinyu Wu, and Jiansong Deng. A boundary-precise method of numerical integration over implicitly defined regions. Journal of Computational and Applied Mathematics, 468(??):??, November 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042725001517.

Zhu:2025:EMS

[ZNL25] Hong Zhu, Chenchen Niu, and Yongjin Liang. Efficient method for symmetric nonnegative matrix factorization with an approximate augmented Lagrangian scheme. Journal of Computational and Applied Mathematics, 454(??):??, January 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042724004679.

Zhang:2025:TBL

[ZPGY25] Xueqing Zhang, Jianwen Peng, Debdas Ghosh, and Jen-Chih Yao. A threeblock linearized generalized ADMM based iterative algorithm for separable convex programming with application to an image compression problem. Journal of Computational and Applied Mathematics, 462(??):??, July 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724007313.

Zakharova:2025:CHA

[ZS25] Yulia Zakharova and Maria Sakhno. Complexity and heuristic algorithms for speed scaling scheduling of parallel jobs with energy constraint. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704272400503X.

Zhang:2025:EAE

[ZSYZ25] Jun Zhang, Fangying Song, Xiaofeng Yang, and Yu Zhang. Error analysis of the explicitinvariant energy quadratization (EIEQ) numerical scheme for solving the Allen– Cahn equation. Journal of Computational and Applied Mathematics, 457(??): ??, March 15, 2025. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724004734.

Zaman:2025:RKF

[ZuI25] Sakhi Zaman and Siraj ul Islam. Reproducing kernel function-based formulation for highly oscillatory integrals. Journal of Computational and Applied Mathematics, 463(??):??, August 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725000226.

Zhou:2025:MIF

[ZWSC25] Sheng-Wei Zhou, Chun Wen, Zhao-Li Shen, and Bruno Carpentieri. The MFPIO iteration and the FPMPE method for multilinear PageRank computations. Journal of Computational and Applied Mathematics, 454(??):??, January 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724004412.

Zhang:2025:IEB

[ZWZJ25] Xin Zhang, Danxia Wang, Jianwen Zhang, and Hongen Jia. An implicit-explicit BDF2 method with variable time steps for the Ericksen–Leslie model. Journal of Computational and Applied Mathematics, 465 (??):??, September 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725000895.

Zhao:2025:PHM

[ZYJ25] Panpan Zhao, Bo Yu, and Libin Jiao. Polynomial homotopy methods for the equally spaced sparse interpolation problem with jump. Journal of Computational and Applied Mathematics, 468(??):??, November 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725001384.

Zhai:2025:OCU

[ZYTH25] Yuwen Zhai, Guolin Yu, Tian Tang, and Wenyan Han. Optimal conditions in uncertain set-valued optimization problems via secondorder subdifferential involving Minkowski difference with application to zero-sum matrix games. Journal of Computational and Applied Mathematics, 464(??):??, August 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725000457.

Zhang:2025:PPH

[ZZ25a] Yang Zhang and Fan Zhang.

A positivity-preserving HLLCbased discontinuous Galerkin method for weakly compressible two-phase flows. Journal of Computational and Applied Mathematics, 462 (??):??, July 2025. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042724007155.

Zheng:2025:GIE

[ZZ25b] Yawen Zheng and Song-Ping Zhu. A generalized integral equation formulation for pricing American options under regime-switching model. Journal of Computational and Applied Mathematics, 453(??):??, January 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042724002668.

Zhang:2025:TSO

[ZZJJ25] Xin Zhang, Yiran Zhang, Qunzhi Jin, and Yuanfeng Jin. Two sixth-order, L_{∞} convergent, and stable compact difference schemes for the generalized Rosenau– KdV–RLW equation. Journal of Computational and Applied Mathematics, 459(??): ??, May 15, 2025. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042724006307.

Zhang:2025:GDI

[ZZM25] Daochang Zhang, Yue Zhao, and Dijana Mosić. The generalized Drazin inverse of the sum of two elements in a Banach algebra. Journal of Computational and Applied Mathematics, 470 (??):??, December 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725002158.

Zhao:2025:TEM

[ZZX25] Jingjun Zhao, Hao Zhou, and Yang Xu. Two explicit methods for one-sided Lipschitz stochastic differential equations driven by fractional Brownian motion. Journal of Computational and Applied Mathematics, 461(??):??, June 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724007106.